### v1

# Koopman for stochastic dynamics: error bounds for kernel extended dynamic mode decomposition

Abstract.
We prove $L^{\infty}$-error bounds for kernel extended dynamic mode decomposition (kEDMD) approximants of the Koopman operator for stochastic dynamical systems. To this end, we establish Koopman invariance of suitably chosen reproducing kernel Hilbert spaces and provide an in-depth analysis of the pointwise error in terms of the data points. The latter is split into two parts by showing that kEDMD for stochastic systems involves a kernel regression step leading to a deterministic error in the fill distance as well as Monte Carlo sampling to approximate unknown expected values yielding a probabilistic error in terms of the number of samples. We illustrate the derived bounds by means of Langevin-type stochastic differential equations involving a nonlinear double-well potential.
M. Hertel gratefully acknowledges funding by Carl-Zeiss-Stiftung, project KI-MSO-O: AI-supported analysis, modeling and synthesis (design) of organoids, P2024-02-019. F.M. Philipp is grateful for the support of the German Research Foundation (DFG), project numbers 554600805 and 519323897. K. Worthmann gratefully acknowledges funding by the DFG, project ALeSCo: Active Learning for Systems and Control, project number 535860958..
Keywords. Extended dynamic mode decomposition (EDMD), kernel EDMD, kernel regression, Koopman operator, Koopman invariance, Monte Carlo sampling, RKHS, stochastic dynamics, $L^{\infty}$-error bound.
Mathematics subject classications. 37M99, 47B32, 65C05, 65D12
1. Introduction
Nowadays, the Koopman operator is an established mathematical and computational tool for modeling, analysis, and control of dynamical systems, see, e.g., the collection [MaurSusu20] or the recent overview papers [brunton:budisic:kaiser:kutz:2022, strasser2026overview]. The underlying idea is to consider nonlinear dynamics through the lens of observable functions using the linear, but infinite-dimensional Koopman operator, see the original work by B.O. Koopman [Koop31, KoopNeum32] and the seminal paper [Igor2005] by I. Mezić. This enabled, among others, to leverage modern techniques from spectral analysis to investigate the complex dynamical behaviour of the original nonlinear system [rowley2009spectral].
Recently, data-driven approaches like dynamic mode decomposition (EDMD; [WillKevr15, Colb23]) were proposed to approximate the Koopman operator, see, e.g., the recent overview [Colb24], which has lead to a plethora of successful applications, including fluid flows [bagheri:2013, mezic:2013], climate forecasting [hogg:fonoberova:mezic:2020, azencot:erichson:lin:mahoney:2020], molecular dynamics [nuske:klus:2023], power systems technology [susuki:mezic:raak:hikihara:2016], quantum systems [klus:nuske:peitz:2022], robotics [mamakoukas:castano:tan:murphey:2019, bruder:fu:gillespie:remy:vasudevan:2020, haggerty:banks:kamenar:cao:curtis:mazic:hawkes:2023, rosenfelder:bold:eschmann:eberhard:worthmann:ebel:2024], and human-centered applications such as grip force prediction for robotic rehabilitation [bazina:kamenar:fonoberova:mezic:2025]. In EDMD, snapshots of the underlying system are employed to approximate the Koopman operator. It has been shown in [KordaMezi18] that EDMD converges in the infinite-data and infinite-dictionary (infinitely many observable functions) corresponding to a complexity solvability index two [colbrook2022foundations]. Then, first finite-data results on the estimation error for ergodic sampling for deterministic systems were provided in [Mezi22] followed by [ZhanZuaz23] analyzing also the $L^{2}$ projection error onto a finite element dictionary, cf. also the works [NuskPeit23, PhilScha24] providing error bounds for stochastic and control systems.
More recently, kernel EDMD approximation schemes, which allow a data-informed choice of the dictionary, were proposed, see, e.g., [WillRowl15, klus:nuske:hamzi:2020]. Several works have already addressed $L^{2}$-error bounds for kernel EDMD: [PhilScha24:kernel] provides these in terms of the Mercer eigenvalues of the chosen kernel in the case of ergodic sampling along a trajectory, while the bounds in [PhilSchaWort2025, OlguOsseRami2025], are derived for i.i.d. sampling. The work [KohnPhil24] shows first pointwise (i.e., $L^{\infty}$) error bounds for approximations using Wendland kernels. However, while kernel EDMD is successfully used for stochastic systems, pointwise error bounds are still missing – to the best of the authors’ knowledge.
In this paper, we propose a readily applicable Monte Carlo framework for estimation of kernel matrices to compute approximations of the Koopman operator for stochastic systems. Then, we provide an in-depth error analysis of the scheme. To this end, we establish Koopman invariance of (fractional) Sobolev spaces based on novel sufficient conditions on a Radon-Nikodym density associated with the stochastic dynamical system. Then we split the approximation error into a deterministic and a probabilistic part, where the latter results from Monte Carlo sampling of the expected value. Combining these ingredients, we then present novel full approximation error bounds for kernel EDMD for stochastic systems using Matérn and Wendland kernels. Finally, we present numerical simulations for systems governed by SDEs as a proof of concept.
The paper is organized as follows. In Section 2, we introduce the Koopman operator for stochastic systems, propose a kernel EDMD approach for its approximation and state the main result in Theorem 2.6, providing a pointwise total error bound for the kernel EDMD approximant with high probability. To prove this error bound, we first derive necessary and sufficient conditions for invariance of Sobolev spaces under the stochastic Koopman operator in Section 3, cf. Theorem 3.2. Then we prove the approximation error bounds in Section 4 and illustrate them using numerical examples in Section 5. Section 6 concludes and gives an outlook.
Notation For a set $\Omega\subset\mathbb{R}^{n}$, we denote the Borel sigma algebra on $\Omega$ (w.r.t. the standard topology) by $\mathcal{B}_{\Omega}$. The Lebesgue measure of a Borel set $\Delta\in\mathcal{B}_{\Omega}$ is denoted by $|\Delta|$; $dx$ or $dy$ indicate integrals w.r.t. Lebesgue measure. $\chi_{A}$ denotes the characteristic/indicator function of a set $A\subset\mathbb{R}^{n}$. The Radon-Nikodym density of a measure $\mu$ w.r.t. the Lebesgue measure will be denoted by $\frac{d\mu}{dx}$. The identity operator on some space is denoted by $\mathcal{I}$, the identity matrix especially is denoted by $I$. For Banach spaces $\mathcal{S}$ and $\mathcal{T}$, $L(\mathcal{S},\mathcal{T})$ denotes the space of bounded linear operators from $\mathcal{S}$ to $\mathcal{T}$. We write $L^{p}(\Omega)$, $p\in[1,\infty]$, for the usual Lebesgue spaces with respect to Lebesgue measure and $C_{b}(\Omega)$ for the space of bounded and continuous functions on $\Omega$; occasionally, we just write $L^{p}$ or $C_{b}$ if the domain is clear from context. For $f:\Omega\to\mathbb{R}$, we denote $\|f\|_{\infty}:=\sup_{x\in\Omega}|f(x)|$. The $\alpha$-th weak derivative of a function $f:\Omega\to\mathbb{R}$, $\alpha\in\mathbb{N}^{n}$, is denoted by $D^{\alpha}f$. In the case of a function $f:\Omega^{L}\to\mathbb{R}$, $L\in\mathbb{N}$, taking multiple arguments, we write $D_{l}^{\alpha}f$ for the $\alpha$-th derivative of $f$ “w.r.t. the $l$-th variable”, $l\in\{1,\dots,L\}$. For $d\in\mathbb{N}$, we abbreviate $\mathbbm{1}=\{v\in\mathbb{R}^{d}:|v_{i}|=1\}$.
2. Koopman operator for stochastic systems: kernel EDMD and error analysis
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded domain with Lipschitz boundary and $\mathcal{B}_{\Omega}$ be the Borel sigma algebra on $\Omega$. Moreover, let $X$ and $Y$ be $\Omega$-valued random variables. It is well known (see, e.g., [Klen2008]) that there exists a function $\rho:\Omega\times\mathcal{B}_{\Omega}\to[0,1]$ such that
  1. i)
    $\rho_{x}:=\rho(x,\cdot)$ is a probability measure for each $x\in\Omega$;
  2. ii)
    For each $\Delta\in\mathcal{B}_{\Omega}$, the function $\rho(\cdot,\Delta):\Omega\to[0,1]$, is measurable;
  3. iii)
    For each $\Delta\in\mathcal{B}_{\Omega}$: $\rho(x,\Delta)=\mathbb{P}(Y\in\Delta\,|\,X=x)$ for $\mathbb{P}^{X}$-a.e. $x\in\Omega$.
A function with the first two properties is called a Markov transition kernel or Markov probability kernel; the third property makes $\rho$ a regular version of the conditional distribution of $Y$ given $X$ (see, e.g., [Kall2001]). In what follows, we consider stochastic dynamical systems of the form
$Y\mid\{X=x\}\sim\rho_{x},\quad x\in\Omega,$ (1)
where the distribution of $Y$ is conditioned on the initial value $X=x$. Specifically, one might consider a discrete-time Markov process, possibly given by the flow of a stochastic differential equation (SDE) for a time step $T>0$, i.e.,
$dX_{t}$ $=b(X_{t})\,dt\,+\sigma(X_{t})\,dW_{t}$ (2)
with drift vector field $b:\mathbb{R}^{n}\to\mathbb{R}^{n}$, diffusion vector field $\sigma:\mathbb{R}^{n}\to\mathbb{R}^{n\times n}$, and $n$-dimensional Brownian motion $W_{t}$.
2.1. The Koopman operator for stochastic systems
The Koopman operator $\mathcal{K}$ associated with $X$ and $Y$ maps functions $f$ on $\Omega$ to functions $\mathcal{K}f$ on $\Omega$ and is defined as
$(\mathcal{K}f)(x):=\mathbb{E}\big[f(Y)\,|\,X=x]=\int_{\Omega}f\,d\rho_{x},\quad x\in\Omega.$ (3)
In the next chapter, we elaborate necessary and sufficient conditions for the invariance of Sobolev spaces $H^{s}(\Omega)$ under the Koopman operator. Those are of special interest since they arise as native spaces, or reproducing kernel Hilbert spaces, of certain kernels, for which there exists an extensive approximation framework, see Section 2.2, that allows for error bounds (Section 4). Before analyzing the Koopman operator on Sobolev spaces, we require the invariance of the space of square-integrable functions $L^{2}(\Omega)=H^{0}(\Omega)$ under $\mathcal{K}$.
Lemma 2.1.
If the condition
$\exists C>0:\quad\int\rho_{x}(\Delta)\,dx\,\leq\,C|\Delta|\qquad\,\forall\Delta\in\mathcal{B}_{\Omega},$ ($L^{1}$)
holds, then $\mathcal{K}L^{2}(\Omega)\subset L^{2}(\Omega)$, and $\mathcal{K}:L^{2}(\Omega)\to L^{2}(\Omega)$ is a bounded linear operator.
Proof.
Indeed, Cauchy-Schwarz yields $|(\mathcal{K}f)(x)|^{2}\leq(\mathcal{K}f^{2})(x)$, so that for simple functions $f=\sum_{j=1}^{m}c_{j}\mathbbm{1}_{\Delta_{j}}$ with mutually disjoint sets $\Delta_{j}$,
$\int|(\mathcal{K}f)(x)|^{2}\,dx\leq\int(\mathcal{K}f^{2})(x)\,dx=\sum_{j=1}^{m}c_{j}^{2}\int\rho_{x}(\Delta_{j})\,dx\,\leq\,C\sum_{j=1}^{m}c_{j}^{2}|\Delta_{j}|=C\|f\|_{L^{2}(\Omega)}^{2}.$
By density of the simple functions in $L^{2}(\Omega)$, it follows that $\mathcal{K}$ is well defined on $L^{2}(\Omega)$ and that $\|\mathcal{K}f\|_{L^{2}(\Omega)}\leq C^{1/2}\|f\|_{L^{2}(\Omega)}$ for all $f\in L^{2}(\Omega)$. ∎
Note that, by the Fubini-Tonelli theorem, ($L^{1}$) implies $L^{1}$-invariance of the Koopman operator. More generally, it can be shown that for a bounded domain $\Omega$, $L^{1}$-invariance (w.r.t. Lebesgue measure) implies $L^{q}$-invariance for $q>1$.
From now on, we consider the Koopman operator acting on functions $\psi\in L^{2}(\Omega)$.
Remark 2.2.
While we consider $L^{2}(\Omega)=L^{2}_{{\boldsymbol{\lambda}}}(\Omega)$ with the Lebesgue measure $\boldsymbol{\lambda}$, our results can be extended to other reference measures – depending on the underlying stochastic system. In the case of $Y=X_{T}$ and $X=X_{0}$ for some Markov process $(X_{t})_{t\geq 0}$, the invariant measure $\mu$ of the process would be an intuitive choice, see, e.g., [NuskPeit23].
2.2. Recap on deterministic systems: kernel EDMD and error bound
Let $\Omega\subset\mathbb{R}^{n}$ as before. In the deterministic case, the dynamics are given by
$x^{+}=F(x),$
with $F:\Omega\to\Omega$, which may be considered as a particular case of the stochastic case (1). In particular, we may consider the SDE (2) with $\sigma\equiv 0$ giving rise to the ordinary differential equation $\dot{x}=b(x)$. In the deterministic case, the transition kernel is given by
$\rho_{x}(A)=\delta_{F(x)}(A)$
such that the Koopman identity (3) reduces to
$\mathcal{K}\psi(x)=\psi(F(x)),$
i.e., $\mathcal{K}\psi=\psi\circ F$. In this case, it is often assumed that $F:\Omega\to\Omega$ is a diffeomorphism with $\inf_{x\in\Omega}|\det F^{\prime}(x)|=c>0$ and $\sup_{x\in\Omega}\|F^{\prime}(x)\|<\infty$ [KohnPhil24]. While the latter condition corresponds to the Lipschitz continuity of the drift function in (2), the assumption on the Jacobi determinant is a special case of ($L^{1}$). Specifically, applying the change of variables formula gives
$\int\delta_{F(x)}(\Delta)\,dx=\left|F^{-1}(\Delta)\right|=\int_{\Delta}|\det(F^{-1})^{\prime}(y)|\,dy\leq\frac{1}{c}\,|\Delta|,$
i.e., ($L^{1}$) also holds in the deterministic case.
We briefly recall Extended Dynamic Mode Decomposition (EDMD; [WillKevr15]) and its extension kernel EDMD, which are well-established numerical methods to approximate the Koopman operator $\mathcal{K}$ from data, see also [KutzBrun16] or the recent overview article [Colb24]. EDMD requires two ingredients: First, let $d\in\mathbb{N}$ snapshots $(x_{i},y_{i})$ with $y_{i}=F(x_{i})$, $i\in\{1,\dots,d\}$, be given. Then, we partition the snapshot data into the sets
$\mathcal{X}=\{x_{1},\dots,x_{d}\}\quad\text{ and }\quad\mathcal{Y}=\{y_{1},\dots,y_{d}\}.$
Second, let $M\in\mathbb{N}$ observable $L^{2}$-functions $\psi_{1},\dots,\psi_{M}$ be given, for which we use the notation $\Psi(x)=(\psi_{1}(x),\dots,\psi_{M}(x))^{\top}$ and define the data matrices
$\Psi_{\mathcal{X}}=(\Psi(x_{1}),\dots,\Psi(x_{d}))\in\mathbb{R}^{M\times d},\qquad\Psi_{\mathcal{Y}}=(\Psi(y_{1}),\dots,\Psi(y_{d}))\in\mathbb{R}^{M\times d}.$
The subspace $\mathcal{V}=\mathrm{span}\{\psi_{1},\dots,\psi_{M}\}\subset L^{2}(\Omega)$ spanned by the observables is a finite-dimensional space. From the Koopman identity, we get
$\mathcal{K}\psi_{i}(x_{j})=\psi_{i}(F(x_{j}))=\psi_{i}(y_{j})\qquad\forall\,(i,j)\in\{1,\ldots,M\}\times\{1,\dots,d\},$ (4)
that is, the Koopman operator maps the components of $\Psi_{\mathcal{X}}$ to the components of $\Psi_{\mathcal{Y}}$. Hence, to derive a data-driven approximation of the Koopman operator, restricted to $\mathcal{V}$, we seek for an approximation $K\in\mathbb{R}^{M\times M}$ that satisfies this identity approximately, that is,
$\Psi_{\mathcal{Y}}\approx K\Psi_{\mathcal{X}}.$ (5)
Assuming that $\Psi_{\mathcal{X}}$ has full row rank (which is equivalent to strong linear independence of the observables $\psi_{1},\dots,\psi_{M}$ w.r.t. Lebesgue measure, see [PhilScha24, Lemma C.3]), an approximation can be computed as a solution to the least squares problem
$\min_{K\in\mathbb{R}^{M\times M}}\|K\Psi_{\mathcal{X}}-\Psi_{\mathcal{Y}}\|_{F}^{2}.$ (6)
Direct calculations using the normal equations yield the EDMD approximation
$K=\Psi_{\mathcal{Y}}\Psi_{\mathcal{X}}^{\dagger}=\Psi_{\mathcal{Y}}\Psi_{\mathcal{X}}^{\top}(\Psi_{\mathcal{X}}\Psi_{\mathcal{X}}^{\top})^{-1}.$ (7)
As illustrated above, EDMD crucially depends on two ingredients: a finite number of data pairs $\{(x_{i},y_{i})\mid y_{i}=F(x_{i}),\,i\in\{1,\dots,d\}\}$ and the choice of the observables. Convergence in the infinite-data limit has been shown in [KordaMezi18] and further specified w.r.t. the estimation error (number of data pairs $d$) in [Mezi22, NuskPeit23, PhilScha24]. The projection error is typically analyzed using finite-element observables as proposed in [ZhanZuaz23] and extended to control systems in [SchaWort23]. However, the latter only yields $L^{2}$-error bounds.
A particular variant of EDMD using a data-informed set of observables is kernel EDMD (kEDMD; [WillRowl15]), sharing two crucial features: First, from an application point of view, the choice of the observable functions is performed in a data-driven manner, eliminating fiddly and time-consuming tuning. Second, due to the rich mathematical structure of reproducing kernel Hilbert spaces, allows for pointwise deterministic error bounds.
[…]
$\psi(x)=\langle\psi,\Phi_{x}\rangle_{\mathbb{H}}\quad\forall x\in\mathbb{R}^{n}$ (8)
with the inner product $\langle\cdot,\cdot\rangle_{\mathbb{H}}$ holds. Then, the function $\Phi_{x}:\mathbb{R}^{n}\to\mathbb{R}$, $\Phi_{x}(y)=k(x,y)\in\mathbb{H}$, commonly termed the canonical feature of $k$ at $x\in\mathbb{R}^{n}$, plays a central role: We choose or sample points $\mathcal{X}=\{x_{1},\dots,x_{d}\}$ as before and let $y_{i}=F(x_{i})$. The dictionary is then given by the canonical features $\Phi_{x}=k(x,\cdot)$, $x\in\Omega$, of the data points, i.e.
$\mathcal{V}=\mathcal{V}_{\mathcal{X}}:=\mathrm{span}(\mathcal{D})=\mathrm{span}(\{\Phi_{x_{1}},\dots,\Phi_{x_{d}}\}).$
We immediately see that $M=d$ holds, i.e., we have exactly as many observables $\Phi_{x_{i}}$ as data points $x_{i}$. The entries in $\Psi_{X}\in\mathbb{R}^{d\times d}$ are easily calculated as $(\Psi_{X})_{i,j}=\psi_{i}(x_{j})=\Phi_{x_{i}}(x_{j})=k(x_{i},x_{j})$. The entries in $\Psi_{Y}$ are obtained via $(\Psi_{Y})_{i,j}=\psi_{i}(y_{j})=\Phi_{x_{i}}(F(x_{j}))=k(x_{i},F(x_{j}))$. Furthermore, we set
$\mathbf{K}_{\mathcal{X}}:=\mathbf{K}_{\mathcal{X},\mathcal{X}}=(k(x_{i},x_{j}))_{i,j=1}^{d}=\Psi_{X},\quad\mathbf{K}_{\mathcal{X},F(\mathcal{X})}=(k(x_{i},F(x_{j}))_{i,j=1}^{d}=\Psi_{Y}$
to which we will refer to as kernel matrices. Plugging these canonical features into (6) yields the regression problem
$\min_{K\in\mathbb{R}^{d\times d}}\left\|K\mathbf{K}_{\mathcal{X}}-\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\right\|_{F}^{2}$ (9)
Assuming that the data points are pairwise distinct, $\mathbf{K}_{\mathcal{X}}$ is symmetric positive definite (hence invertible) due to the symmetry and strict positive definiteness of the underlying kernel $\mathbf{k}$. [^2] We explicitly speak of strict positive definiteness above, since positive definiteness of kernels is often defined via positive semi-definiteness of kernel matrices in the literature, see [KohnPhil24]. The regression problem (9) now reduces to (exactly) solving a linear system, whose solution $K=\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\mathbf{K}_{\mathcal{X}}^{-1}$ corresponds to (7). This matrix not only represents the action of the Koopman operator for our data, but gives the matrix representation of $S_{\mathcal{X}}\mathcal{K}|_{\mathcal{V}}$ in the basis of the canonical features, where $S_{\mathcal{X}}:=S_{V_{\mathcal{X}}}$ denotes the orthogonal projection onto $\mathcal{V}$. Specifically, for some $g\in\mathbb{H}$, the projection onto $\mathcal{V}_{\mathcal{X}}$ is given by
$S_{\mathcal{X}}g=\sum_{i=1}^{d}\left(\mathbf{K}_{\mathcal{X}}^{-1}g_{\mathcal{X}}\right)_{i}\Phi_{x_{i}}=g_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}$ (10)
with $g_{\mathcal{X}}=(g(x_{1}),\dots,g(x_{d}))^{\top}$ and $\mathbf{k}_{\mathcal{X}}=(\Phi_{x_{1}},\dots,\Phi_{x_{d}})^{\top}$, see [KohnPhil24, Proposition 2.1]. For $f\in\mathcal{V}_{\mathcal{X}}$, we have some $\alpha_{f}\in\mathbb{R}^{d}$ such that $f=\alpha_{f}^{\top}\mathbf{k}_{\mathcal{X}}$. Setting $g=\mathcal{K}|_{\mathcal{V}_{\mathcal{X}}}f$, we get $g_{\mathcal{X}}=\mathbf{K}_{\mathcal{X},F(\mathcal{X})}^{\top}\alpha_{f}$ and
$S_{\mathcal{X}}\mathcal{K}|_{\mathcal{V}_{\mathcal{X}}}f=S_{\mathcal{X}}g=g_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}=\alpha_{f}^{\top}\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}.$ (11)
This means the kEDMD-approximant $K=\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\mathbf{K}_{\mathcal{X}}^{-1}$ [KohnPhil24, Proposition 3.2] maps (the coefficients of) $f\in\mathcal{V}_{\mathcal{X}}$ to (the coefficients of) $S_{\mathcal{X}}\mathcal{K}|_{\mathcal{V}_{\mathcal{X}}}f$ linearly (in the respective canonical bases). To extend this operator to the entire space $\mathbb{H}$, one may simply add another orthogonal projection from the right, i.e., replace $f\in\mathcal{V}_{\mathcal{X}}$ by $S_{\mathcal{X}}g$ for any $g\in\mathbb{H}$. From (10), we get the representation $S_{\mathcal{X}}g=\alpha_{S_{\mathcal{X}}g}^{\top}\mathbf{k}_{\mathcal{X}}=g_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}$. In total, this yields the Koopman approximant
$\widehat{\mathcal{K}}=S_{\mathcal{X}}\mathcal{K}S_{\mathcal{X}}:\mathbb{H}\mapsto V_{\mathcal{X}},$ (12)
which, using (11), may be evaluated by
$\widehat{\mathcal{K}}f=S_{\mathcal{X}}\mathcal{K}S_{\mathcal{X}}f=S_{\mathcal{X}}\mathcal{K}|_{\mathcal{V}_{\mathcal{X}}}(S_{\mathcal{X}}f)=\alpha_{S_{\mathcal{X}}f}^{\top}\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}.$ (13)
The main result regarding this approximation obtained in [KohnPhil24, Theorem 3.4] may be summarized as follows.
Theorem 2.3.
Suppose that $\mathcal{K}\mathbb{H}\subset\mathbb{H}$. Then $\mathcal{K}:\mathbb{H}\to\mathbb{H}$ is a bounded operator, and the approximant $\widehat{\mathcal{K}}=S_{\mathcal{X}}\mathcal{K}S_{\mathcal{X}}$ satisfies
$\sup_{\|f\|_{\mathbb{H}}\leq 1}\|(\mathcal{K}-\widehat{\mathcal{K}})f\|_{\infty}\leq\,(\|\mathcal{K}\|_{C_{b}\to C_{b}}+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}.$
The above error bound depends on two crucial features:
  1. (1)
    Invariance property. First, it is assumed that the Koopman operator is indeed a bounded operator on $\mathbb{H}$. Due to closedness of the Koopman operator on RKHS, it suffices to show that $\mathcal{K}\mathbb{H}\subset\mathbb{H}$ (see [KohnPhil24, Lemma 3.1] and Proposition 2.4 below).
  2. (2)
    Approximation error. Second, for kEDMD, the regression problem is indeed solved by interpolation and the interpolation error $\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}$ has to be controlled; this is usually done in terms of the fill distance $h_{\mathcal{X}}$ of the data set $\mathcal{X}$, i.e., one usually has
    $\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}\leq Ch_{\mathcal{X}}^{r}\qquad\text{with}\qquad h_{\mathcal{X}}=\underset{z\in\Omega\;x\in\mathcal{X}}{\sup\min}\,\|z-x\|_{2}$ (14)
    for a suitable rate $r>0$ and $C\geq 0$ using existing results from kernel-based approximation theory, see e.g. [Wend04, Section 11.2].
The invariance property (1) is a subtle detail that has to be verified rigorously and cannot be taken for granted. In particular, there are RKHS for which $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ implies that the underlying dynamics is affine linear, such as the Gaussian RKHS, see [GonzAbud23]. For (fractional) Sobolev spaces $H^{s}$ (see Section 3.2 for an introduction), however, which are RKHS if $s>n/2$ and are generated by Wendland or Matérn kernels (see below for a precise definition), Koopman invariance was proven in [KohnPhil24, Theorem 4.2] in the deterministic setting. The central assumption is that the flow is smooth enough, such that the Koopman composition operator preserves Sobolev regularity. On the other hand, invariance of some RKHS under the Koopman operator is sufficient for boundedness, as shown in [KohnPhil24, Lemma 3.1] for the deterministic case, and analogously in Proposition 2.4 below for the stochastic case.
Proposition 2.4.
Let $\mathbb{H}\subset C_{b}(\Omega)$ be a RKHS that is invariant under $\mathcal{K}$. Then $\mathcal{K}:\mathbb{H}\to\mathbb{H}$ is bounded.
Proof.
Since $\mathbb{H}\subset C_{b}(\Omega)$, the corresponding kernel is bounded. Let $f_{n},f,g\in\mathbb{H}$ such that $f_{n}\to f$ and $\mathcal{K}f_{n}\to g$ in $\mathbb{H}$. Recall the canonical features $\Phi_{x}$ for $x\in\Omega$. Then, the reproducing property (8) yields
$g(x)$ $=\langle g,\Phi_{x}\rangle_{\mathbb{H}}=\lim_{n\to\infty}\langle\mathcal{K}f_{n},\Phi_{x}\rangle_{\mathbb{H}}=\lim_{n\to\infty}\mathcal{K}f_{n}(x)=\lim_{n\to\infty}\int f_{n}(y)\,d\rho_{x}(y).$
Now, $f_{n}(y)=\langle f_{n},\Phi_{y}\rangle_{\mathbb{H}}\to\langle f,\Phi_{y}\rangle_{\mathbb{H}}=f(y)$ for all $y\in\Omega$ as $n\to\infty$ by continuity of the scalar product, and $|\langle f_{n},\Phi_{y}\rangle_{\mathbb{H}}|\leq\|f_{n}\|_{\mathbb{H}}\|\Phi_{y}\|_{\mathbb{H}}\leq C$ for all $n\in\mathbb{N}$. Thus, by dominated convergence, $g(x)=\mathcal{K}f(x)$, which proves that $\mathcal{K}$ is closed and, thus, bounded. ∎
To sum up, the assumption of Theorem 2.3 is met for Sobolev RKHS, and error bounds of the projection error (14) are readily available from the literature for the generating kernels, leading to a pointwise and deterministic error bound for EDMD for deterministic systems.
2.3. Kernel EDMD for stochastic systems
In contrast to the deterministic case, the stochastic Koopman operator is defined by means of conditional expectations. These are usually not readily available in applications with unknown dynamics such that the analogue to the regression problem (6) is not directly solvable if only observations are available. More precisely, to implement the counterpart of the approximate identity (5), i.e.,
$K\Psi_{\mathcal{X}}\approx\Psi_{\mathcal{Y}}=(\mathbb{E}[\Psi_{i}(Y)\mid X=x_{j}])_{(i,j)\in[M]\times[d]},$ (15)
or the corresponding regression problem (6), the data matrix $\Psi_{\mathcal{Y}}$ has to be further approximated. To this end, we build upon the rich theory of Monte Carlo methods to perform an entry-wise approximation, where we draw $m_{j}\in\mathbb{N}$ i.i.d. samples $y_{j}^{(l)}\sim\rho_{x_{j}}$, $l=1,\dots,m_{j}$, and invoke the law of large numbers such that
$\mathbb{E}[\Psi_{i}(Y)\mid X=x_{j}]\approx\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}\Psi_{i}(y_{j}^{(l)}).$ (16)
For the kernel case, we then have, in view of $\Psi_{i}=k(x_{i},\cdot)$,
$K\mathbf{K}_{\mathcal{X}}\approx\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\qquad\mathrm{with}\qquad(\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}})_{ij}=\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}k(x_{i},y_{j}^{(l)})$ (17)
This yields the matrix approximant $K=\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\mathbf{K}_{\mathcal{X}}^{-1}$, which corresponds to the operator approximant
$\widehat{\mathcal{K}}^{\mathrm{MC}}f=f_{\mathcal{X}}^{\top}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{k}_{\mathcal{X}}$ (18)
which mirrors the deterministic approximant from (13). However, the entry-wise Monte Carlo approximation $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ introduces a further estimation error since (16) only holds approximately. In the ideal case where $\Psi_{\mathcal{Y}}=\mathbf{K}_{\mathcal{X},\mathcal{Y}}$ is known exactly, this estimation error vanishes, and the error bound from Theorem 2.3 holds.
Remark 2.5.
Another possible design choice more in line with the covariance operator approach circumvents the estimation of estimating the expected value under certain assumptions on the underlying stochastic system and sampling scheme (see [NuskPeit23, PhilSchaWort2025] for details). In the setting of a governing stochastic differential equation, the authors either sample the design points in $\mathcal{X}$ from the invariant measure $\mu$ or assume ergodicity and sample from a single long trajectory $x_{0},x_{1},\dots,x_{d}$. In this paper, we pursue the first approach. Future work utilizing the second approach could be the subject of further work and will be discussed later.
2.4. Main contribution: error analysis of kEDMD for stochastic systems
In this part, we summarize our main results. In addition, we provide a roadmap for the subsequent sections, where the derivations are presented in full detail. Our first main result is an abstract error bound for general RKHS resulting in a pointwise and probabilistic error bound on the approximation with kEDMD for stochastic systems.
Theorem 2.6 ($L^{\infty}$-error bound).
Let $\mathbb{H}$ be an RKHS of functions over $\Omega$, and assume $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b}(\Omega),C_{b}(\Omega))$. Furthermore, for $\mathcal{X}=\{x_{1},\ldots,x_{d}\}$, let $m_{j}=m\in\mathbb{N}$ i.i.d. samples from every $\rho_{x_{j}}$, $j\in\{1,\dots,d\}$ be given. Then, for every $\varepsilon>0$ and for all $f\in\mathbb{H}$, $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
$\|(\mathcal{K}-\widehat{\mathcal{K}}^{\mathrm{MC}})f\|_{\infty}\leq\big[1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\big]\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}+\varepsilon\cdot\|k\|_{\infty}^{1/2}\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$ (19)
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right)$.
The proof of the above theorem consists of two parts. First, the finite number of data points and observables given by $\mathcal{X}\subset\Omega$ introduces a deterministic error. Second, we can only estimate the entries of the “propagation matrix” $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ which are conditional expected values. Our estimator $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ is random itself, resulting in a probabilistic estimation error. Thus, we split the approximation error as follows:
$\|\mathcal{K}-\widehat{\mathcal{K}}^{\mathrm{MC}}\|_{\mathbb{H}\to C_{b}}\leq\underbrace{\|\mathcal{K}-\widehat{\mathcal{K}}\|_{\mathbb{H}\to C_{b}}}_{\text{deterministic error}}+\underbrace{\|\widehat{\mathcal{K}}-\widehat{\mathcal{K}}^{\mathrm{MC}}\|_{\mathbb{H}\to C_{b}}}_{\text{probabilistic error}}.$
The probabilistic error is controlled by the accuracy of the entry-wise Monte Carlo estimator (5), i.e., the number $m$ of i.i.d. samples from $\rho_{x_{j}}$. Besides, the two main components in Theorem 2.6 are the operator norms (which are assumed to be finite) and the projection error governed by $\mathcal{I}-S_{\mathcal{X}}$. These two components strongly rely on the chosen kernel and thus deserve separate treatments which constitute the second and third main result of this work.
Boundedness of the Koopman operator. Invariance of reproducing kernel Hilbert spaces for Koopman operators is a non-trivial issue. Hence, in Section 3 we provide characterizations and sufficient conditions for the invariance of Sobolev spaces. Here, we present the most important and practically relevant result. Its proof follows from the more general result of Theorem 3.2.
Theorem 2.7.
If the transition kernel $\rho_{x}$ is given by a density $p:\Omega\times\Omega\to[0,\infty)$, i.e.,
$\rho_{x}(\Delta)=\int_{\Delta}p(x,y)\,dy\qquad\forall x\in\Omega,\,\Delta\in\mathcal{B}_{\Omega},$
that satisfies
$\sum_{|\alpha|\leq\ell}\left\|D_{1}^{\alpha}p(\cdot,\cdot)\right\|_{L^{2}(\Omega\times\Omega)}\leq C$ (20)
for some constant $C>0$ and $\ell\in\mathbb{N}$, then $\mathcal{K}\in L(H^{s},H^{s})$ for all $s\in(0,\ell]$.
We point out that Koopman invariance, as shown in [KohnPhil24] for suitably chosen reproducing kernel Hilbert spaces, was extended to stochastic systems in the preprint [OlguOsseRami2025], under the stronger regularity assumption $p\in C^{m}(\Omega\times\Omega)$ for the transition density. On the one hand, we generalize this result by significantly weakening the assumptions. On the other hand, we also provide necessary conditions and, thus, a complete characterization – even without assuming that a density for the transition kernel $\rho_{x}$ exists.
Projection error bounds in terms of the fill distance. In this work, we apply our results to two popular choices of kernels, Wendland kernels [Wend04] and Matérn kernels [BerlThom2004] which are in particular consistent with the above mentioned invariance as they generate Sobolev spaces (see [Adams2003] and Section 3.2 for an introduction). We briefly present the error bounds for Wendland kernels, as the bounds for Matérn kernels (see Appendix C for an introduction) are structurally very similar.
The Wendland kernel $k_{n,l}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ is bounded, continuous and strictly positive-definite (see [Wend04] and Appendix B for details). For $\sigma_{n,l}=\frac{n+1}{2}+l$, it generates the (fractional) Sobolev space $H^{\sigma_{n,l}}(\mathbb{R}^{n})$ as RKHS with equivalence of the standard Sobolev norm and the induced (native) RKHS norm. Restricted to a bounded domain $\Omega\subset\mathbb{R}^{n}$ with Lipschitz boundary, as is of interest in the setting of this paper, the Wendland kernel generates the corresponding Sobolev space $H^{\sigma_{n,l}}(\Omega)$ [KohnPhil24, Theorem 4.1]. Results from kernel approximation theory allow for a more explicit bound on the projection error in terms of the fill distance.
Theorem 2.8 ($L^{\infty}$-error bound for Wendland native spaces).
Let $l\in\mathbb{N}$, define $\sigma_{n,l}=\frac{n+1}{2}+l$ and denote by $\mathbb{H}$ the Sobolev space $H^{\sigma_{n,l}}(\Omega)$ induced by the Wendland kernel $k_{n,l}$ as RKHS. Assume that $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b}(\Omega),C_{b}(\Omega))$. Then, for every $\varepsilon>0$ and all $f\in\mathbb{H}$ with $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
$\|(\mathcal{K}-\mathcal{K}^{\mathrm{MC}})f\|_{\infty}\leq$ $C_{1}\,(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot h_{\mathcal{X}}^{l+1/2}+\varepsilon\cdot C_{2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$ (21)
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2C_{2}^{2}}\right)$, with constants $C_{1},C_{2}>0$ only depending on the kernel $k_{n,l}$.
Note that the bounds of Theorems 2.6 and 2.8 depend on the smallest eigenvalue of the kernel matrix $\mathbf{K}_{\mathcal{X}}$. Tikhonov regularization is a means to control the latter, for the price of introducing a deterministic bias Section 4 addresses the impact of regularization on the deterministic (Proposition 4.5), probabilistic (Corollary 4.9) and total error bounds (Corollary 4.12).
3. $H^{s}$-invariance under the Koopman operator
Since Wendland (and Matérn) kernels generate the Sobolev spaces as their respective native spaces, we study the invariance of Sobolev spaces $H^{s}$ under the stochastic Koopman operator. As those spaces are defined via partial integration, a condition for preserving weak derivatives under Koopman operator action is paramount. The central ingredient for this is the following signed measure associated with the transition kernel $\rho_{x}$. For $g\in L^{\infty}(\Omega)$, define $\nu_{g}:\mathcal{B}_{\Omega}\to\mathbb{R}$ by
$\nu_{g}(\Delta):=\int_{\Omega}\rho_{x}(\Delta)g(x)\,dx.$ (22)
The following lemma links the signed measure $\nu_{g}$ to the action of the dual of the Koopman operator on $L^{1}(\Omega)$ and $L^{2}(\Omega)$. Its proof is given in Section 3.1.
Lemma 3.1.
Assume that ($L^{1}$) holds and consider $\mathcal{K}_{1}:L^{1}(\Omega)\to L^{1}(\Omega)$ as well as $\mathcal{K}_{2}:L^{2}(\Omega)\to L^{2}(\Omega)$. Then, the measure $\nu_{g}$ defined in (22) is absolutely continuous w.r.t. the Lebesgue measure, and its density $\frac{d\nu_{g}}{dx}\in L^{\infty}(\Omega)$ satisfies
$\mathcal{K}_{1}^{*}g=\mathcal{K}_{2}^{*}g=\frac{d\nu_{g}}{dx},\quad\mathrm{for\ all}\quad g\in L^{\infty}(\Omega).$
where $\mathcal{K}_{1}^{\ast}$ and $\mathcal{K}_{2}^{\ast}$ denote the adjoints w.r.t. the respective dual pairing.
In Section 3.1, we state our main result Theorem 3.2 regarding Sobolev space invariance, which turns out to be closely linked to the properties of such densities $\frac{d\nu_{g}}{dx}$. By Proposition 2.4, this RKHS invariance implies boundedness of the Koopman operator, which proves crucial for the error bound provided later in Section 4. Section 3.2 briefly introduces relevant notation and defines integer-order and fractional-order Sobolev spaces. The concluding Section 3.3 is dedicated to the proofs of the results of Section 3.1.
3.1. Conditions for $H^{s}$-invariance and boundedness
The following theorem provides conditions for $H^{s}$-invariance in different settings. While (i) fully characterizes $H^{\ell}$-invariance for integer order $\ell\in\mathbb{N}$, (ii) gives a sufficient condition for $H^{s}$-invariance with real-valued $s\geq 0$. (iii) covers the important special case of a (weakly) differentiable transition density and was presented as Theorem 2.7 in Section 2. The proof is deferred to Section 3.3.
Theorem 3.2 (Koopman invariance of Sobolev spaces).
Assume that the condition ($L^{1}$) holds.
  • (i)
    Let $\ell\in\mathbb{N}_{\geq 1}$. Then $H^{\ell}$ is invariant under $\mathcal{K}$ if and only if there is $C\geq 0$ such that
    $\sum_{|\alpha|\leq\ell}\left\|\frac{d\nu_{D^{\alpha}\varphi}}{dx}\right\|_{H^{-|\alpha|}}\leq C\|\varphi\|_{L^{2}}\qquad\forall\varphi\in C_{c}^{\infty}.$ (23)
  • (ii)
    If (23) holds, then $H^{s}$ is invariant under $\mathcal{K}$ for all $s\in(0,\ell]$.
  • (iii)
    Assume the transition kernel is given by a density $p:\Omega\times\Omega\to[0,\infty)$ such that
    $\rho_{x}(\Delta)=\int_{\Delta}p(x,y)\,dy\quad\mathrm{for\ all}\quad x\in\Omega,\,\Delta\in\mathcal{B}_{\Omega}.$
    If for some constant $C>0$ and $\ell\in\mathbb{N}_{\geq 1}$
    $\sum_{|\alpha|\leq\ell}\left\|D_{1}^{\alpha}\,p(\cdot,\cdot)\right\|_{L^{2}(\Omega\times\Omega)}\leq C$ (24)
    then $H^{s}$ is invariant under $\mathcal{K}$ for all $s\in(0,\ell]$.
Note that part (iii) of Theorem 3.2 is precisely the statement of Theorem 2.7. In the following result, we exemplarily verify its sufficient condition (24) for $H^{s}$-invariance in case of a particular smooth transition density. Similar results can be obtained for higher-order Sobolev spaces or other (truncated) distributions with smooth densities on bounded domains $\Omega$.
Proposition 3.3 (Truncated Normal Distribution).
For $n=1$ and $\Omega=(0,1)\subset\mathbb{R}$, assume $Y\mid\{X=x\}\sim\rho_{x}:=\mathcal{N}_{(0,1)}(x,1)$ where $\mathcal{N}_{(0,1)}(x,1)$ denotes the (doubly) truncated normal distribution on $(0,1)$ with location parameter $x\in(0,1)$ and scale parameter set to $1$. Then, $H^{1}(\Omega)$ is invariant under the associated Koopman operator that acts on $f\in H^{1}(\Omega)$ via $f\in H^{1}\bigl((0,1)\bigr)$ via $\mathcal{K}f(x)=\mathbb{E}[f(Y)\mid X=x]$.
Proof.
The doubly truncated normal distribution $\rho_{x}$ is the distribution of a random variable $Z\sim\mathcal{N}(x,1)$ conditioned on the event $0<Z<1$. To apply part (iii) of Theorem 3.2, we study. the density of $\rho_{x}$. Denote the cumulative distribution function of the standard normal distribution by $F(\xi)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\xi}e^{-t^{2}/2}\,dt$. Then, the distribution of $Y$ conditioned on $\{X=x\}$ with $x\in(0,1)$ has density $p_{x}=p(x,\cdot):(0,1)\to[0,\infty)$ given by
$p(x,y)=\frac{\frac{1}{\sqrt{2\pi}}e^{-(y-x)^{2}/2}}{F(1-x)-F(-x)}=\frac{G(x,y)}{H(x)}$
Numerator and denominator both depend smoothly on $x$, so we can easily compute derivatives with respect to $x$ and bound them. Using $F^{\prime}(\xi)=\frac{1}{\sqrt{2\pi}}e^{-\xi^{2}/2}$, we get
$\partial_{1}p(x,y)=\frac{\partial_{1}G(x,y)H(x)-G(x,y)H^{\prime}(x)}{H(x)^{2}}=G(x,y)\left((y-x)-\frac{H^{\prime}(x)}{H(x)^{2}}\right)$
Note that $0<C_{1}\leq H(x)\leq C_{2}$ since $H(x)=\mathbb{P}(U\in(-x,1-x))$ for some $U\sim\mathcal{N}(0,1)$ and $x\in(0,1)$. Using the fact that $G(x,y)\leq\frac{1}{\sqrt{2\pi}}$ for all $x,y\in(0,1)$ and $|H^{\prime}(x)|\leq\frac{1}{\sqrt{2\pi}}$, we obtain the basic bound
$|\partial_{1}p(x,y)|\leq\frac{1}{\sqrt{2\pi}}\left(1+\frac{1}{\sqrt{2\pi}C_{1}}\right)$
for all $x,y\in(0,1)$. Thus, condition (24) is satisfied with $C=\frac{1}{\sqrt{2\pi}}\left(1+\frac{1}{\sqrt{2\pi}C_{1}}\right)$. By part (iii) of Theorem 3.2, $H^{1}(\Omega)$ is invariant under the associated Koopman operator $\mathcal{K}$. ∎
3.2. Sobolev Spaces
For self-containedness, we briefly recapitulate the notion of Sobolev spaces following [Adams2003]. Readers who are familiar with Sobolev spaces may only skim this section for notation or skip it entirely.
Recall the bounded domain $\Omega\subset\mathbb{R}^{n}$ with Lipschitz boundary and write $C_{c}^{\infty}(\Omega)$ for the space of infinitely differentiable and compactly supported functions on $\Omega$. If the domain is clear from context, we might write $C_{c}^{\infty}=C_{c}^{\infty}(\Omega)$. Given some multi-index $\alpha\in\mathbb{N}_{0}^{n}$, we write $D^{\alpha}\varphi$ for the $\alpha$-th partial derivative of a sufficiently smooth function $\varphi$. Note that $\varphi\in C_{c}^{\infty}(\Omega)$ implies $D^{\alpha}\varphi(x)=0$ for all $x\in\partial\Omega$ and $\alpha\in\mathbb{N}^{n}$ since $\varphi\equiv 0$ in an open neighborhood of the boundary.
Next, recall the notion of the weak derivative of a function: Consider a locally integrable function $f\in L_{\mathrm{loc}}^{1}(\Omega)$; that is, $f$ is integrable on every compact set $K\subset\Omega$, for which $f\in L^{2}(\Omega)$ is clearly sufficient. If there exists a function $g\in L_{\mathrm{loc}}^{1}(\Omega)$ satisfying
$\int_{\Omega}f(x)\,(D^{\alpha}\varphi)(x)\,dx=(-1)^{|\alpha|}\int_{\Omega}g(x)\,\varphi(x)\,dx$ (25)
for every test function $\varphi\in C_{c}^{\infty}(\Omega)$, we call $g$ a $\alpha$-th weak derivative of $f$. The weak derivative is unique up to a set of measure zero, i.e., two weak derivatives of $f$ agree almost everywhere, and needs not be defined pointwise (the elements of $L_{\mathrm{loc}}^{1}(\Omega)$ are equivalence classes, after all). In the case of a “conventionally” partially differentiable function, the weak derivative coincides (almost everywhere) with the respective partial derivative. Hence, in a slight abuse of notation, we write $g=D^{\alpha}f$ for the weak derivative.
Using the notion of a weak derivative, given some $\ell\in\mathbb{N}$ we introduce
$H^{\ell}(\Omega)=W^{\ell,2}(\Omega)=\{f\in L^{2}(\Omega):D^{\alpha}f\in L^{2}(\Omega)~\text{for all}~\alpha\in\mathbb{N}^{n}~\text{with}~|\alpha|\leq\ell\}$
as the space of $L^{2}$-functions whose weak derivatives up to degree $\ell$ exist and lie in $L^{2}$. If the domain $\Omega$ is clear from context, we briefly write $H^{\ell}=H^{\ell}(\Omega)$. Equipped with the norm
$\|f\|_{H^{\ell}}=\left(\sum_{|\alpha|\leq\ell}\|D^{\alpha}f\|_{L^{2}}^{2}\right)^{1/2},$
$H^{\ell}$ is a Banach space. Its dual space with respect to the $L^{2}$-scalar product will be denoted by $H^{-\ell}(\Omega)$ or $H^{-\ell}$. We briefly remind the reader that we have the following chain of inclusions with continuous embeddings $H^{1}\,\subset\,C_{b}\,\subset\,L^{2}\,\subset\,C_{b}^{*}=\mathcal{M}\,\subset\,H^{-1}$, where $C_{b}^{*}=\mathcal{M}$ is the space of all finite regular signed measures on $\Omega$.
Sobolev spaces can be generalized to real-valued regularity order $s\geq 0$. Those fractional Sobolev spaces arise, for example, as reproducing kernel Hilbert spaces of certain kernels. There are two predominant definitions of fractional Sobolev spaces in the literature, either via the Fourier transform or via weak derivatives and interpolation between integer-order Sobolev spaces. For a domain with Lipschitz boundary, as is $\Omega\subset\mathbb{R}^{n}$, those definitions coincide [McLe2000, Theorem 3.30], so we proceed with the more suitable second definition for the remainder of this paper. Given $s\geq 0$, write $s=\lfloor s\rfloor+r$ with $r\in[0,1)$ and define the norm
$\|f\|_{H^{s}(\Omega)}^{2}:=\sum_{|\alpha|\leq\,s}\|D^{\alpha}f\|_{L^{2}(\Omega)}^{2}+\chi_{(0,1)}(r)\int_{\Omega}\int_{\Omega}\frac{\left|D^{\alpha}f(x)-D^{\alpha}f(y)\right|^{2}}{\|x-y\|_{2}^{n+2r}}\,dx\,dy.$ (26)
Under this definition, the Sobolev space $H^{s}(\Omega)=\{f\in L^{2}(\Omega):\|f\|_{H^{s}(\Omega)}<\infty\}$ presents itself as the interpolation space between $H^{\lfloor s\rfloor}(\Omega)$ and $H^{\lceil s\rceil}(\Omega)$.
3.3. Proof of Lemma 3.1 and Theorem 3.2
We start by proving the important auxiliary result from the beginning of this section. While Lemma 3.1 essentially follows from [PhilScha24, Appendix B], we provide a concise proof below for self-containedness.
Proof of Lemma 3.1.
Let $g\in L^{\infty}(\Omega)$ and $\Delta\in\mathcal{B}_{\Omega}$. If ($L^{1}$) holds, the definition of (22) immediately gives
$|\nu_{g}(\Delta)|\,\leq\,\|g\|_{L^{\infty}}\int\rho_{x}(\Delta)\,dx\,\leq\,C\|g\|_{L^{\infty}}|\Delta|.$
Hence, $\nu_{g}$ is absolutely continuous w.r.t. Lebesgue measure with density $\frac{d\nu_{g}}{dx}\in L^{\infty}(\Omega)$. Moreover, for $f\in L^{1}(\Omega)$ and the corresponding Koopman operator $\mathcal{K}_{1}:L^{1}(\Omega)\to L^{1}(\Omega)$, we have
$\langle\mathcal{K}_{1}f,g\rangle_{L^{1},L^{\infty}}=\int\mathcal{K}_{1}f\cdot g\,dx=\int\int f\,d\rho_{x}g(x)\,dx=\int f\,d\nu_{g}=\left\langle f,\frac{d\nu_{g}}{dx}\right\rangle_{L^{1},L^{\infty}}.$
This shows $\mathcal{K}_{1}^{*}g=\frac{d\nu_{g}}{dx}$. Finally, if $f,g\in L^{\infty}(\Omega)$, we have
$\langle\mathcal{K}_{2}^{*}f,g\rangle_{L^{1},L^{\infty}}=\langle\mathcal{K}_{2}^{*}f,g\rangle_{L^{2}}=\langle f,\mathcal{K}_{2}g\rangle_{L^{2}}=\langle f,\mathcal{K}_{1}g\rangle_{L^{1},L^{\infty}}=\langle\mathcal{K}_{1}^{*}f,g\rangle_{L^{1},L^{\infty}}.$
It follows that $\mathcal{K}_{2}^{*}f=\mathcal{K}_{1}^{*}f$. ∎
Building on Lemma 3.1, the remainder of the proof of Theorem 3.2 consists of five steps.
Proof of Theorem 3.2.
Step 1. $H^{1}$-invariance. As the first step, we prove the statement of for $\ell=s=1$.
We start with the sufficiency of invariance in (i): Assume that $H^{1}$ is invariant under $\mathcal{K}$ and fix $\alpha\in\mathbb{N}^{n}$ with $|\alpha|=1$. Then $\mathcal{K}:H^{1}\to H^{1}$ is bounded by Proposition 2.4. Specifically, there exists $C_{1}>0$ such that
$\|\mathcal{K}f\|_{H^{1}}\leq C_{1}\|f\|_{H^{1}}$ (27)
for all $f\in H^{1}$. Moreover, Lemma 3.1 gives $\frac{d\nu_{D^{\alpha}\varphi}}{dx}=\mathcal{K}_{2}^{\ast}(D^{\alpha}\varphi)$, so we study the $H^{-1}$-norm of the latter. For all $f\in H^{1}$ and all $\varphi\in C_{c}^{\infty}$, we have
$|\langle f,\mathcal{K}_{2}^{*}(D^{\alpha}\varphi)\rangle_{H^{1},H^{-1}}|$ $=\left|\langle\mathcal{K}f,D^{\alpha}\varphi\rangle\right|=\left|\langle D^{\alpha}(\mathcal{K}f),\varphi\rangle\right|\,\leq\,\|D^{\alpha}(\mathcal{K}f)\|_{L^{2}}\|\varphi\|_{L^{2}}$ (28)
$\leq\,\|\mathcal{K}f\|_{H^{1}}\|\varphi\|_{L^{2}}\,\leq\,C_{1}\|f\|_{H^{1}}\|\varphi\|_{L^{2}}.$
The first equality follows from the definition of the dual operator, while the second equality leverages the definition of the weak derivative $D^{\alpha}$. The inequalities follow from the Cauchy-Schwarz inequality, the definition of the norm in the Sobolev space $H^{1}$ and (27). The inequality above implies
$\left\|\frac{d\nu_{D^{\alpha}\varphi}}{dx}\right\|_{H^{-1}}=\|\mathcal{K}_{2}^{*}(D^{\alpha}\varphi)\|_{H^{-1}}$ $=\sup\left\{|\langle f,\mathcal{K}_{2}^{*}(D^{\alpha}\varphi)\rangle_{H^{1},H^{-1}}|:f\in H^{1},\,\|f\|_{H^{1}}=1\right\}\,$
$\leq\,C_{1}\|\varphi\|_{L^{2}}.$
for every $|\alpha|=1$. For $|\alpha|=0$, we have $H^{0}=L^{2}$ and can directly apply the Cauchy-Schwarz inequality and Lemma 2.1 to get $\left|\langle\mathcal{K}f,\varphi\rangle\right|\leq\|\mathcal{K}f\|_{L^{2}}\|\varphi\|_{L^{2}}\leq C_{2}\|f\|_{L^{2}}\|\varphi\|_{L^{2}}$ and hence
$\left\|\frac{d\nu_{\varphi}}{dx}\right\|_{H^{-1}}\leq C_{2}\|\varphi\|_{L^{2}}$
for some $C_{2}\geq 0$. Since those inequalities hold for all $\varphi\in C_{c}^{\infty}$, (23) follows for $\ell=1$ with appropriately defined $C\geq 0$.
Conversely, suppose that (23) holds for all $\varphi\in C_{c}^{\infty}$. By Lemma 3.1, this gives
$\left\|\mathcal{K}_{2}^{\ast}(D^{\alpha}\varphi)\right\|_{H^{-1}}=\left\|\frac{d\nu_{D^{\alpha}\varphi}}{dx}\right\|_{H^{-1}}\leq C\|\varphi\|_{L^{2}}$
for every $|\alpha|=1$. Hence, $|\langle\mathcal{K}f,D^{\alpha}\varphi\rangle|=|\langle f,K_{2}^{\ast}(D^{\alpha}\varphi)\rangle|\leq C_{\alpha}\|f\|_{H^{1}}\|\varphi\|_{L^{2}}$ holds for all $f\in H^{1}$ and all $\varphi\in C_{c}^{\infty}$. Therefore, for fixed $f\in H^{1}$, the functional $\varphi\mapsto\langle\mathcal{K}f,D^{\alpha}\varphi\rangle$ is bounded on $C_{c}^{\infty}$ and can thus be extended to a bounded functional $F_{f}$ on $L^{2}$. By Riesz’ representation theorem, $F_{f}\varphi=\langle Lf,\varphi\rangle$ for all $\varphi\in L^{2}$ with some $Lf\in L^{2}$. Equating those expressions gives $\langle\mathcal{K}f,D^{\alpha}\varphi\rangle=F_{f}\varphi=\langle Lf,\varphi\rangle$ for all $f\in H^{1}$ and all $\varphi\in C_{c}^{\infty}$. This shows that $\mathcal{K}f$ has the weak derivative $Lf=D^{\alpha}(\mathcal{K}f)\in L^{2}$ for all $f\in H^{1}$ and all $|\alpha|=1$, hence $\mathcal{K}f\in H^{1}$ and $H^{1}$ is invariant under $\mathcal{K}$. This shows the equivalence from (i) in Theorem 3.2.
Step 2: $H^{\ell}$-invariance, $\ell\in\mathbb{N}$. The characterization from step 1 is easily generalized to Sobolev spaces of higher smoothness. Fix $\ell\in\mathbb{N}$. By definition, we have $f\in H^{\ell}$ if and only if $D^{\alpha}f\in L^{2}$ for every $\alpha\in\mathbb{N}^{n}$ with $|\alpha|\leq\ell$. For the sufficiency, assume the invariance of $H^{\ell}$, i.e., $\mathcal{K}f\in H^{\ell}$ for every $f\in H^{\ell}$. Again, this gives the boundedness of $\mathcal{K}:H^{\ell}\to H^{\ell}$ by Proposition 2.4, and for all $\alpha\in\mathbb{N}^{n}$ with $|\alpha|\leq\ell$ we have $D^{\alpha}(\mathcal{K}f)\in L^{2}$. By the definition of the weak derivative (25)
$\left|\left\langle\mathcal{K}f,D^{\alpha}\varphi\right\rangle\right|=\left|\left\langle D^{\alpha}(\mathcal{K}f),\varphi\right\rangle\right|$
for all $f\in H^{\ell}$ and $\varphi\in C_{c}^{\infty}$. Replacing $H^{1}$ by $H^{\ell}$ in the arguments from step 1 from (27) onward, we get (23) for $\ell\in\mathbb{N}$.
Conversely, in analogy to step 1, (23) gives the boundedness of the functional $\varphi\mapsto\langle\mathcal{K}f,D^{\alpha}\varphi\rangle$ for fixed $f\in H^{\ell}$ and every $|\alpha|\leq\ell$ on $C_{c}^{\infty}$ and its extension on $L^{2}$. By Riesz’ representation theorem, this functional is of form $\langle Lf,\varphi\rangle$ with $Lf\in L^{2}$. We obtain the weak derivative $D^{\alpha}(\mathcal{K}f)=(-1)^{|\alpha|}Lf$ for all $|\alpha|\leq\ell$ and hence the invariance of $H^{\ell}$ under $\mathcal{K}$.
Thus, the invariance of $H^{\ell}$ under $\mathcal{K}$ and (23) are equivalent, concluding the proof of part (i) of Theorem 3.2.
In the next step, we use an interpolation argument to deduce a sufficient condition for fractional Sobolev space invariance.
Step 3: $H^{s}$-invariance, $s\in(0,\ell]$. By [BrenScot2008, Theorem 14.2.3], for real-valued $s\in(0,\ell]$, the space $H^{s}$ coincides with the real interpolation space
$H^{s}(\Omega)=[L^{2}(\Omega),H^{\ell}(\Omega)]_{\theta},\quad\theta=s/\ell\in(0,1)$
with equivalent norms. Under condition ($L^{1}$), we have the invariance of $L^{2}(\Omega)$ under $\mathcal{K}$. Let (23) hold for $\ell\in\mathbb{N}$. Then $H^{\ell}(\Omega)$ is also invariant under $\mathcal{K}$ by step 2. [Tart2007, Theorem 22.3] now implies that the Koopman operator
$\mathcal{K}:[L^{2}(\Omega),H^{\ell}(\Omega)]_{\theta}\to[L^{2}(\Omega),H^{\ell}(\Omega)]_{\theta}$
mapping from this interpolation space is well-defined and bounded. Thus, $H^{s}(\Omega)$ in indeed invariant under $\mathcal{K}$, and part (ii) of Theorem 3.2 is proven.
The final two steps of the proof deal with the special case that the transition kernel is given by a differentiable transition density, in which case the Radon-Nikodym derivative $\frac{d\nu_{g}}{dx}$ can be explicitly calculated.
Step 4: Deriving an explicit Radon-Nikodym density. Assume that the transition kernel $\rho_{x}$ admits a density w.r.t. the Lebesgue measure for every $x\in\Omega$, i.e., there exists some $p_{x}:\Omega\to[0,\infty)$ such that $\rho_{x}(\Delta)=\int_{\Delta}p_{x}(y)\,dy$ for every $\Delta\in\mathcal{B}_{\Omega}$. Consider the family of densities as mapping $p:\Omega\times\Omega\to[0,\infty),p(x,y)=p_{x}(y)$ and write $d\rho_{x}(y)=p(x,y)\,dy$. This gives $\rho_{x}(\Delta)=\mathbb{P}(Y\in\Delta\mid X=x)=\int_{\Omega}\chi{}_{\Delta}(y)\,p(x,y)\,dy$ and for $g\in L^{\infty}(\Omega)$, Fubini’s theorem yields
$\nu_{g}(\Delta)=\int_{\Omega}\biggl(\int_{\Omega}\chi_{\Delta}(y)p(x,y)\,dy\biggr)\,g(x)\,dx=\int_{\Omega}\chi_{\Delta}(y)\biggl(\int_{\Omega}p(x,y)g(x)\,dx\biggr)\,dy,$
so the Radon-Nikodym density $z_{g}=\frac{d\nu_{g}}{dx}$ of $\nu_{g}$ takes the form
$z_{g}(y)=\int_{\Omega}p(x,y)g(x)\,dx.$ (29)
Step 5: Sufficient condition via density derivatives. To apply part (ii) of Theorem 3.2, we take $g=D^{\alpha}\varphi$ with $\alpha\in\mathbb{N}^{n}$, $|\alpha|\leq\ell\in\mathbb{N}$, and $\varphi\in C_{c}^{\infty}$ in (29). Integration by parts yields the expression
$z_{D^{\alpha}\varphi}(y)=\int_{\Omega}p(x,y)(D^{\alpha}\varphi)(x)\,dx=(-1)^{|\alpha|}\int_{\Omega}D_{1}^{\alpha}\,p(x,y)\varphi(x)\,dx$
for the Radon-Nikodym density of $\nu_{D^{\alpha}\varphi}$, where $D_{1}^{\alpha}$ denotes the $\alpha$-th weak derivative with respect to the first argument. Now, to compute $\left\|\frac{d\nu_{D^{\alpha}\varphi}}{dx}\right\|_{H^{-\ell}}=\|z_{D^{\alpha}\varphi}\|_{H^{-\ell}}$, we take $f\in H^{\ell}$ and calculate
$\left|\left\langle f,z_{D^{\alpha}\varphi}\right\rangle_{H^{\ell},H^{-\ell}}\right|$ $=\left|\int_{\Omega}f(y)z_{D^{\alpha}\varphi}(y)\,dy\right|$
$=\left|\int_{\Omega}f(y)\int_{\Omega}D_{1}^{\alpha}\,p(x,y)\varphi(x)\,dx\,dy\right|$
$\leq\|f\|_{L^{2}}\left\|\int_{\Omega}D_{1}^{\alpha}\,p(x,\cdot)\,\varphi(x)\,dx\right\|_{L^{2}}$
$\leq\|f\|_{L^{2}}\|\varphi\|_{L^{2}}\|D_{1}^{\alpha}\,p(\cdot,\cdot)\|_{L^{2}(\Omega\times\Omega)}\leq\|f\|_{H^{\ell}}\,\|\varphi\|_{L^{2}}\|D_{1}^{\alpha}\,p(\cdot,\cdot)\|_{L^{2}(\Omega\times\Omega)}.$
In this, we applied the Cauchy-Schwarz inequality twice and then used that $\|f\|_{L^{2}}\leq\|f\|_{H^{\ell}}$. Recall that
$\left\|z_{D^{\alpha}\varphi}\right\|_{H^{-\ell}}=\sup\left\{|\langle f,z_{D^{\alpha}\varphi}\rangle_{H^{\ell},H^{-\ell}}|:f\in H^{\ell},\,\|f\|_{H^{\ell}}=1\right\}\,\leq\,\|D_{1}^{\alpha}\,p(\cdot,\cdot)\|_{L^{2}(\Omega\times\Omega)}\,\|\varphi\|_{L^{2}}.$
Thus, if (24) holds for some $C>0$, we have
$\sum_{|\alpha|\leq\ell}\left\|\frac{d\nu_{D^{\alpha}\varphi}}{dx}\right\|_{H^{-\ell}}\leq\sum_{|\alpha|\leq\ell}\|D_{1}^{\alpha}\,p(\cdot,\cdot)\|_{L^{2}(\Omega\times\Omega)}\,\|\varphi\|_{L^{2}}\leq C\,\|\varphi\|_{L^{2}}.$
By part (ii) of Theorem 3.2, this implies that $H^{s}$ is invariant under $\mathcal{K}$ for all $s\in(0,\ell]$. This is precisely the statement of part (iii) of Theorem 3.2, concluding the proof. ∎
4. Error bounds for approximations of Koopman operators of stochastic systems
This section is concerned with deriving error bounds for the approximation of the stochastic Koopman operator defined by
$(\mathcal{K}f)(x):=\mathbb{E}\big[f(Y)\,|\,X=x].$ (30)
As our proof strategy reveals, we will rely on the following standing assumption.
Assumption 4.1.
Let $k:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ be a continuous strictly positive-definite symmetric kernel and $\mathbb{H}$ the induced reproducing kernel Hilbert space on $\Omega$ with $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b}(\Omega),C_{b}(\Omega))$}.
To verify especially the condition $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ in the assumption above, we refer to the sufficient conditions of the previous Section 3, where Theorem 3.2 guarantees the invariance of $\mathbb{H}$ since Sobolev spaces $\mathbb{H}=H^{s}$ are induced by Wendland kernels as RKHS [KohnPhil24, Theorem 4.1].
We start by constructing the kernel-based approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$ of the stochastic Koopman operator from (18) more thoroughly. In an intermediate step, we derive a kernel EDMD approximant $\widehat{\mathcal{K}}$, significantly building upon the analogue from [KohnPhil24] in the deterministic setting. This allows to split the approximation error as follows:
[…]
The remainder of this section works towards the proof of Theorem 2.6 by analyzing both errors independently. The intermediary approximation $\widehat{\mathcal{K}}$ is employed in Section 4.1 to prove a bound on the deterministic projection error, see Theorem 4.2. Therein, we appeal to similar arguments as [KohnPhil24], where the invariance of a reproducing kernel Hilbert space under the Koopman operator is assumed as in our Assumption 4.1. Since $\widehat{\mathcal{K}}$ requires the exact knowledge of propagated observables under the stochastic Koopman operator, i.e., conditional expectation, Section 4.2 proposes the random, but computable approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$. The approximation of the “exact”, but not computable $\widehat{\mathcal{K}}$ introduces a probabilistic error that is dealt with in Theorem 4.6.
4.1. Deterministic approximation step: interpolation
The first (deterministic) approximation mimicks the approach of [KohnPhil24]. However, this approximation will not be useful in practice as it involves expectation values and thus is not computable, in general. Therefore, we merely use it as a theoretical tool for an error bound of a computable estimator which we shall propose in the next subsection.
Let $\mathcal{X}=\{x_{1},\ldots,x_{d}\}\subset\Omega$ be a set of points. As described in Section 2.3, we define a kernel-based approximation by
$\widehat{\mathcal{K}}=S_{\mathcal{X}}\mathcal{K}S_{\mathcal{X}},$ (31)
where $S_{\mathcal{X}}$ is the $\mathbb{H}$-orthogonal projection onto
$V_{\mathcal{X}}=\operatorname{span}\{\Phi_{x_{i}}:i\in\{1,\dots,d\}\}\subset\mathbb{H}$
with the canonical features $\Phi_{x_{i}}=k(x_{i},\cdot)\in\mathbb{H}$. With the notation from Section 2.3, recall the explicit projection formula (10) given by $S_{\mathcal{X}}g=g_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}$ for some $g\in\mathbb{H}$.
Hence, application of the abstract operator defined in (31) may explicitly be computed via
$\widehat{\mathcal{K}}f=S_{\mathcal{X}}\mathcal{K}S_{\mathcal{X}}f=(\mathcal{K}S_{\mathcal{X}}f)_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}=\sum_{i=1}^{d}\big[\mathbf{K}_{\mathcal{X}}^{-1}\big(\underbrace{\mathbb{E}[(S_{\mathcal{X}}f)(Y)|X=\cdot]}_{\mathcal{K}S_{\mathcal{X}}f}\big)_{\mathcal{X}}\big]_{i}\Phi_{x_{i}}.$
Using linearity of the expected value and the projection formula (10), we get, for each $j\in\{1,\dots,d\}$,
$\mathbb{E}[S_{\mathcal{X}}f(Y)|X=x_{j}]=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbb{E}[\mathbf{k}_{\mathcal{X}}(Y)|X=x_{j}]=\mathbb{E}[\mathbf{k}_{\mathcal{X}}^{\top}(Y)|X=x_{j}]\mathbf{K}_{\mathcal{X}}^{-1}f_{\mathcal{X}}.$
Hence, defining the coefficients
$\alpha_{\widehat{\mathcal{K}}f}=\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}f_{\mathcal{X}},\qquad\mathrm{where}\qquad(\mathbf{K}_{\mathcal{X},\mathcal{Y}})_{i,j}=\mathbb{E}(k(Y,x_{i})|X=x_{j}),$
yields the approximant
$\widehat{\mathcal{K}}f=\sum_{i=1}^{d}(\alpha_{\widehat{K}f})_{i}\Phi_{x_{i}}=\alpha_{\widehat{K}f}^{\top}\mathbf{k}_{\mathcal{X}}=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}.$ (32)
In what follows, we shall assume that the probability kernel has the continuity-like property
$\forall\,x_{0}\in\Omega\;\forall\,\Delta\in\mathcal{B}_{\Omega}(x_{0}):\quad\rho_{x}(\Delta)\to\rho_{x_{0}}(\Delta)\quad\text{ as $x\to x_{0}$},$ (33)
where $\mathcal{B}_{\Omega}(x)=\{\Delta\in\mathcal{B}_{\Omega}:\rho_{x}(\partial\Delta)=0\}$ for $x\in\Omega$, where $\partial\Delta$ denotes the boundary of the Borel-set $\Delta$. Note that this holds in the deterministic case where $\rho_{x}=\delta_{F(x)}$ with continuous $F:\Omega\to\Omega$: Then, $\rho_{x_{0}}(\partial\Delta)=\delta_{F(x_{0})}(\partial\Delta)=0$ is equivalent to $F(x_{0})\notin\partial\Delta$, so $F(x_{0})$ lies either in the interior or the exterior of $\Delta$, and continuity of $F$ gives the convergence. By the Portmanteau theorem [Klen2008, Theorem 13.16], the condition (33) is equivalent to the fact that $\mathcal{K}f$ is continuous for each $f\in C_{b}(\Omega)$. Hence, if (33) is satisfied, the Koopman operator is a contraction from $C_{b}(\Omega)$ into itself as for $f\in C_{b}(\Omega)$,
$\|\mathcal{K}f\|_{\infty}=\sup_{x\in\Omega}\left|\int_{\Omega}f\,d\rho_{x}\right|\leq\sup_{x\in\Omega}\int_{\Omega}\|f\|_{\infty}\,d\rho_{x}=\|f\|_{\infty}.$
Combining the sufficient conditions of Theorem 3.2 with (33), Assumption 4.1 holds, under which the approximant $\widehat{\mathcal{K}}$ from (32) satisfies the following error bound.
Theorem 4.2 (Deterministic error bound).
Under Assumption 4.1, i.e. $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b},C_{b})$, the approximant $\widehat{\mathcal{K}}$ from (32) admits the error bound
$\big\|\mathcal{K}-\widehat{\mathcal{K}}\big\|_{\mathbb{H}\to C_{b}}\,\leq\,(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}.$ (34)
Proof.
The proof is a particular case of the proof of [KohnPhil24, Theorem 3.4] and we repeat it here for the sake of clarity. As $\mathcal{K}-\widehat{\mathcal{K}}=\mathcal{K}-S_{\mathcal{X}}\mathcal{K}S_{\mathcal{X}}=\mathcal{K}(\mathcal{I}-S_{\mathcal{X}})+(\mathcal{I}-S_{\mathcal{X}})\mathcal{K}S_{\mathcal{X}}$, we get
$\big\|\mathcal{K}-\widehat{\mathcal{K}}\big\|_{\mathbb{H}\to C_{b}}$ $\leq\|\mathcal{K}(\mathcal{I}-S_{\mathcal{X}})\|_{\mathbb{H}\to C_{b}}+\|(\mathcal{I}-S_{\mathcal{X}})\mathcal{K}S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}$
$\leq\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}(\|\mathcal{K}\|_{C_{b}\to C_{b}}+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\|S_{\mathcal{X}}\|_{\mathbb{H}\to\mathbb{H}}).$
Under the assumption $\mathcal{K}\in L(C_{b},C_{b})$, $\mathcal{K}$ is a contraction, and for constant $f\equiv c\in\Omega$
$\mathcal{K}f(x)=\mathbb{E}[f(Y)\mid X=x]=\mathbb{E}[c\mid X=x]=c=f(x)\quad\forall x\in\Omega,$
so $\|\mathcal{K}\|_{C_{b}\to C_{b}}=1$. Since $S_{\mathcal{X}}$ is a projection, we also have $\|S_{\mathcal{X}}\|_{\mathbb{H}\to\mathbb{H}}=1$, and the claim follows. ∎
By Proposition 2.4, the invariance assumption $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ implies that the operator norm $\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}$ is bounded. Thus, it is in order to focus on the projection error $\mathcal{I}-S_{\mathcal{X}}$ in dependence of the point set $\mathcal{X}$. The projection error in (34) attains a particularly intuitive form when considering Sobolev spaces as RKHS. As stated before, Wendland and Matérn kernels both generate such Sobolev spaces, allowing for the application of results of approximation theory ([Wend04, Theorem 11.13]) to further specify Theorem 4.2 in these cases. Under the appropriate sufficient condition from Section 3, the order of convergence can be quantified in terms of the smoothness of the underlying kernels. We exemplarily state the result for Wendland kernels below.
Corollary 4.3 (Deterministic error bound in Wendland native spaces).
Let $l\in\mathbb{N}$ and define $\sigma_{n,l}=\frac{n+1}{2}+l$. Consider the Wendland kernel $k_{n,l}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ that generates the Sobolev space $H^{\sigma_{n,l}}(\Omega)$ as reproducing kernel Hilbert space with equivalence between the native norm $\|\cdot\|_{\mathbb{H}}$ and the classical Sobolev norm $\|\cdot\|_{H^{\sigma_{n,l}}}$. Assume that $\mathcal{K}\in L(C_{b},C_{b})$ and that (23) is satisfied for some $\ell\in\mathbb{N}$ with $\ell>\sigma_{n,l}$. Then, the Koopman approximation $\widehat{\mathcal{K}}$ from (31) satisfies
$\|\mathcal{K}-\widehat{\mathcal{K}}\|_{\mathbb{H}\to C_{b}}\leq C(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot h_{\mathcal{X}}^{l+1/2}$ (35)
with some constant $C>0$ and the fill distance
$h_{\mathcal{X}}=\sup_{x\in\Omega}\,\min_{i\in\{1,\dots,d\}}\|x-x_{i}\|_{2}.$
Proof.
Under (23), the Sobolev space $H^{\sigma_{n,l}}$ is invariant under the Koopman operator $\mathcal{K}$. By [Wend04, Theorem 11.17] or the restatement [KohnPhil24, Theorem 5.1], we may bound the projection error by
$\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}\leq Ch_{\mathcal{X}}^{l+1/2}.$
Theorem 4.2 now yields (35). ∎
Remark 4.4.
A similar result for Matérn kernels can be readily obtained; we state it in Appendix C. For equal smoothness parameters, the bound on the projection error in the case of Wendland kernels gains an extra $1/2$ in the exponent of the fill distance $h_{\mathcal{X}}$ compared to Matérn kernels. This is because the locally supported Wendland kernels allow for a bound on the derivatives that scales with the fill distance, while the Matérn kernels are globally supported and stationary. In other words, the compact support introduces boundary effects that do not play a role for the global Matérn kernels.
For a large data set $\mathcal{X}$, the kernel matrix $K_{\mathcal{X}}$ is typically badly conditioned, leading to numerical instabilities. To alleviate this, one might regularize the kernel-based interpolation problem as in [BoldPhil24] to get $R_{\mathcal{X}}^{\lambda}f:={\arg\min}_{{g\in V_{\mathcal{X}}}}\sum_{i=1}^{d}\left|f(x_{i})-g(x_{i})\right|^{2}+\lambda\|g\|_{\mathbb{H}}^{2}$ with regularization parameter $\lambda\geq 0$ and solution operator $R_{\mathcal{X}}^{\lambda}$. The linear solution operator satisfies $R_{\mathcal{X}}^{\lambda}f=f_{\mathcal{X}}^{\top}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}$} for $f\in\mathbb{H}$, so for $\lambda=0$ we recover $S_{\mathcal{X}}$, but for $\lambda>0$ $R_{\mathcal{X}}^{\lambda}:\mathbb{H}\to\mathbb{H}$ is no longer a projection. However, as is shown in [BoldPhil24, Appendix A], in analogy to (31) we can still define the approximant
$\widehat{\mathcal{K}}_{\lambda}:=R_{\mathcal{X}}^{\lambda}\mathcal{K}R_{\mathcal{X}}^{\lambda}$ (36)
that acts on $f\in\mathbb{H}$ via
$\widehat{\mathcal{K}}_{\lambda}f=f_{\mathcal{X}}^{\top}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}.$ (37)
For this regularized approximant, we get a similar error bound as in Theorem 4.2.
Proposition 4.5 (Regularized deterministic error bound).
Under Assumption 4.1, the regularized approximant (37) with $\lambda\geq 0$ satisfies
$\bigl\|\mathcal{K}-\widehat{\mathcal{K}}_{\lambda}\bigr\|_{\mathbb{H}\to C_{b}}\leq\left(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\right)\|\mathcal{I}-R_{\mathcal{X}}^{\lambda}\|_{\mathbb{H}\to C_{b}}.$ (38)
Specifically, for the RKHS $\mathbb{H}=H^{\sigma_{n,l}}$ generated by the Wendland kernel $k_{n,l}$, where $\sigma_{n,l}=\frac{n+1}{2}+l$, the deterministic error admits the bound
$\|(\mathcal{K}-\widehat{\mathcal{K}}_{\lambda})f\|_{\infty}\leq C\left(h_{\mathcal{X}}^{l+1/2}+\lambda^{1/2}\right)\|f\|_{\mathbb{H}}.$ (39)
Proof.
The proof of (38) follows the lines of the proof of Theorem 4.2 and uses the fact that $\bigl\|R_{\mathcal{X}}^{\lambda}\bigr\|_{\mathbb{H}\to\mathbb{H}}\leq 1$ as shown in [BoldPhil24, Proposition B.1 (iii)]. The special case covered in (39) follows from [BoldPhil24, Theorem 2.4]. ∎
The second part of Proposition 4.5 indicates the bias introduced by the regularization parameter $\lambda$. For decreasing fill distance, the regularization bias becomes dominant in the error bound. As a tradeoff, regularization allows for a control on the probabilistic part of the error by bounding the smallest eigenvalue of $\mathbf{K}_{\mathcal{X}}$ from below, see Corollary 4.9 in the following section.
4.2. Probabilistic approximation step: estimating the expected value
In practice, we can not form $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$, as we do not have access to the expected values. For this reason, we introduce a random matrix ${\mathbf{K}}^{\mathrm{MC}}_{\mathrm{prop}}$, whose entries are Monte Carlo approximations of the expected values in $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$. We denote the $j$-th column of $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$ by $(\mathbf{K}_{\mathcal{X},\mathcal{Y}})_{j}$. The $j$-th column of $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ is then defined by
$\big(\mathbf{K}^{\mathrm{MC}}_{\mathcal{X},\mathcal{Y}}\big)_{j}:=\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}\mathbf{k}_{\mathcal{X}}(y_{l,j})\approx\mathbb{E}[\mathbf{k}_{\mathcal{X}}(Y)|X=x_{j}],$ (40)
where $y_{l,j}$, $l\in\{1,\dots,m_{j}\}$, are drawn i.i.d. from $\rho(x_{j},\cdot)$, $j\in\{1,\dots,d\}$.
This leads to a computable approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$ which is analogously defined as $\widehat{\mathcal{K}}$ in (32) with the only difference that we replace $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$ with $\mathbf{K}^{\mathrm{MC}}_{\mathcal{X},\mathcal{Y}}$ as defined in (40), that is
$\widehat{\mathcal{K}}^{\mathrm{MC}}f=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}.$ (41)
Decomposing the full error by inserting the auxiliary approximation $\widehat{\mathcal{K}}$, i.e,
[…]
we are left with bounding the second term, after analyzing the deterministic first term in Section 4.1. Due to the random sampling employed in $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$, this error is probabilistic. In the remainder of this part, we derive a probabilistic bound on the error
$(\widehat{\mathcal{K}}-\widehat{\mathcal{K}}^{\mathrm{MC}})f=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}(\mathbf{K}_{\mathcal{X},\mathcal{Y}}-\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}})\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}.$
For this, let us define the vectors
$g^{\top}:=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}},\qquad(g^{\mathrm{MC}})^{\top}=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}.$
Via Hölder’s inequality, we rewrite the probabilistic error term in terms of those vectors. Subsequently, we bound the maximum norm of the Monte Carlo estimation error between $g$ and $g^{\mathrm{MC}}$ via the Hoeffding inequality.
Theorem 4.6 (Monte Carlo error bound).
Let $\varepsilon>0$ and choose $m_{j}=m\in\mathbb{N}$ i.i.d. samples from every $\rho_{x_{j}}$, $j=1,\dots,d$. Then, for all $f\in\mathbb{H}$, the Koopman approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$ from (41) satisfies
$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|k\|_{\infty}^{1/2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}\|f\|_{\mathbb{H}}$ (42)
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right)$.
Proof.
For all $x\in\Omega$, we may interpret this as a perturbed interpolation problem, where the perturbation of the function values stems from approximation of the expected value:
$|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f(x)|=|(g-g^{\mathrm{MC}})^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)|\leq\|g-g^{\mathrm{MC}}\|_{\infty}\,\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}.$
As shown in the proof of [BoldPhil24, Theorem 4.2], the deterministic second factor is bounded by $\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}\leq\|k\|_{\infty}^{1/2}\cdot\big(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big)^{1/2}$, where $\mathbbm{1}=\{v\in\mathbb{R}^{d}:|v_{i}|=1\}$. Moreover,
$\|g-g^{\mathrm{MC}}\|_{\infty}=\max_{x_{j}\in\mathcal{X}}\left|\mathbb{E}[(S_{\mathcal{X}}f)(Y)|X=x_{j}]-\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}(S_{\mathcal{X}}f)(y_{l,j})\right|.$
Setting $\mathbb{E}_{j}:=g_{j}=\mathbb{E}[(S_{\mathcal{X}}f)(Y)|X=x_{j}]$, we obtain $\mathbb{E}[(S_{\mathcal{X}}f)(y_{l,j})]=\int(S_{\mathcal{X}}f)(y)\,\rho(x_{j},dy)=(\mathcal{K}S_{\mathcal{X}}f)(x_{j})=\mathbb{E}_{j}$. Moreover,
$|(S_{\mathcal{X}}f)(y)|=|\langle S_{\mathcal{X}}f,\Phi_{y}\rangle_{\mathbb{H}}|\,\leq\,\|S_{\mathcal{X}}f\|_{H}\|\Phi_{y}\|_{\mathbb{H}}\,\leq\,\|k\|_{\infty}^{1/2}\|f\|_{\mathbb{H}}$
for all $y\in\Omega$, where the last inequality follows from $\|S_{\mathcal{X}}\|_{H\to H}=1$ since $S_{\mathcal{X}}$ is a projection. Hence, the random variables $(S_{\mathcal{X}}f)(y_{l,j})$ are bounded and i.i.d., so that we may apply Hoeffding’s inequality:
$\mathbb{P}\left(\left|\mathbb{E}_{j}-\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}(S_{\mathcal{X}}f)(y_{l,j})\right|\,\geq\,\varepsilon\|f\|_{H}\right)$ $=\mathbb{P}\left(\left|\sum_{l=1}^{m_{j}}\big[(S_{\mathcal{X}}f)(y_{l,j})-\mathbb{E}_{j}\big]\right|\,\geq\,\varepsilon m_{j}\|f\|_{H}\right)$
$\leq 2\exp\left(-\frac{2m_{j}^{2}\varepsilon^{2}\|f\|_{H}^{2}}{4m_{j}\|k\|_{\infty}\|f\|_{H}^{2}}\right)=2\exp\left(-\frac{m_{j}\varepsilon^{2}}{2\|k\|_{\infty}}\right).$
Plugging this entry-wise result into a union bound, we readily obtain:
$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{\infty}\geq\varepsilon\|f\|_{H})$ $=\mathbb{P}\left(\exists j\in\{1,\dots,d\}:\left|\mathbb{E}_{j}-\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}(S_{\mathcal{X}}f)(y_{l,j})\right|\,\geq\,\varepsilon\|f\|_{H}\right)$
$\leq\sum_{j=1}^{d}\mathbb{P}\left(\left|\mathbb{E}_{j}-\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}(S_{\mathcal{X}}f)(y_{l,j})\right|\,\geq\,\varepsilon\|f\|_{H}\right)$
$\leq 2\sum_{j=1}^{d}\exp\left(-\frac{m_{j}\varepsilon^{2}}{2\|k\|_{\infty}}\right).$
Hence, if $m_{j}=m$ for all $j\in\{1,\dots,d\}$, then
$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{\infty}\geq\varepsilon\|f\|_{H})$ $\leq 2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right).$
In total, for every $f\in H$, with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right)$, we obtain
$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|k\|_{\infty}^{1/2}\cdot\big(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big)^{1/2}\|f\|_{H},$
i.e., (42). ∎
Note that in the bound (42) we have $\|k\|_{\infty}<\infty$ by Assumption 4.1, since the kernel $k$ is continuous on $\mathbb{R}^{n}\times\mathbb{R}^{n}$ and hence bounded on the bounded open set $\Omega\times\Omega$.
Remark 4.7 (Alternative probabilistic bound).
Alternatively, one might use Hölder’s inequality with $p=q=2$:
$|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f(x)|\leq\|g-g^{\mathrm{MC}}\|_{2}\,\|K_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{2}.$
In contrast to Theorem 4.6, which required bounding the maximum norm, this approach necessitates a bound on $\|g-g^{\mathrm{MC}}\|_{2}$ with high probability. Applying the Hilbert-space Hoeffding inequality circumvents the union bound used in the proof of Theorem 4.6 to directly get
$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{2}\geq\varepsilon\|f\|_{H})\leq 2\exp\left(-\frac{m\varepsilon^{2}}{8d\|k\|_{\infty}}\right).$ (43)
By the $\ell^{1}$-$\ell^{2}$-inequality, the deterministic second term is smaller than the one in Theorem 4.6, i.e., $\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{2}\leq\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}$. We refer the reader to Appendix A for detailed calculations. In total, with probability at least $1-2\exp\left(-\frac{m\varepsilon^{2}}{8d\|k\|_{\infty}}\right)$, the bound
$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{2}\,\|f\|_{H}.$
holds. However, comparing the error probabilities, we notice that in Theorem 4.6 the factor $d$ appears in front of the exponential term, while in (43), the factor $d$ appears inside the exponential term. From a practical standpoint, the former is preferable, since we only have to increase the number of Monte Carlo samples $m$ by a logarithmic factor in $d$ to achieve the same error probability. In the latter case, the growth of $m$ would have to be linear in $d$.
Remark 4.8 (Impact of kernel matrix eigenvalues).
The bound
$\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}\leq\|k\|_{\infty}^{1/2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$
from the proof of Theorem 4.6 warrants further discussion. While the second term might prove sharper, a more explicit upper bound involving the Rayleigh quotient can be calculated in the following way: As a first step, apply the $\ell^{1}$-$\ell^{2}$-norm inequality to get
$\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}^{2}\leq d\cdot\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{2}^{2}$ $=d\cdot\mathbf{k}_{\mathcal{X}}(x)^{\top}\mathbf{K}_{\mathcal{X}}^{-2}\,\mathbf{k}_{\mathcal{X}}(x)$
$\leq d\cdot\lambda_{\mathrm{max}}(\mathbf{K}_{\mathcal{X}}^{-1})\,\mathbf{k}_{\mathcal{X}}(x)^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{k}_{\mathcal{X}}(x).$
From the positive definiteness of the kernel $k$, we get the positive semidefiniteness of the augmented kernel matrix $\mathbf{K}_{\mathcal{X}\cup\{x\}}=\begin{pmatrix}\mathbf{K}_{\mathcal{X}}&\mathbf{k}_{\mathcal{X}}\\ \mathbf{k}_{\mathcal{X}}^{\top}&k(x,x)\end{pmatrix}$ with positive semidefinite (thus, non-negative) Schur complement $k(x,x)-\mathbf{k}_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}\geq 0$. With $\lambda_{\mathrm{max}}(\mathbf{K}_{\mathcal{X}}^{-1})=\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})^{-1}$, where $\lambda_{\max}$ and $\lambda_{\min}$ are the largest and smallest eigenvalue of the respective matrices, we get
$\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}^{2}\leq k(x,x)\cdot\frac{d}{\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})}.$
The term in the denominator usually decays rapidly as the condition of the kernel matrix gets worse. If our kernel is given by a radial basis function (RBF) $B:\mathbb{R}^{n}\to\mathbb{R}$, $k(x,x)=B(0)$ is constant, so
$\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}^{2}\leq B(0)\frac{d}{\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})}.$
For example, for the Wendland kernel $k_{n,l}$ with RBF $B_{n,l}$, we have $B_{n,l}(0)=\frac{1}{2^{l}}\binom{\lfloor n/2\rfloor+2l+1}{l}$, while for Matérn kernels $k_{\nu}$ with RBF $\mathcal{M}_{\nu}$, we have $\mathcal{M}_{\nu}(0)=1$.
To balance the decay of the eigenvalue $\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})$ to zero, the computable Monte Carlo approximation (41) also lends itself to a Tikhonov regularization in style of (36). Hence, we define the computable regularized approximation
$\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}f=f_{\mathcal{X}}^{\top}(K_{\mathcal{X}}+\lambda I)^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}(K_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}.$ (44)
A similar error bound can be computed for the Monte Carlo estimation error $\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}-\widehat{\mathcal{K}}_{\lambda}$ in the regularized case. For this, we define
$g_{\lambda}^{\top}:=f_{\mathcal{X}}^{\top}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}},\qquad(g_{\lambda}^{\mathrm{MC}})^{\top}=f_{\mathcal{X}}^{\top}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}.$
In similar fashion to Theorem 4.6, we obtain the following error bound.
Corollary 4.9 (Regularized Monte Carlo error bound).
Let the assumptions of of Theorem 4.6 hold. For $f\in\mathbb{H}$, the regularized approximant $\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}$ from (44) satisfies
$\|(\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}-\widehat{\mathcal{K}}_{\lambda})f\|_{\infty}\,\leq\,\varepsilon\cdot\|k\|_{\infty}^{1/2}\cdot\sqrt{\frac{d}{\lambda}}\|f\|_{H}$ (45)
with probability at least $1-2de^{-m\varepsilon^{2}/(2\|k\|_{\infty})}$.
Proof.
Again, we factor out the probabilistic part of the Monte Carlo error:
$|(\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}-\widehat{\mathcal{K}}_{\lambda})f(x)|$ $=|(g_{\lambda}-g_{\lambda}^{\mathrm{MC}})^{\top}(K_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}(x)|$
$\leq\|g_{\lambda}-g_{\lambda}^{\mathrm{MC}}\|_{\infty}\|(K_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}.$
Regarding the deterministic second part of the error, from the positive definiteness of $\mathbf{K}_{\mathcal{X}}$, we get $\lambda_{\min}(\mathbf{K}_{\mathcal{X}}+\lambda I)=\lambda_{\min}(\mathbf{K}_{\mathcal{X}})+\lambda\geq\lambda$. Following Remark 4.8, this yields
$\|(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}^{2}\leq k(x,x)\frac{d}{\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}}+\lambda I)}\leq k(x,x)\frac{d}{\lambda}.$
For the other term, we have
$\|g_{\lambda}-g_{\lambda}^{\mathrm{MC}}\|_{\infty}=\max_{x_{j}\in\mathcal{X}}\left|\mathbb{E}[(R_{\mathcal{X}}^{\lambda}f)(Y)|X=x_{j}]-\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}(R_{\mathcal{X}}^{\lambda}f)(y_{l,j})\right|$
with the operator $R_{\mathcal{X}}^{\lambda}$ from (37). By [BoldPhil24, Proposition B.1], that “projection-like” operator satisfies $\|R_{\mathcal{X}}^{\lambda}\|_{\mathbb{H}\to\mathbb{H}}\leq\|S_{\mathcal{X}}\|_{\mathbb{H}\to\mathbb{H}}=1$. Thus, the arguments from the proof of Theorem 4.6 carry over, and for every $f\in\mathbb{H}$, we get
$\mathbb{P}\left(\|g_{\lambda}-g_{\lambda}^{\mathrm{MC}}\|_{\infty}\geq\varepsilon\|f\|_{\mathbb{H}}\right)\leq 2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right).$
In total, this yields (45). ∎
The appeal of Corollary 4.9 lies in the more explicit bound in terms of the number $d$ of data points in $\mathcal{X}$. The regularization parameter $\lambda$ puts a “floor” on the conditioning of the kernel matrix $\mathbf{K}_{\mathcal{X}}$, making the tradeoff between deterministic regularization bias and kernel matrix conditioning evident.
In the final subsection, we combine the deterministic and Monte Carlo error bounds to get a bound on the full $L^{\infty}$-error of the approximations $\widehat{\mathcal{K}}^{\mathrm{MC}}$ and $\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}$.
4.3. Proof of Theorems 2.6 and 2.8
In this part, we combine the results of Theorems 4.2 and 4.6 to obtain the full error bound from Theorem 2.6. After the proof, we provide the respective result for the regularized approximant $\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}$ and a more explicit version in the central special case of a Sobolev space $\mathbb{H}=H^{s}$ generated by a Wendland kernel. First, let us restate our main result for the stochastic kEDMD approximant.
Theorem 0 ($L^{\infty}$-error bound, restated).
Let $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b},C_{b})$. Furthermore, for $\mathcal{X}=\{x_{1},\ldots,x_{d}\}$, let $m_{j}=m\in\mathbb{N}$ i.i.d. samples from every $\rho_{x_{j}}$, $j\in\{1,\dots,d\}$. Then, for every $\varepsilon>0$ and for all $f\in\mathbb{H}$, $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
$\|(\mathcal{K}-\widehat{\mathcal{K}}^{\mathrm{MC}})f\|_{\infty}\leq\big[1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\big]\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}+\varepsilon\cdot\|k\|_{\infty}^{1/2}\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right)$.
Proof.
Let $f\in\mathbb{H}$ with $\|f\|_{\mathbb{H}}\leq 1$. We decompose the full approximation error into its deterministic and probabilistic part by
$\|(\mathcal{K}-\widehat{\mathcal{K}}^{\mathrm{MC}})f\|_{\infty}\leq\|(\mathcal{K}-\widehat{\mathcal{K}})f\|_{\infty}+\|(\widehat{\mathcal{K}}-\widehat{\mathcal{K}}^{\mathrm{MC}})f\|_{\infty}.$
As $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b},C_{b})$, Assumption 4.1 is fulfilled. Theorem 4.2 now bounds the first summand by
$\big\|(\mathcal{K}-\widehat{\mathcal{K}}\big)f\|_{\mathbb{H}\to C_{b}}\,\leq\,(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}\|f\|_{\mathbb{H}}.$
By Theorem 4.6, the second summand admits the bound
$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|k\|_{\infty}^{1/2}\cdot\big(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big)^{1/2}\|f\|_{H}$
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right)$. With $\|f\|_{\mathbb{H}}\leq 1$, those two inequalities yield the claim. ∎
The projection error $\mathcal{I}-S_{\mathcal{X}}$ and the expression involving the kernel matrix $\mathbf{K}_{\mathcal{X}}^{-1}$ warrant a closer inspection in certain special cases. For Sobolev spaces and Wendland RKHS, the projection error in Theorem 2.8 is bounded in terms of the fill distance $h_{\mathcal{X}}$.
Theorem 0 ($L^{\infty}$-error bound for Wendland native spaces, restated).
Let $l\in\mathbb{N}$, define $\sigma_{n,l}=\frac{n+1}{2}+l$ and denote by $\mathbb{H}$ the Sobolev space $H^{\sigma_{n,l}}$ induced by the Wendland kernel $k_{n,l}$ as RKHS. Assume that $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b},C_{b})$. Then, for every $\varepsilon>0$ and all $f\in\mathbb{H}$ with $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
$\|(\mathcal{K}-\mathcal{K}^{\mathrm{MC}})f\|_{\infty}\leq$ $C_{1}\,(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot h_{\mathcal{X}}^{l+1/2}+\varepsilon\cdot C_{2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2C_{2}^{2}}\right)$, with constants $C_{1},C_{2}>0$ only depending on the kernel $k_{n,l}$.
Proof.
The result follows analogously to the proof of Theorem 2.6 by the error decomposition
$\|(\mathcal{K}-\widehat{\mathcal{K}}^{\mathrm{MC}})f\|_{\infty}\leq\|(\mathcal{K}-\widehat{\mathcal{K}})f\|_{\infty}+\|(\widehat{\mathcal{K}}-\widehat{\mathcal{K}}^{\mathrm{MC}})f\|_{\infty}.$
for $f\in\mathbb{H}$. By Corollary 4.3,
$\|\mathcal{K}-\widehat{\mathcal{K}}\|_{\mathbb{H}\to C_{b}}\leq C_{1}(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot h_{\mathcal{X}}^{l+1/2}$
for some $C_{1}>0$. The second summand is bounded by Theorem 4.6, which gives
$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|k_{n,l}\|_{\infty}^{1/2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}\|f\|_{H}$
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k_{n,l}\|_{\infty}}\right)$. Setting $C_{2}=\|k_{n,l}\|_{\infty}^{1/2}$ and taking both inequalities for $\|f\|_{\mathbb{H}}\leq 1$ yields the claim. ∎
Note that Wendland kernels are defined via $k_{n,l}(x,y)=B_{n,l}(\|x-y\|_{2})$ with the radial basis function $B_{n,l}:[0,\infty)\to\mathbb{R}$, see Appendix B. It can be shown that $\|k_{n,l}\|_{\infty}=\sup_{x\in\mathbb{R}^{n}}|k_{n,l}(x,x)|=B_{n,l}(0)=p_{n,l}(0)$ (see, e.g., [PhilSchaWort2025]), where $p_{n,l}$ is the recursively defined polynomial underlying the Wendland kernel [Wend04]. Thus, we get the explicitly computable expression $C_{2}=\sqrt{p_{n,l}(0)}$ for the constant from (21) that depends on dimension $n$ and smoothness $l$. A similar bound as in Theorem 2.8 is easily proven for Matérn kernels, which we state in Corollary C.2 in Appendix C.
To make use of the error bounds from Theorem 2.8 in practice, one would first choose a large enough $d$ to ensure that the fill distance $h_{\mathcal{X}}$ controlling the deterministic projection error is sufficiently small. This might result in a badly conditioned kernel matrix $\mathbf{K}_{\mathcal{X}}$, making the expression $\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$ large. Thus, to control the probabilistic error, one must choose a small enough $\varepsilon>0$ that compensates for the conditioning. Lastly, we want our error bound to hold with probability close to one, so the amount of Monte Carlo samples $m$ for the entries of $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ has to guarantee that $2d\exp(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}})$ is close to zero. Since the factor involving the inverse kernel matrix $\mathbf{K}_{\mathcal{X}}^{-1}$ depends on the size $d$ of data set $\mathcal{X}$, regularization as indicated in Corollary 4.9 prevents this expression from blowing up with $d$. However, regularization introduces a bias to the deterministic error, which is why the final result of this section explicitly juxtaposes deterministic and probabilistic error for the regularized approximant $\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}$.
Corollary 4.12 (Regularized $L^{\infty}$-error bounds).
Let the assumptions of Theorem 2.6 hold and let $\lambda\geq 0$. Then, for every $\varepsilon>0$ and every $f\in\mathbb{H}$ with $\|f\|_{\mathbb{H}}\leq 1$, we have
$\|(\mathcal{K}-\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}})f\|_{\infty}\leq\left(\|1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\right)\,\|\mathcal{I}-R_{\mathcal{X}}^{\lambda}\|_{\mathbb{H}\to C_{b}}\,+\varepsilon\cdot\|k\|_{\infty}^{1/2}\cdot\sqrt{\frac{d}{\lambda}}$ (46)
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right)$. Specifically, for the RKHS $\mathbb{H}=H^{\sigma_{n,l}}$ induced by the Wendland kernel $k_{n,l}$, where $\sigma_{n,l}=\frac{n+1}{2}+l$, we have the pointwise error bound
$\|(\mathcal{K}-\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}})f\|_{\infty}\leq C_{1}\left(h_{\mathcal{X}}^{l+1/2}+\lambda^{1/2}\right)+\varepsilon\cdot C_{2}\cdot\sqrt{\frac{d}{\lambda}}$ (47)
for every $\varepsilon>0$ and all $f\in H^{\sigma_{n,l}}$ with $\|f\|_{H^{\sigma_{n,l}}}\leq 1$ with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2C_{2}^{2}}\right)$ and constants $C_{1},C_{2}>0$ only depending on $k_{n,l}$.
Proof.
The proof of the first claim is analogous to the proof of Theorem 2.6 by splitting
$\|(\mathcal{K}-\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}})f\|_{\infty}\leq\|(\mathcal{K}-\widehat{\mathcal{K}}_{\lambda})f\|_{\infty}+\|(\widehat{\mathcal{K}}_{\lambda}-\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}})f\|_{\infty}$
and bounding the deterministic first part by Proposition 4.5. The estimation error in the second part is bounded by Corollary 4.9 with the respective probability and yields (46).
The second claim combines the second part of Proposition 4.5 for the deterministic error bound with setting $C_{2}=\|k_{n,l}\|_{\infty}^{1/2}$, obtaining (47). ∎
The Matérn kernel analogue to the Wendland special case of Corollary 4.12 is stated in Corollary C.3 in Appendix C.
5. Numerical examples
In this part, we verify the kernel EDMD method for stochastic systems $Y\mid\{X=x\}$ by considering numerical examples from stochastic differential equations (SDEs). For this, we assume that Assumption 4.1 holds and that we can sample from the conditional distribution $\rho_{x}$ of $Y$ given $X=x$. We simulate from SDEs of the form
$dX_{t}$ $=b(X_{t})\,dt\,+\sigma(X_{t})\,dW_{t},$ (48)
$X_{0}$ $=x_{0}\in\Omega$
with Lipschitz continuous drift $b:\mathbb{R}\to\mathbb{R}$ and diffusion $\sigma:\mathbb{R}\to\mathbb{R}$, and Brownian motion $(W_{t})_{t\geq 0}$, via the Euler-Maruyama method (see, e.g., [KloePlat1992]). [Okse2000, Theorem 5.2.1] guarantees the existence of a unique solution $(X_{t})_{t\geq 0}$. A basic example for this case is the Ornstein-Uhlenbeck process (OU process) given by the linear SDE
$dX_{t}$ $=-\theta X_{t}\,dt\,+\sigma\,dW_{t},$ (49)
$X_{0}$ $=x_{0}\in\Omega$
with drift coefficient $\theta>0$ and diffusion coefficient $\sigma>0$.
We fix a timestep $T>0$, set $X=X_{0}$ and $Y=X_{T}$. To apply the setting of Sections 3 and 4, the solution process should remain in some bounded domain $\Omega=(a,b)$ until time $T$. [PhilSchaWort2025] elaborates on a sufficient condition on the drift and diffusion for the invariance of a compact interval under the SDE flow. Another way to guarantee the boundedness is by reflection at the boundaries. We consider this doubly reflected OU process on $\Omega=(0,1)$. It is formally given by the SDE
$dX_{t}$ $=-\theta X_{t}\,dt+\sigma\,dW_{t}+dL_{t}-dU_{t},$ (50)
$X_{0}$ $=x_{0}\in(0,1),$
where the $L_{t}\geq 0$ and $U_{t}\geq 0$ are the minimal nondecreasing processes that ensure $X_{t}\geq 0$ and $X_{t}\leq 1$ with reflection at the boundaries for $t\geq 0$, respectively. The processes $L$ and $U$ can be interpreted as “idleness process” and “loss process”; they satisfy $\int_{0}^{T}\chi_{\{X_{t}>0\}}\,dL_{t}=0$ and $\int_{0}^{T}\chi_{\{X_{t}<1\}}\,dU_{t}=0$ for any $T>0$, respectively. We refer to [HuanMandSpre2014] for details.
As an example of a nonlinear stochastic differential equation, we consider the double-well potential $V(x)=\frac{1}{2}x^{2}-\frac{1}{4}x^{4}$ and set the drift to $b(x)=V^{\prime}(x)$, yielding the SDE
$dX_{t}$ $=(X_{t}-X_{t}^{3})\,dt\,+\sigma\,dW_{t},$ (51)
$X_{0}$ $=x_{0}\in\Omega$
with $\sigma>0$. To this process, we apply an affine transformation to encapsulate the behavior at the wells $\pm 1$ inside the boundaries $a=0$ and $b=1$. We also reflect at those boundaries as described above, thereby bounding the derivative of the drift function, ensuring its Lipschitz continuity.
Algorithm 1 summarizes our method to approximate the action of the stochastic Koopman operator on some observable $f:\Omega\to\mathbb{R}$.
Algorithm 1 Kernel EDMD for Stochastic Koopman Operator Approximation
1:Design points $\mathcal{X}=\{x_{1},\dots,x_{d}\}$; kernel $k$; observable $f$; training parameter $m_{\mathrm{train}}$; regularization parameter $\lambda$; evaluation grid $(x_{\ell}^{\mathrm{grid}})_{\ell=1}^{L}$.
2:Data matrices
3:Form the kernel matrix
$\mathbf{K}_{\mathcal{X}}=\bigl(k(x_{i},x_{j})\bigr)_{i,j=1}^{d}.$
4:Define $f_{\mathcal{X}}=(f(x_{1}),\dots,f(x_{d}))^{\top}$.
5:Sampling step
6:For each $j=1,\dots,d$ draw i.i.d. samples
$y_{j}^{(m)}\sim\rho_{x_{j}},\qquad m=1,\dots,m_{\mathrm{train}}.$
7:Estimate the propagated kernel matrix
$(\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}})_{ij}=\frac{1}{m_{\mathrm{train}}}\sum_{m=1}^{m_{\mathrm{train}}}k(x_{i},y_{j}^{(m)}).$
8:Koopman estimate
9:for $\ell=1,\dots,L$ do
10:  Form kernel vector
$k_{\mathcal{X}}(x_{\ell}^{\mathrm{grid}})=(k(x_{1},x_{\ell}^{\mathrm{grid}}),\dots,k(x_{d},x_{\ell}^{\mathrm{grid}}))^{\top}.$
11:  Compute the estimated Koopman action
$\widehat{\mathcal{K}}^{\mathrm{MC}}f(x_{\ell}^{\mathrm{grid}})=f_{\mathcal{X}}^{\top}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\,\mathbf{K}_{\mathcal{X}}^{-1}\,k_{\mathcal{X}}(x_{\ell}^{\mathrm{grid}}).$
12:end for
13:return $\widehat{\mathcal{K}}^{\mathrm{MC}}f(x_{\ell}^{\mathrm{grid}})$ for $\ell=1,\dots,L$.
In the setting of the stochastic Koopman operator (3), we set $X=X_{0}$ and $Y=X_{T}$. This means that we are interested in the evolution of observables $f:\Omega\to\mathbb{R}$ under the stochastic process that solves (48), i.e.
$\mathcal{K}f(x)=\mathbb{E}[f(X_{T})\mid X_{0}=x].$ (52)
For numerical testing, we can generally proceed as usual in kernel EDMD methods using the approximation (41), that is,
$\widehat{\mathcal{K}}^{\mathrm{MC}}f=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}.$
or the regularized approximant (44), i.e.,
$\widehat{\mathcal{K}}^{\mathrm{MC}}f=f_{\mathcal{X}}^{\top}(\mathbf{K}_{\mathcal{X}}^{-1}+\lambda I)^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}.$
Aside from the observable $f:\Omega\to\mathbb{R}$ and design points $\mathcal{X}=\{x_{1},\dots,x_{d}\}\subset\Omega$ that need to be chosen beforehand, we introduce the parameter $m_{\mathrm{train}}\in\mathbb{N}$ that specifies the number of samples we draw from the conditional distributions $\rho_{x_{j}}$, $j=1,\dots,d$. We note that the error analysis in the stochastic kEDMD setting presents several challenges compared to the deterministic setting:
  1. (1)
    Conditional expectations of the form $\mathcal{K}f(x_{0})=\mathbb{E}[f(X_{T})\mid X_{0}=x_{0}]$ for some $x_{0}\in\Omega$ usually do not necessarily have an analytical expression. Therefore, we do not have access to the exact action of the Koopman operator on the observable $f$ (”ground truth”). Instead, we use a further Monte Carlo approximation
    $\mathcal{K}f(x_{\mathrm{test}})\approx\frac{1}{m_{\mathrm{truth}}}\sum_{l=1}^{m_{\mathrm{test}}}f(y_{l})$ (53)
    where $y_{l}\overset{\text{i.i.d.}}{\sim}\rho_{x_{\mathrm{test}}}$ are i.i.d. samples of the simulated SDE solution at time $T$ and $0\ll m_{\mathrm{truth}}\in\mathbb{N}$ is our “ground truth parameter”. It should also be noted that we introduce a discretization error by simulating the SDE via the Euler-Maruyama method. To control this error, a fine enough grid on $\Omega=(0,1)$ is employed.
  2. (2)
    Even for an analytically calculated ground truth, we can only approximate the $L^{\infty}$-error over some grid on $\{x_{1}^{\mathrm{grid}},\dots,x_{L}^{\mathrm{grid}}\}\subset\Omega$. This gives us an “empirical” approximated $L^{\infty}$-error $\widehat{\mathrm{err}}=\max_{l}|\mathcal{K}f(x_{l}^{\mathrm{grid}})-\widehat{\mathcal{K}}^{\mathrm{MC}}f(x_{l}^{\mathrm{grid}})|$.
  3. (3)
    The randomness of the propagated matrix $K_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ necessitates a comparison over multiple realizations of the Koopman approximation. In our SDE setting, this means that we build matrices $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC,n}}$, $n=1,\dots,n_{\mathrm{pred}}$ and use the mean values and standard deviations of the empirical $L^{\infty}$-errors $\widehat{\mathrm{err}}^{(n)}$ per realization to judge the performance of the approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$.
Theorem 2.6 decomposes the $L^{\infty}$-error of the stochastic Koopman approximation into a deterministic part controlled by the fill distance $h_{\mathcal{X}}$ and a probabilistic part influenced by the Monte Carlo sample size, which we denote by $m_{\mathrm{train}}$ to differentiate it from the number of Monte Carlo samples $m_{\mathrm{truth}}$ used for ground truth approximation. For numerical testing, we specify $h_{\mathcal{X}}$ (or corresponding design points $\mathcal{X}$) and $m_{\mathrm{train}}$. Throughout the testing procedure, we frequently sample the SDE solution with initial values $x_{0}\in\Omega$ at time $T>0$ from the conditional distribution $\rho_{x_{0},T}$. First, we start trajectories at chosen evaluation grid points $\mathcal{X}^{\mathrm{grid}}\subset\Omega$ for ground truth computation. Later, we repeatedly simulate trajectories starting at the design points $\mathcal{X}$ to construct realizations $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ in the Koopman approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$ Algorithm 2 illustrates the procedure.
Algorithm 2 Error Analysis for the Stochastic Kernel EDMD Koopman Approximation
1: number of Koopman approximation realizations $n_{\mathrm{pred}}\in\mathbb{N}$; skEDMD parameters = {fill distance $h_{\mathcal{X}}$; kernel $k$; training parameter $m_{\mathrm{train}}$; observable $f$; regularization parameter $\lambda$; evaluation grid $\mathcal{X}^{\mathrm{grid}}=\{x_{\ell}^{\mathrm{grid}}\}_{\ell=1}^{L}$}; simulation parameters = {SDE parameters; SDE simulator; timestep T}; ground truth parameters.
2:Ground truth
3:for $\ell=1,\dots,L$ do
4:  For $m=1,\dots,m_{\mathrm{truth}}$, simulate $y_{\ell}^{(m)}\sim\rho_{x_{\ell},T}$ with simulation parameters and set
$\mathcal{K}_{\mathrm{truth}}f(x_{\ell}^{\mathrm{grid}})=\frac{1}{m_{\mathrm{truth}}}\sum_{m=1}^{m_{\mathrm{truth}}}y_{\ell}^{(m)}$
5:end for
6:Design points
7:Generate equidistant $\mathcal{X}=\{x_{i}\}_{i=1}^{d}$ for fill distance (at least) $h_{\mathcal{X}}$.
8:Kernel EDMD iterations
9:for $n=1,\dots,n_{\mathrm{pred}}$ do
10:  Apply Algorithm 1 with skEDMD and simulation parameters to obtain approximations $\widehat{\mathcal{K}}^{\mathrm{MC}}f(x_{\ell}^{\mathrm{grid}})$ for $l=1,\dots,L$
11:  Compute approximation of the $L^{\infty}$-error over the evaluation grid:
$\widehat{\mathrm{err}}_{n}=\max_{\ell}\,\lvert\mathcal{K}_{\mathrm{truth}}f(x_{\ell}^{\mathrm{grid}})-\widehat{\mathcal{K}}^{\mathrm{MC}}f(x_{\ell}^{\mathrm{grid}})\rvert$
12:end for
13:return mean and standard deviation of $\widehat{\mathrm{err}}_{n}$, $n=1,\dots,n_{\mathrm{pred}}$.
For numerical stability, we apply a Tikhonov regularization with parameter $\lambda$ as in (44) to alleviate the bad conditioning of kernel matrices $\mathbf{K}_{\mathcal{X}}$, . The parameters $m_{\mathrm{train}}$ and $\lambda$ are occasionally adapted to the chosen fill distance $h_{\mathcal{X}}$ as indicated in the table of Figure 1.
To settle on a number of realizations $n_{\mathrm{pred}}$ to control for the randomness in $\widehat{\mathcal{K}}^{\mathrm{MC}}$, we conducted numerical tests for different choices of $h_{\mathcal{X}}$ and $m_{\mathrm{train}}$. Our tests showed that the standard deviation of the empirical $L^{\infty}$-error seems to stabilize after as little as $n_{\mathrm{pred}}=30$ realizations. Therefore, we average approximation errors over $30$ realizations in the following.
The remainder of this section studies the error behavior for the linear Ornstein-Uhlenbeck process and the nonlinear process governed by the double-well potential in dependence of the fill distance $h_{\mathcal{X}}$ and Monte Carlo sample size $m_{\mathrm{train}}$. In all cases, we consider the nonlinear observable $f(x)=e^{x}$. We chose to compare the
$\text{Wendland kernel}\quad k_{1,1}^{\text{wendland}}(x,y)$ $=(1-r)_{+}^{3}\cdot(3r+1),~\text{and}$
[…]
where we denote $r=\|x-y\|_{2}$.
5.1. Linear SDE: Ornstein-Uhlenbeck process
In the linear Ornstein-Uhlenbeck SDE (49), we set $\theta=1$, $\sigma=0.2$ and $T=1$ to simulate from
[…]
$X_{0}$ $=x_{0}\in\Omega$
with initial value $x_{0}$ and auxiliary processes $L$ and $U$ to achieve reflection at the boundaries of $\Omega=(0,1)$.
As laid out in Algorithm 2, we approximate the $L^{\infty}$-error per realization as maximal absolute error over all evaluation grid points. We then average over all $n_{\mathrm{pred}}$ realizations to obtain an empirical $L^{\infty}$-error and compute its standard deviation.
Adaptive parameters
$h_{\mathcal{X}}$ $\lambda$ $m_{\mathrm{train}}$
$\cdot 10^{-1} 11$$\cdot 10^{-3}$ 20
$5\cdot 10^{-2}$ $\frac{5}{4}\cdot 10^{-4}$ 30
$2\cdot 10^{-2}$ $8\cdot 10^{-6}$ 127
$\cdot 10^{-2} 11$$\cdot 10^{-6}$ 387
$5\cdot 10^{-3}$ $\frac{5}{4}\cdot 10^{-7}$ 1183
$2\cdot 10^{-3}$ $8\cdot 10^{-9}$ 5166
$\cdot 10^{-3} 11$$\cdot 10^{-9}$ 15684
$5\cdot 10^{-4}$ $\frac{5}{4}\cdot 10^{-10}$ 47428
Figure 1. Ornstein–Uhlenbeck process: mean empirical $L^{\infty}$-error of $n_{\mathrm{pred}}=30$ realizations of the stochastic Koopman approximation as a function of the fill distance $h_{\mathcal{X}}$. The plot compares constant and adaptive choices of $m_{\mathrm{train}}$ and $\lambda$; the table reports the adaptive parameter values used for each $h_{\mathcal{X}}$. Error bars indicate standard deviation over the realizations.
The left panel in Figure 1 is a first display of the dependence of the deterministic part of the error on the fill distance for both Matérn and Wendland kernel. However, after an initially steeper decline of error with fill distance, the curve seems to flatten where we would expect a linear decay in the log-log-plot from the deterministic error bound of Theorem 4.2.
Figure 2. Ornstein-Uhlenbeck process: mean empirical $L^{\infty}$-error with standard deviation of $n_{\mathrm{pred}}=30$ realizations of the stochastic Koopman approximation in dependence of the number of Monte Carlo simulations $m_{\mathrm{train}}$. $\star$ and $\diamond$ indicate Wendland and Matérn kernels, respectively.
In order to guarantee decay with the fill distance with the suggested rate, we have to adapt $m_{\mathrm{train}}$ to the fill distance to ensure the probabilistic error behaves similarly and does not become dominant for smaller fill distances. Appendix D expands on that balancing. We heuristically chose $m_{\mathrm{train}}(h_{\mathcal{X}})=O(h_{\mathcal{X}}^{-3/2})=O(d^{3/2})$ and $\lambda(h_{\mathcal{X}})=O(h_{\mathcal{X}}^{3})=O(d^{-3})$. The right panel of Figure 1 depicts how decreasing fill distances combined with adapted Monte Carlo parameters $m_{\mathrm{train}}$ and regularization parameters $\lambda$ result in smaller empirical $L^{\infty}$-errors. The chosen growth rate of the Monte Carlo parameter $m_{\mathrm{train}}$ in dependence of the fill distance $h_{\mathcal{X}}$ is far more conservative than the theoretically derived one in Appendix D. Still, after adapting the parameters $m_{\mathrm{train}}$ and $\lambda$, the decay appears close to linear on the log-scale. The Wendland kernel and Matérn kernel perform similarly well and both achieve a comparable decay of the $L^{\infty}$-error.
Another insight is gained by fixing a fill distance and only increasing $m_{\mathrm{train}}$, decreasing the probabilistic error with fixed deterministic error bound. The probabilistic effect is illustrated in Figure 2 for various fill distances. For a fixed $m_{\mathrm{train}}$, a smaller fill distance results in a smaller $L^{\infty}$-error. This is consistent with the error bounds on the deterministic, probabilistic and full $L^{\infty}$-error proven above. Again, the performance of Wendland and Matérn kernel is comparable.
5.2. Nonlinear SDE: Rescaled double-well potential
For the nonlinear SDE with double-well potential, we also set $\sigma=0.2$ and $T=1$. For better comparability, especially in terms of the fill distance, we want to restrict the corresponding SDE to the same domain $\Omega=(0,1)$. Since the wells in the regular double-well SDE (51) are at $\pm 1$, we apply the affine transformation $Z_{t}=(X_{t}+2)/4$ to mimic the behavior of the original double-well potential on (2,2)(-2,2) on our domain $\Omega=(0,1)$. Solving for $X_{t}$ gives $X_{t}=4Z_{t}-2$, so that the chain rule yields the SDE
$dZ_{t}$ $=\frac{1}{4}\left[(4Z_{t}-2)-(4Z_{t}-2)^{3}\right]\,dt\,+\frac{\sigma}{4}\,dW_{t}+dL_{t}-dU_{t}$ (55)
$Z_{0}$ $=z_{0}\in\Omega$
with the auxiliary processes $L$ and $U$ achieving reflection at the boundaries $a=0$ and $b=1$.
Figure 3. Double-well SDE: mean empirical $L^{\infty}$-error with standard deviation of $n_{\mathrm{pred}}=30$ realizations of the stochastic Koopman approximation in dependence of fill distance $h_{\mathcal{X}}$ and Monte Carlo sample size $m_{\mathrm{train}}$. Parameters in the left panel were adapted as indicated by the table in Figure 1. In the right panel, $\star$ and $\diamond$ indicate Wendland and Matérn kernels, respectively.
As Figure 3 shows, the double-well SDE demonstrates similar error behavior. Mirroring the adaptive parameter regime from the linear OU process above, we scaled $m_{\mathrm{train}}$ and $\lambda$ with the fill distance to demonstrate the decay of the $L^{\infty}$-error. The decrease of the probabilistic error for various fixed fill distances also follows a similar pattern as in Figure 2. Furthermore, we see that, while comparable in performance, Matérn kernels seem to have a slight edge over Wendland kernels for each given fill distance.
6. Conclusions and Outlook
We proposed a Monte Carlo-based estimator in a kernel EDMD approach for the Koopman operator of stochastic systems and analyzed its pointwise error. Extending the results of [KohnPhil24], we decompose the error into the well-known deterministic projection error and the probabilistic error introduced by the entry-wise estimation of expected values in the exact “propagation matrix” $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$. To apply known projection error bounds, we derive sufficient conditions for the invariance of (fractional) Sobolev spaces $H^{s}$ under the stochastic Koopman operator, which are induced as RKHS by Wendland and Matérn kernels. Novel probabilistic error bounds for our estimator are proven to obtain a pointwise bound on the full approximation error with high probability.
In this work, we only considered two exemplary kernels, i.e., Mátern and Wendland. In future work, the respective error bounds may be derived for other kernels and their corresponding RKHSs. Furthermore, a canonical (and important) extension is the one to multilevel Monte Carlo (MLMC) to increase sample and, thus, computational efficiency by variance reduction.
Appendix A Proof of Monte Carlo error bound from Remark 4.7
Applying Hölder’s inequality according to Remark 4.7 with $p=q=2$ instead of the approach $p=\infty$, $q=1$ from Theorem 4.6, one obtains
$|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f(x)|\leq\|g-g^{\mathrm{MC}}\|_{2}\|K_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{2}.$
We then have
$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{2}\geq\varepsilon\|f\|_{H})$ $=\mathbb{P}\left(\sum_{j=1}^{d}\left|\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}\big[(S_{\mathcal{X}}f)(y_{l,j})-\mathbb{E}_{j}\big]\right|^{2}\,\geq\,\varepsilon^{2}\|f\|_{H}^{2}\right)$
$\leq\mathbb{P}\left(\exists j\in\{1,\dots,d\}:\left|\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}\big[(S_{\mathcal{X}}f)(y_{l,j})-\mathbb{E}_{j}\big]\right|^{2}\,\geq\,\frac{\varepsilon^{2}\|f\|_{H}^{2}}{d}\right)$
$\leq\sum_{j=1}^{d}\mathbb{P}\left(\left|\sum_{l=1}^{m_{j}}\big[(S_{\mathcal{X}}f)(y_{l,j})-\mathbb{E}_{j}\big]\right|\,\geq\,\frac{\varepsilon m_{j}\|f\|_{H}}{\sqrt{d}}\right)$
$\leq 2\sum_{j=1}^{d}\exp\left(-\frac{m_{j}\varepsilon^{2}}{2d\|k\|_{\infty}}\right).$
Hence, if $m_{j}=m$ for all $j\in\{1,\dots,d\}$, then
$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{2}\geq\varepsilon\|f\|_{H})\,\leq\,2d\exp\left(-\frac{m\varepsilon^{2}}{2d\|k\|_{\infty}}\right).$
We may improve this by using the Hilbert-space Hoeffding inequality. For this, let $m_{j}=m$ for all $j$, and set $\xi_{l}=[(S_{\mathcal{X}}f)(y_{l,j})]_{j=1}^{d}-g\in\mathbb{R}^{d}$, $l\in\{1,\dots,m\}$. Then $\mathbb{E}[\xi_{l}]=0$ and
$\|\xi_{l}\|_{2}^{2}=\sum_{j=1}^{d}\left|(S_{\mathcal{X}}f)(y_{l,j})-\int(S_{\mathcal{X}}f)(y)\,\rho(x,dy)\right|^{2}.$
Since $|(S_{\mathcal{X}}f)(y)|\leq\|k\|_{\infty}^{1/2}\|f\|_{H}$, we obtain $\|\xi_{l}\|_{2}^{2}\leq 4d\|k\|_{\infty}\|f\|_{H}^{2}$ for all $l\in\{1,\dots,m\}$. Furthermore,
$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{2}\geq\varepsilon\|f\|_{H})$ $=\mathbb{P}\left(\sum_{j=1}^{d}\left|\frac{1}{m}\sum_{l=1}^{m}\xi_{l,j}\right|^{2}\,\geq\,\varepsilon^{2}\|f\|_{H}^{2}\right)$
$=\mathbb{P}\left(\left\|\frac{1}{m}\sum_{l=1}^{m}\xi_{l}\right\|\,\geq\,\varepsilon\|f\|_{H}\right)$
$\leq 2\exp\left(-\frac{m\varepsilon^{2}}{8d\|k\|_{\infty}}\right).$
We are left with estimating $\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{2}$. The map $\tau:V_{\mathcal{X}}\to\mathbb{R}^{d}$, $\tau f=f_{\mathcal{X}}$, is surjective (namely, if $v\in\mathbb{R}^{d}$ is given, then $\tau(v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}})=v$), hence
$\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{2}$ $=\sup_{\|v\|_{2}=1}|v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)|=\sup_{f\in V_{\mathcal{X}},\,f\neq 0}\frac{|f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)|}{\|f_{\mathcal{X}}\|_{2}}=\sup_{f\in V_{\mathcal{X}},\,f\neq 0}\frac{|f(x)|}{\|f_{\mathcal{X}}\|_{2}}$
$=\sup_{f\in V_{\mathcal{X}},\,f\neq 0}\frac{|\langle f,\Phi_{x}\rangle_{H}|}{\|f_{\mathcal{X}}\|_{2}}\,\leq\,\|k\|_{\infty}^{1/2}\sup_{f\in V_{\mathcal{X}},\,f\neq 0}\frac{\|f\|_{H}}{\|f_{\mathcal{X}}\|_{2}}=\|k\|_{\infty}^{1/2}\|\mathbf{K}_{\mathcal{X}}^{-1}\|_{2\to 2}^{1/2}.$
Alternatively, by the $\ell^{1}$-$\ell^{2}$-inequality, one may bound the deterministic term as in Theorem 4.6 by $\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{2}\leq\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}\leq\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$.
Appendix B Recap: Wendland kernels
Wendland kernels are constructed using particular compactly supported radial basis functions, which are constructed in the following way: For some integer $m\in\mathbb{N}$, define the a function $\beta_{m}:\mathbb{R}_{+}\to\mathbb{R},\beta_{m}(r)=(1-r)_{+}^{m}=\max(1-r,0)^{m}$. Given the operator $\mathcal{I}$ that acts on measurable non-negative functions $\beta:\mathbb{R}_{+}\to\mathbb{R}_{+}$ by
$\mathcal{I}\beta(r)=\int_{r}^{\infty}t\beta(t)\,dt$
and dimension $n\geq 1$, we choose a smoothness $l\in\mathbb{N}$ and set $\beta_{n,l}=\mathcal{I}^{l}\beta_{\lfloor\frac{n}{2}\rfloor+l+1}$. This function is of the form
$\beta_{n,l}(r)=\begin{cases}p_{n,l}(r),&0\leq r\leq 1,\\ 0,&r>1\end{cases}$
with a univariate polynomial $p_{n,l}$ of degree $\lfloor\frac{n}{2}\rfloor+3l+1$ whose coefficients can be computed recursively by [Wend04, Theorem 9.12]. We can now define the Wendland radial basis function $B_{n,m}:\mathbb{R}^{n}\to\mathbb{R}$ of dimension $n$ and smoothness degree $m$ by
$B_{n,l}(x):=\beta_{n,l}(\|x\|_{2})$
and the corresponding Wendland kernel by $k_{n,l}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R},k_{n,l}(x,y)=B_{n,l}(x-y)=\beta_{n,l}(\|x-y\|_{2})$.
Appendix C Error bounds for Matérn kernels
Similarly to the Wendland kernels, Matérn kernels are constructed via radial basis functions. Specifically, the Matérn covariance function $\mathcal{M}_{\nu,\alpha}:\mathbb{R}_{+}\to\mathbb{R}$ with smoothness $\nu>0$ and length-scale parameter $\alpha>0$ is given by
$\mathcal{M}_{\nu,\alpha}(r)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}\,r}{\alpha}\right)^{\nu}K_{\nu}\left(\frac{\sqrt{2\nu}\,r}{\alpha}\right)$
where $\Gamma$ denotes the gamma function and $K_{\nu}$ denotes the modified Bessel function of the second kind. Denoting the Euclidean distance of $x,y\in\mathbb{R}^{n}$ by $\|\cdot\|_{2}$, the Matérn kernel $k_{\mu,\alpha}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ is given by
$k_{\nu,\alpha}(x,y)=\mathcal{M}_{\nu,\alpha}(\|x-y\|_{2}).$
The parameter $\nu$ controls the differentiability of the Matérn kernel at $x=y$ through $k_{\nu,\alpha}\in C^{2\lfloor\nu\rfloor}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ – with increasing $\nu$, the Matérn kernels become smoother. Equivalently, the sample paths of a Gaussian process with the Matérn kernel as covariance function are $\lfloor\nu\rfloor$-times mean-square differentiable, which provides a flexible way to control smoothness and makes Matérn kernels a popular choice in this setting. Another central property is the convergence
$\mathcal{M}_{\nu,\alpha}(r)\xrightarrow[]{\nu\to\infty}\exp\left(-\frac{r^{2}}{2\alpha^{2}}\right)$
to the squared exponential (Gaussian) kernel, uniformly on compact sets. Of particular interest are the cases $\nu=s+1/2$ for $s\in\mathbb{N}$: In that case, the Matérn covariance function is given as the product of a negative exponential and a polynomial of order $s$ [PorcBeviSchabOate2024]. Up to scaling constants, the Matérn kernel $k_{\nu,\alpha}$ on $\mathbb{R}^{n}$ induces the Sobolev space $H^{\nu+n/2}$ with equivalent norms; for the dimensionless Matérn kernel $k_{\nu}:=k_{\nu,\sqrt{2\nu}}$ the RKHS norm even coincides with the standard Sobolev norm of order $\nu+n/2$ up to a constant factor. We refer the reader to [FassYe11, PorcBeviSchabOate2024] for more insights into Matérn kernels.
Matérn kernels $k_{\nu}(x,y)=\mathcal{M}_{\nu,1}(\|x-y\|_{2})$ of smoothness $\nu>0$ induce the Sobolev spaces $H^{\nu+n/2}$. Denote the native norm of the RKHS by $\|f\|_{\mathbb{H}}$. The Matérn kernel is $2\lfloor\nu\rfloor$-times differentiable. Furthermore, (normalized) Matérn kernels are uniformly bounded by $1$. Thus, according to the discussion after [Wend04, Theorem 11.13], the bound
$\|(\mathcal{I}-S_{\mathcal{X}})f\|_{\infty}\leq Ch_{\mathcal{X}}^{\lfloor\nu\rfloor}\|f\|_{\mathbb{H}}$
with a constant $C>0$ holds for all $f\in\mathbb{H}$.
Using the above properties, we obtain quasi-analogous results as in Section 4 for kernel EDMD approximations with underlying Matérn kernels. Following the derivations for Wendland kernels, we start with the deterministic error bound characterized by the projection error.
Corollary C.1 (Deterministic error bound in Matérn native spaces).
Let $\nu>0$. Denote by $k_{\nu}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ the (dimensionless) Matérn kernel generating the Sobolev space $H^{\nu+n/2}(\Omega)$ with equivalence between the native norm $\|\cdot\|_{\mathbb{H}}$ and the classical Sobolev norm $\|\cdot\|_{H^{\nu+n/2}}$. Assume that $\mathcal{K}\in L(C_{b},C_{b})$ and that (23) is satisfied for some $\ell\in\mathbb{N}$ with $\ell>\nu+n/2$. Then, the Koopman approximation $\widehat{\mathcal{K}}$ from (31) satisfies
$\|\mathcal{K}-\widehat{\mathcal{K}}\|_{\mathbb{H}\to C_{b}}\leq C(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot h_{\mathcal{X}}^{\lfloor\nu\rfloor}$ (56)
Proof.
The Matérn kernel satisfies $k_{\nu}\in C^{\lfloor 2\nu\rfloor}(\mathbb{R}^{n}\times\mathbb{R}^{n})$. According to [Wend04, Theorem 11.13] and the subsequent discussion, we have
$\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}\leq h_{\mathcal{X}}^{\lfloor\nu\rfloor}$
since the Matérn kernel’s derivatives of order $2\lfloor\nu\rfloor$ are continuous on $\overline{\Omega}\times\overline{\Omega}$, allowing for a uniform bound on those derivatives on that compact set. The assertion (56) then follows analogously as in the proof of Corollary 4.3. ∎
Since the probabilistic error bound for the Matérn case is virtually identical to the Wendland case, merely inserting $\|k_{\nu}\|_{\infty}=1$, we state the analogue to Theorem 2.8 for the Matérn case below.
Corollary C.2 ($L^{\infty}$-error bound for Matérn native spaces).
Let the assumptions of Theorem 2.6 hold. Let $\nu>0$ and denote by $\mathbb{H}=H^{\nu+n/2}$ the reproducing kernel Hilbert space induced by the Matérn kernel $k_{\nu}$. Then, for every $\varepsilon>0$ and all $f\in\mathbb{H}$ with $\|f\|_{\mathbb{H}}$, the full error admits the bound
$\|(\mathcal{K}-\mathcal{K}^{\mathrm{MC}})f\|_{\infty}\leq$ $C\,(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot h_{\mathcal{X}}^{\lfloor\nu\rfloor}+\varepsilon\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$ (57)
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2}\right)$.
Proof.
the Matérn covariance function is monotonically decreasing for $r>0$, admitting its maximum at $\mathcal{M}_{\nu,\alpha}(0)=1$, so $\|k_{\nu}\|_{\infty}=1$. Similarly to Theorem 2.8, the result now follows immediately from Theorem 2.6. ∎
Applying Tikhonov regularization with some $\lambda>0$, as is usual in practice, gives rise to the error bound
$\|\mathcal{K}f-\widehat{\mathcal{K}}_{\lambda}f\|_{\infty}\leq C\left(h_{\mathcal{X}}^{\lfloor\nu\rfloor}+\lambda^{1/2}\right)\|f\|_{\mathbb{H}}$ (58)
for $\mathbb{H}=H^{\nu+n/2}$, following the line of argument in Proposition 4.5. To conclude, we state the analogue to the second part Corollary 4.12 for Matérn native spaces.
Corollary C.3 (Regularized $L^{\infty}$-error bounds for Matérn native spaces).
Let the assumptions of Theorem 2.6 hold and let $\lambda\geq 0$. Then, for the RKHS $\mathbb{H}=H^{\nu+n/2}$ induced by the Matérn kernel $k_{\nu}$, we have the pointwise error bound
$\|(\mathcal{K}-\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}})f\|_{\infty}\leq C\left(h_{\mathcal{X}}^{l+1/2}+\lambda^{1/2}\right)+\varepsilon\cdot\sqrt{\frac{d}{\lambda}}$ (59)
for every $\varepsilon>0$ and all $f\in H^{\sigma_{n,l}}$ with $\|f\|_{\mathbb{H}}\leq 1$ with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2}\right)$ and constant $C>0$ only depending on $k_{\nu}$.
Proof.
The proof is analogous to the second part of Corollary 4.12 and leverages the uniform bound $\|k_{\nu}\|_{\infty}=1$. ∎
Appendix D Balancing between deterministic and probabilistic error
We discuss here the practical choice of the parameters $h_{\mathcal{X}}$, $m_{\mathrm{train}}$ and $\lambda$ for the stochastic kernel EDMD method when using the Wendland $C^{2}$ kernel ($l=1$) with spatial dimension $n=1$. The derivation for the Matérn kernel can be done in a similar fashion.
The deterministic error associated with the projection onto the span of the kernel translates, for Wendland smoothness $l=1$, into the estimate
$\|\mathcal{K}f-\widehat{\mathcal{K}}f\|_{\infty}\leq C\|f\|_{\mathbb{H}}\,h_{\mathcal{X}}^{\,l+1/2}=C\|f\|_{\mathbb{H}}\,h_{\mathcal{X}}^{3/2}.$
from Corollary 4.3. Thus, decreasing fill distance improves the deterministic accuracy at rate $O(h_{\mathcal{X}}^{3/2})$. As $h_{\mathcal{X}}$ decreases, the number $d$ of design points increases, and the smallest eigenvalue of the kernel matrix $\mathbf{K}_{\mathcal{X}}$ decreases. For the unregularized inverse, the bound $\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\leq\frac{d}{\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})}$ and $\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})=O(h_{\mathcal{X}}^{3})=O(d^{-3})$ give
$\left(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X},\lambda}^{-1}v\right)^{1/2}=O\left(\sqrt{\frac{d}{d^{-3}}}\right)=O(d^{2})=O(h_{\mathcal{X}}^{-2}).$
Under Tikhonov regularization, Corollary 4.12 shows that $\left(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X},\lambda}^{-1}v\right)^{1/2}=O\!\left(\sqrt{\frac{d}{\lambda}}\right)$. However, regularization introduces an additional deterministic bias:
$\|\mathcal{K}f-\widehat{\mathcal{K}}_{\lambda}f\|_{\infty}\leq C\left(h_{\mathcal{X}}^{3/2}+\lambda^{1/2}\right)\|f\|_{\mathbb{H}}.$
To prevent the regularization bias from dominating, one should require $\lambda=O(h_{\mathcal{X}}^{3})$.
Monte Carlo estimation of $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$ introduces a stochastic error. From Theorem 4.6, with probability at least $1-2de^{-m\varepsilon^{2}/(2\|k\|_{\infty})}$,
$\|\widehat{\mathcal{K}}f-\widehat{\mathcal{K}}^{\mathrm{MC}}f\|_{\infty}\leq\varepsilon\,C_{\mathcal{X}}\|f\|_{\mathbb{H}}.$
Solving the probability expression for $\varepsilon$ gives
$\varepsilon=O(m^{-1/2}),$
again ignoring logarithmic terms in $d$. Since equispaced grids satisfy $d=O(h_{\mathcal{X}}^{-1})$, the Tikhonov control factor becomes
$\sqrt{\max_{v}v^{\top}\mathbf{K}_{\mathcal{X},\lambda}^{-1}v}=O\!\left(\sqrt{\frac{d}{\lambda}}\right)=O\!\left(\frac{h_{\mathcal{X}}^{-1/2}}{h_{\mathcal{X}}^{3/2}}\right)=O(h_{\mathcal{X}}^{-2}).$
To ensure the probabilistic error is of the same order as the deterministic error $O(h_{\mathcal{X}}^{3/2})$, we solve $\varepsilon\cdot O(h_{\mathcal{X}}^{-2})=O(h_{\mathcal{X}}^{3/2})$, which gives $\varepsilon=O(h_{\mathcal{X}}^{7/2})$. Since $\varepsilon=O(m^{-1/2})$, this yields
$m_{\mathrm{train}}(h_{\mathcal{X}})=O\!\left(h_{\mathcal{X}}^{-7}\right)=O(d^{7}).$
Since this is infeasible in practice, we conducted numerical experiments to test more moderate orders of growth of $m_{\mathrm{train}}$. We found the orders
$m_{\mathrm{train}}(h_{\mathcal{X}})=O\!\left(h_{\mathcal{X}}^{-3/2}\right),\quad\lambda(h_{\mathcal{X}})=O\!\left(h_{\mathcal{X}}^{-3}\right)$
to produce satisfying balancing of the two error sources, as displayed in Figure 1.

### v2

@@
 1. Introduction
-Nowadays, the Koopman operator is an established mathematical and computational tool for modeling, analysis, and control of dynamical systems, see, e.g., the collection [MaurSusu20] or the recent overview papers [brunton:budisic:kaiser:kutz:2022, strasser2026overview].
+Nowadays, the Koopman operator is an established mathematical and computational tool for modeling, analysis, and control of dynamical systems, see, e.g., the collection [30] or the recent overview papers [10, 47].
 The underlying idea is to consider nonlinear dynamics through the lens of observable functions using the linear, but infinite-dimensional Koopman operator, see the original work by B.O.
-Koopman [Koop31, KoopNeum32] and the seminal paper [Igor2005] by I.
+Koopman [26, 25] and the seminal paper [32] by I.
 Mezić.
-This enabled, among others, to leverage modern techniques from spectral analysis to investigate the complex dynamical behaviour of the original nonlinear system [rowley2009spectral].
-Recently, data-driven approaches like dynamic mode decomposition (EDMD; [WillKevr15, Colb23]) were proposed to approximate the Koopman operator, see, e.g., the recent overview [Colb24], which has lead to a plethora of successful applications, including fluid flows [bagheri:2013, mezic:2013], climate forecasting [hogg:fonoberova:mezic:2020, azencot:erichson:lin:mahoney:2020], molecular dynamics [nuske:klus:2023], power systems technology [susuki:mezic:raak:hikihara:2016], quantum systems [klus:nuske:peitz:2022], robotics [mamakoukas:castano:tan:murphey:2019, bruder:fu:gillespie:remy:vasudevan:2020, haggerty:banks:kamenar:cao:curtis:mazic:hawkes:2023, rosenfelder:bold:eschmann:eberhard:worthmann:ebel:2024], and human-centered applications such as grip force prediction for robotic rehabilitation [bazina:kamenar:fonoberova:mezic:2025].
+This enabled, among others, to leverage modern techniques from spectral analysis to investigate the complex dynamical behaviour of the original nonlinear system [44].
+Recently, data-driven approaches like dynamic mode decomposition (EDMD; [53, 11]) were proposed to approximate the Koopman operator, see, e.g., the recent overview [13], which has lead to a plethora of successful applications, including fluid flows [4, 33], climate forecasting [17, 3], molecular dynamics [35], power systems technology [48], quantum systems [23], robotics [29, 9, 16, 43], and human-centered applications such as grip force prediction for robotic rehabilitation [5].
 In EDMD, snapshots of the underlying system are employed to approximate the Koopman operator.
-It has been shown in [KordaMezi18] that EDMD converges in the infinite-data and infinite-dictionary (infinitely many observable functions) corresponding to a complexity solvability index two [colbrook2022foundations].
-Then, first finite-data results on the estimation error for ergodic sampling for deterministic systems were provided in [Mezi22] followed by [ZhanZuaz23] analyzing also the $L^{2}$ projection error onto a finite element dictionary, cf.
-also the works [NuskPeit23, PhilScha24] providing error bounds for stochastic and control systems.
-More recently, kernel EDMD approximation schemes, which allow a data-informed choice of the dictionary, were proposed, see, e.g., [WillRowl15, klus:nuske:hamzi:2020].
-Several works have already addressed $L^{2}$-error bounds for kernel EDMD: [PhilScha24:kernel] provides these in terms of the Mercer eigenvalues of the chosen kernel in the case of ergodic sampling along a trajectory, while the bounds in [PhilSchaWort2025, OlguOsseRami2025], are derived for i.i.d.
+It has been shown in [27] that EDMD converges in the infinite-data and infinite-dictionary (infinitely many observable functions) limit, corresponding to a complexity solvability index of two [12].
+Then, first finite-data results on the estimation error for ergodic sampling for deterministic systems were provided in [34] followed by [54] analyzing the $L^{2}$-projection error onto a finite element dictionary, cf.
+also the works [36, 39] providing error bounds for stochastic and control systems.
+More recently, kernel EDMD approximation schemes, which allow a data-informed choice of the dictionary, were proposed, see, e.g., [52, 22].
+Several works have already addressed $L^{2}$-error bounds for kernel EDMD: [40] provides these in terms of the Mercer eigenvalues of the chosen kernel in the case of ergodic sampling along a trajectory, while the bounds in [41, 38], are derived for i.i.d.
 sampling.
-The work [KohnPhil24] shows first pointwise (i.e., $L^{\infty}$) error bounds for approximations using Wendland kernels.
+The work [24] shows first pointwise (i.e., $L^{\infty}$) error bounds for approximations using Wendland kernels.
 However, while kernel EDMD is successfully used for stochastic systems, pointwise error bounds are still missing – to the best of the authors’ knowledge.
@@
 To this end, we establish Koopman invariance of (fractional) Sobolev spaces based on novel sufficient conditions on a Radon-Nikodym density associated with the stochastic dynamical system.
-Then we split the approximation error into a deterministic and a probabilistic part, where the latter results from Monte Carlo sampling of the expected value.
+Then we split the approximation error into a deterministic and a probabilistic part, where the latter results from Monte Carlo estimation of the expected value.
 Combining these ingredients, we then present novel full approximation error bounds for kernel EDMD for stochastic systems using Matérn and Wendland kernels.
@@
 Section 6 concludes and gives an outlook.
-Notation For a set $\Omega\subset\mathbb{R}^{n}$, we denote the Borel sigma algebra on $\Omega$ (w.r.t.
+Notation.
+For a set $\Omega\subset\mathbb{R}^{n}$, we denote the Borel sigma algebra on $\Omega$ (w.r.t.
 the standard topology) by $\mathcal{B}_{\Omega}$.
@@
 The identity operator on some space is denoted by $\mathcal{I}$, the identity matrix especially is denoted by $I$.
-For Banach spaces $\mathcal{S}$ and $\mathcal{T}$, $L(\mathcal{S},\mathcal{T})$ denotes the space of bounded linear operators from $\mathcal{S}$ to $\mathcal{T}$.
-We write $L^{p}(\Omega)$, $p\in[1,\infty]$, for the usual Lebesgue spaces with respect to Lebesgue measure and $C_{b}(\Omega)$ for the space of bounded and continuous functions on $\Omega$; occasionally, we just write $L^{p}$ or $C_{b}$ if the domain is clear from context.
+For Banach spaces $\mathcal{X}$ and $\mathcal{Y}$, $L(\mathcal{X},\mathcal{Y})$ denotes the space of bounded linear operators from $\mathcal{X}$ to $\mathcal{Y}$.
+The norm of $T\in L(\mathcal{X},\mathcal{Y})$ is denoted by $\|T\|_{\mathcal{X}\to\mathcal{Y}}$.
+We write $L^{p}(\Omega)$, $p\in[1,\infty]$, for the usual Lebesgue spaces with respect to Lebesgue measure, $C_{b}(\Omega)$ for the space of bounded and continuous functions, and $C_{c}^{\infty}(\Omega)$ for the space of infinitely differentiable and compactly supported functions, respectively, on $\Omega$; occasionally, we just write $L^{p}$, $C_{b}$, or $C_{c}^{\infty}$ if the domain is clear from context.
 For $f:\Omega\to\mathbb{R}$, we denote $\|f\|_{\infty}:=\sup_{x\in\Omega}|f(x)|$.
@@
 the $l$-th variable”, $l\in\{1,\dots,L\}$.
-For $d\in\mathbb{N}$, we abbreviate $\mathbbm{1}=\{v\in\mathbb{R}^{d}:|v_{i}|=1\}$.
 2. Koopman operator for stochastic systems: kernel EDMD and error analysis
@@
 Moreover, let $X$ and $Y$ be $\Omega$-valued random variables.
-It is well known (see, e.g., [Klen2008]) that there exists a function $\rho:\Omega\times\mathcal{B}_{\Omega}\to[0,1]$ such that
-  1. i)
+It is well known (see, e.g., [20]) that there exists a function $\rho:\Omega\times\mathcal{B}_{\Omega}\to[0,1]$ such that
+  1. (i)
     $\rho_{x}:=\rho(x,\cdot)$ is a probability measure for each $x\in\Omega$;
-  2. ii)
+  2. (ii)
     For each $\Delta\in\mathcal{B}_{\Omega}$, the function $\rho(\cdot,\Delta):\Omega\to[0,1]$, is measurable;
-  3. iii)
+  3. (iii)
     For each $\Delta\in\mathcal{B}_{\Omega}$: $\rho(x,\Delta)=\mathbb{P}(Y\in\Delta\,|\,X=x)$ for $\mathbb{P}^{X}$-a.e. $x\in\Omega$.
-A function with the first two properties is called a Markov transition kernel or Markov probability kernel; the third property makes $\rho$ a regular version of the conditional distribution of $Y$ given $X$ (see, e.g., [Kall2001]).
+A function with the first two properties is called a (Markov) transition kernel or probability kernel; the third property makes $\rho$ a regular version of the conditional distribution of $Y$ given $X$ (see, e.g., [19]).
 In what follows, we consider stochastic dynamical systems of the form
@@
 $(\mathcal{K}f)(x):=\mathbb{E}\big[f(Y)\,|\,X=x]=\int_{\Omega}f\,d\rho_{x},\quad x\in\Omega.$ (3)
-In the next chapter, we elaborate necessary and sufficient conditions for the invariance of Sobolev spaces $H^{s}(\Omega)$ under the Koopman operator.
-Those are of special interest since they arise as native spaces, or reproducing kernel Hilbert spaces, of certain kernels, for which there exists an extensive approximation framework, see Section 2.2, that allows for error bounds (Section 4).
+In the next section, we elaborate necessary and sufficient conditions for the invariance of Sobolev spaces $H^{s}(\Omega)$ under the Koopman operator.
+These spaces are of special interest since they arise as native spaces, or reproducing kernel Hilbert spaces (RKHS), of certain kernels for which there exists an extensive approximation framework, see Section 2.2, that allows for error bounds (Section 4).
 Before analyzing the Koopman operator on Sobolev spaces, we require the invariance of the space of square-integrable functions $L^{2}(\Omega)=H^{0}(\Omega)$ under $\mathcal{K}$.
@@
 If the condition
-$\exists C>0:\quad\int\rho_{x}(\Delta)\,dx\,\leq\,C|\Delta|\qquad\,\forall\Delta\in\mathcal{B}_{\Omega},$ ($L^{1}$)
+$\exists C>0:\quad\int\rho_{x}(\Delta)\,dx\,\leq\,C|\Delta|\qquad\,\forall\Delta\in\mathcal{B}_{\Omega},$ (4)
 holds, then $\mathcal{K}L^{2}(\Omega)\subset L^{2}(\Omega)$, and $\mathcal{K}:L^{2}(\Omega)\to L^{2}(\Omega)$ is a bounded linear operator.
@@
 ∎
-Note that, by the Fubini-Tonelli theorem, ($L^{1}$) implies $L^{1}$-invariance of the Koopman operator.
+Note that, by the Fubini-Tonelli theorem, (4) implies $L^{1}$-invariance of the Koopman operator.
 More generally, it can be shown that for a bounded domain $\Omega$, $L^{1}$-invariance (w.r.t.
@@
 While we consider $L^{2}(\Omega)=L^{2}_{{\boldsymbol{\lambda}}}(\Omega)$ with the Lebesgue measure $\boldsymbol{\lambda}$, our results can be extended to other reference measures – depending on the underlying stochastic system.
-In the case of $Y=X_{T}$ and $X=X_{0}$ for some Markov process $(X_{t})_{t\geq 0}$, the invariant measure $\mu$ of the process would be an intuitive choice, see, e.g., [NuskPeit23].
+In the case of $Y=X_{T}$ and $X=X_{0}$ for some Markov process $(X_{t})_{t\geq 0}$, the invariant measure $\mu$ of the process would be an intuitive choice, see, e.g., [36].
 2.2. Recap on deterministic systems: kernel EDMD and error bound
@@
 In particular, we may consider the SDE (2) with $\sigma\equiv 0$ giving rise to the ordinary differential equation $\dot{x}=b(x)$.
-In the deterministic case, the transition kernel is given by
-$\rho_{x}(A)=\delta_{F(x)}(A)$
-such that the Koopman identity (3) reduces to
+In this case, the transition kernel is given by $\rho_{x}(A)=\delta_{F(x)}(A)$, such that the Koopman identity (3) reduces to
 $\mathcal{K}\psi(x)=\psi(F(x)),$
 i.e., $\mathcal{K}\psi=\psi\circ F$.
-In this case, it is often assumed that $F:\Omega\to\Omega$ is a diffeomorphism with $\inf_{x\in\Omega}|\det F^{\prime}(x)|=c>0$ and $\sup_{x\in\Omega}\|F^{\prime}(x)\|<\infty$ [KohnPhil24].
-While the latter condition corresponds to the Lipschitz continuity of the drift function in (2), the assumption on the Jacobi determinant is a special case of ($L^{1}$).
-Specifically, applying the change of variables formula gives
+As in [24], we assume that $F:\Omega\to\Omega$ is a diffeomorphism satisfying $\inf_{x\in\Omega}|\det F^{\prime}(x)|=c>0$ and $\sup_{x\in\Omega}\|F^{\prime}(x)\|<\infty$ .
+While the latter condition corresponds to the Lipschitz continuity of the drift function in (2), the assumption on the Jacobi determinant is a special case of (4).
+Applying the change of variables formula gives
 $\int\delta_{F(x)}(\Delta)\,dx=\left|F^{-1}(\Delta)\right|=\int_{\Delta}|\det(F^{-1})^{\prime}(y)|\,dy\leq\frac{1}{c}\,|\Delta|,$
-i.e., ($L^{1}$) also holds in the deterministic case.
-We briefly recall Extended Dynamic Mode Decomposition (EDMD; [WillKevr15]) and its extension kernel EDMD, which are well-established numerical methods to approximate the Koopman operator $\mathcal{K}$ from data, see also [KutzBrun16] or the recent overview article [Colb24].
+i.e., (4) also holds in the deterministic case.
+We briefly recall Extended Dynamic Mode Decomposition (EDMD; [53]) and its extension kernel EDMD, which are well-established numerical methods to approximate the Koopman operator $\mathcal{K}$ from data, see also [28] or the recent overview article [13].
 EDMD requires two ingredients: First, let $d\in\mathbb{N}$ snapshots $(x_{i},y_{i})$ with $y_{i}=F(x_{i})$, $i\in\{1,\dots,d\}$, be given.
@@
 $\mathcal{X}=\{x_{1},\dots,x_{d}\}\quad\text{ and }\quad\mathcal{Y}=\{y_{1},\dots,y_{d}\}.$
-Second, let $M\in\mathbb{N}$ observable $L^{2}$-functions $\psi_{1},\dots,\psi_{M}$ be given, for which we use the notation $\Psi(x)=(\psi_{1}(x),\dots,\psi_{M}(x))^{\top}$ and define the data matrices
+Second, let $M\in\mathbb{N}$ observable $L^{2}$-functions $\psi_{1},\dots,\psi_{M}$ be given, for which we use the notation $\Psi(x)=(\psi_{1}(x),\dots,\psi_{M}(x))^{\top}$, and define the data matrices
 $\Psi_{\mathcal{X}}=(\Psi(x_{1}),\dots,\Psi(x_{d}))\in\mathbb{R}^{M\times d},\qquad\Psi_{\mathcal{Y}}=(\Psi(y_{1}),\dots,\Psi(y_{d}))\in\mathbb{R}^{M\times d}.$
-The subspace $\mathcal{V}=\mathrm{span}\{\psi_{1},\dots,\psi_{M}\}\subset L^{2}(\Omega)$ spanned by the observables is a finite-dimensional space. From the Koopman identity, we get
+The subspace $\mathcal{V}=\mathrm{span}\{\psi_{1},\dots,\psi_{M}\}\subset L^{2}(\Omega)$ spanned by the observables is finite-dimensional. From the Koopman identity, we get
 $\mathcal{K}\psi_{i}(x_{j})=\psi_{i}(F(x_{j}))=\psi_{i}(y_{j})\qquad\forall\,(i,j)\in\{1,\ldots,M\}\times\{1,\dots,d\},$ (5)
 that is, the Koopman operator maps the components of $\Psi_{\mathcal{X}}$ to the components of $\Psi_{\mathcal{Y}}$.
-Hence, to derive a data-driven approximation of the Koopman operator, restricted to $\mathcal{V}$, we seek for an approximation $K\in\mathbb{R}^{M\times M}$ that satisfies this identity approximately, that is,
+Since the Koopman operator acts linearly on observables, we seek a data-driven linear approximation on $\mathcal{V}$ given by a matrix $K\in\mathbb{R}^{M\times M}$ approximately satisfying (5), that is,
 $\Psi_{\mathcal{Y}}\approx K\Psi_{\mathcal{X}}.$ (6)
 Assuming that $\Psi_{\mathcal{X}}$ has full row rank (which is equivalent to strong linear independence of the observables $\psi_{1},\dots,\psi_{M}$ w.r.t.
-Lebesgue measure, see [PhilScha24, Lemma C.3]), an approximation can be computed as a solution to the least squares problem
+Lebesgue measure, see [39, Lemma C.3]), an approximation can be computed as a solution to the least squares problem
 $\min_{K\in\mathbb{R}^{M\times M}}\|K\Psi_{\mathcal{X}}-\Psi_{\mathcal{Y}}\|_{F}^{2}.$ (7)
@@
 As illustrated above, EDMD crucially depends on two ingredients: a finite number of data pairs $\{(x_{i},y_{i})\mid y_{i}=F(x_{i}),\,i\in\{1,\dots,d\}\}$ and the choice of the observables.
-Convergence in the infinite-data limit has been shown in [KordaMezi18] and further specified w.r.t.
-the estimation error (number of data pairs $d$) in [Mezi22, NuskPeit23, PhilScha24].
-The projection error is typically analyzed using finite-element observables as proposed in [ZhanZuaz23] and extended to control systems in [SchaWort23].
+Convergence in the infinite-data limit has been shown in [27] and further specified w.r.t.
+the estimation error (number of data pairs $d$) in [34, 36, 39].
+The projection error is typically analyzed using finite-element observables as proposed in [54] and extended to control systems in [46].
 However, the latter only yields $L^{2}$-error bounds.
-A particular variant of EDMD using a data-informed set of observables is kernel EDMD (kEDMD; [WillRowl15]), sharing two crucial features: First, from an application point of view, the choice of the observable functions is performed in a data-driven manner, eliminating fiddly and time-consuming tuning.
-Second, due to the rich mathematical structure of reproducing kernel Hilbert spaces, allows for pointwise deterministic error bounds.
+A particular variant of EDMD using a data-informed set of observables is kernel EDMD (kEDMD; [52]), sharing two crucial features: First, from an application point of view, the choice of the observable functions is performed in a data-driven manner, eliminating fiddly and time-consuming tuning.
+Second, the rich mathematical structure of reproducing kernel Hilbert spaces allows for pointwise deterministic error bounds.
 […]
 $\psi(x)=\langle\psi,\Phi_{x}\rangle_{\mathbb{H}}\quad\forall x\in\mathbb{R}^{n}$ (9)
-with the inner product $\langle\cdot,\cdot\rangle_{\mathbb{H}}$ holds.
-Then, the function $\Phi_{x}:\mathbb{R}^{n}\to\mathbb{R}$, $\Phi_{x}(y)=k(x,y)\in\mathbb{H}$, commonly termed the canonical feature of $k$ at $x\in\mathbb{R}^{n}$, plays a central role: We choose or sample points $\mathcal{X}=\{x_{1},\dots,x_{d}\}$ as before and let $y_{i}=F(x_{i})$.
-The dictionary is then given by the canonical features $\Phi_{x}=k(x,\cdot)$, $x\in\Omega$, of the data points, i.e.
-$\mathcal{V}=\mathcal{V}_{\mathcal{X}}:=\mathrm{span}(\mathcal{D})=\mathrm{span}(\{\Phi_{x_{1}},\dots,\Phi_{x_{d}}\}).$
-We immediately see that $M=d$ holds, i.e., we have exactly as many observables $\Phi_{x_{i}}$ as data points $x_{i}$.
-The entries in $\Psi_{X}\in\mathbb{R}^{d\times d}$ are easily calculated as $(\Psi_{X})_{i,j}=\psi_{i}(x_{j})=\Phi_{x_{i}}(x_{j})=k(x_{i},x_{j})$.
-The entries in $\Psi_{Y}$ are obtained via $(\Psi_{Y})_{i,j}=\psi_{i}(y_{j})=\Phi_{x_{i}}(F(x_{j}))=k(x_{i},F(x_{j}))$.
-Furthermore, we set
-$\mathbf{K}_{\mathcal{X}}:=\mathbf{K}_{\mathcal{X},\mathcal{X}}=(k(x_{i},x_{j}))_{i,j=1}^{d}=\Psi_{X},\quad\mathbf{K}_{\mathcal{X},F(\mathcal{X})}=(k(x_{i},F(x_{j}))_{i,j=1}^{d}=\Psi_{Y}$
-to which we will refer to as kernel matrices. Plugging these canonical features into (6) yields the regression problem
-$\min_{K\in\mathbb{R}^{d\times d}}\left\|K\mathbf{K}_{\mathcal{X}}-\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\right\|_{F}^{2}$ (9)
+holds, where $\Phi_{x}\in\mathbb{H}$, defined by $\Phi_{x}(y)=\text{\sf{k}}(x,y)$, is commonly termed the canonical feature of k at $x\in\Omega$.
+In kEDMD, these features play a central role: We choose or sample points $\mathcal{X}=\{x_{1},\dots,x_{d}\}$ as before and let $y_{i}=F(x_{i})$.
+The dictionary is given by the canonical features $\Phi_{x}=\text{\sf{k}}(x,\cdot)$, $x\in\Omega$, at the data points, i.e.
+$\mathcal{V}=\mathcal{V}_{\mathcal{X}}:=\mathrm{span}(\mathcal{D})=\mathrm{span}\{\Phi_{x_{1}},\dots,\Phi_{x_{d}}\}.$
+Clearly $M=d$ holds, i.e., we have exactly as many observables $\Phi_{x_{i}}$ as data points $x_{i}$.
+The data matrix entries are easily calculated as $(\Psi_{\mathcal{X}})_{i,j}=\psi_{i}(x_{j})=\Phi_{x_{i}}(x_{j})=\text{\sf{k}}(x_{i},x_{j})$ and $(\Psi_{\mathcal{Y}})_{i,j}=\psi_{i}(y_{j})=\Phi_{x_{i}}(F(x_{j}))=\text{\sf{k}}(x_{i},F(x_{j}))$.
+We write them as kernel matrices
+$\mathbf{K}_{\mathcal{X}}:=\mathbf{K}_{\mathcal{X},\mathcal{X}}=(\text{\sf{k}}(x_{i},x_{j}))_{i,j=1}^{d}=\Psi_{\mathcal{X}},\qquad\mathbf{K}_{\mathcal{X},F(\mathcal{X})}:=(\text{\sf{k}}(x_{i},F(x_{j}))_{i,j=1}^{d}=\Psi_{\mathcal{Y}}.$
+Plugging the kernel matrices into (7) yields the regression problem
+$\min_{K\in\mathbb{R}^{d\times d}}\left\|K\mathbf{K}_{\mathcal{X}}-\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\right\|_{F}^{2}.$ (10)
 Assuming that the data points are pairwise distinct, $\mathbf{K}_{\mathcal{X}}$ is symmetric positive definite (hence invertible) due to the symmetry and strict positive definiteness of the underlying kernel $\mathbf{k}$.
-[^2] We explicitly speak of strict positive definiteness above, since positive definiteness of kernels is often defined via positive semi-definiteness of kernel matrices in the literature, see [KohnPhil24].
+[^2] We explicitly speak of strict positive definiteness above, since positive definiteness of kernels is often defined via positive semi-definiteness of kernel matrices in the literature, see [24].
 The regression problem (10) now reduces to (exactly) solving a linear system, whose solution $K=\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\mathbf{K}_{\mathcal{X}}^{-1}$ corresponds to (8).
@@
 $S_{\mathcal{X}}g=\sum_{i=1}^{d}\left(\mathbf{K}_{\mathcal{X}}^{-1}g_{\mathcal{X}}\right)_{i}\Phi_{x_{i}}=g_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}$ (11)
-with $g_{\mathcal{X}}=(g(x_{1}),\dots,g(x_{d}))^{\top}$ and $\mathbf{k}_{\mathcal{X}}=(\Phi_{x_{1}},\dots,\Phi_{x_{d}})^{\top}$, see [KohnPhil24, Proposition 2.1].
-For $f\in\mathcal{V}_{\mathcal{X}}$, we have some $\alpha_{f}\in\mathbb{R}^{d}$ such that $f=\alpha_{f}^{\top}\mathbf{k}_{\mathcal{X}}$.
-Setting $g=\mathcal{K}|_{\mathcal{V}_{\mathcal{X}}}f$, we get $g_{\mathcal{X}}=\mathbf{K}_{\mathcal{X},F(\mathcal{X})}^{\top}\alpha_{f}$ and
+with $g_{\mathcal{X}}:=(g(x_{1}),\dots,g(x_{d}))^{\top}$ and $\mathbf{k}_{\mathcal{X}}:=(\Phi_{x_{1}},\dots,\Phi_{x_{d}})^{\top}$, see [24, Proposition 2.1].
+For $f\in\mathcal{V}_{\mathcal{X}}$, there exists $\alpha_{f}\in\mathbb{R}^{d}$ such that $f=\alpha_{f}^{\top}\mathbf{k}_{\mathcal{X}}$.
+With $g=\mathcal{K}|_{\mathcal{V}_{\mathcal{X}}}f$ we get $g_{\mathcal{X}}=\mathbf{K}_{\mathcal{X},F(\mathcal{X})}^{\top}\alpha_{f}$ and
 $S_{\mathcal{X}}\mathcal{K}|_{\mathcal{V}_{\mathcal{X}}}f=S_{\mathcal{X}}g=g_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}=\alpha_{f}^{\top}\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}.$ (12)
-This means the kEDMD-approximant $K=\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\mathbf{K}_{\mathcal{X}}^{-1}$ [KohnPhil24, Proposition 3.2] maps (the coefficients of) $f\in\mathcal{V}_{\mathcal{X}}$ to (the coefficients of) $S_{\mathcal{X}}\mathcal{K}|_{\mathcal{V}_{\mathcal{X}}}f$ linearly (in the respective canonical bases).
+This means the kEDMD-approximant $K=\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\mathbf{K}_{\mathcal{X}}^{-1}$ [24, Proposition 3.2] maps (the coefficients of) $f\in\mathcal{V}_{\mathcal{X}}$ to (the coefficients of) $S_{\mathcal{X}}\mathcal{K}|_{\mathcal{V}_{\mathcal{X}}}f$ linearly (in the respective canonical bases).
 To extend this operator to the entire space $\mathbb{H}$, one may simply add another orthogonal projection from the right, i.e., replace $f\in\mathcal{V}_{\mathcal{X}}$ by $S_{\mathcal{X}}g$ for any $g\in\mathbb{H}$.
@@
 $\widehat{\mathcal{K}}f=S_{\mathcal{X}}\mathcal{K}S_{\mathcal{X}}f=S_{\mathcal{X}}\mathcal{K}|_{\mathcal{V}_{\mathcal{X}}}(S_{\mathcal{X}}f)=\alpha_{S_{\mathcal{X}}f}^{\top}\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}.$ (14)
-The main result regarding this approximation obtained in [KohnPhil24, Theorem 3.4] may be summarized as follows.
+The main result regarding this deterministic Koopman approximation [24, Theorem 3.4] may be summarized as follows.
 Theorem 2.3.
@@
 First, it is assumed that the Koopman operator is indeed a bounded operator on $\mathbb{H}$.
-Due to closedness of the Koopman operator on RKHS, it suffices to show that $\mathcal{K}\mathbb{H}\subset\mathbb{H}$ (see [KohnPhil24, Lemma 3.1] and Proposition 2.4 below).
+Due to closedness of the Koopman operator on RKHS, it suffices to show that $\mathcal{K}\mathbb{H}\subset\mathbb{H}$ (see [24, Lemma 3.1] and Proposition 2.4 below).
   2. (2)
@@
     $\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}\leq Ch_{\mathcal{X}}^{r}\qquad\text{with}\qquad h_{\mathcal{X}}=\underset{z\in\Omega\;x\in\mathcal{X}}{\sup\min}\,\|z-x\|_{2}$ (15)
-    for a suitable rate $r>0$ and $C\geq 0$ using existing results from kernel-based approximation theory, see e.g. [Wend04, Section 11.2].
+    for a suitable rate $r>0$ and $C\geq 0$ using existing results from kernel-based approximation theory, see e.g. [50, Section 11.2].
 The invariance property (1) is a subtle detail that has to be verified rigorously and cannot be taken for granted.
-In particular, there are RKHS for which $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ implies that the underlying dynamics is affine linear, such as the Gaussian RKHS, see [GonzAbud23].
-For (fractional) Sobolev spaces $H^{s}$ (see Section 3.2 for an introduction), however, which are RKHS if $s>n/2$ and are generated by Wendland or Matérn kernels (see below for a precise definition), Koopman invariance was proven in [KohnPhil24, Theorem 4.2] in the deterministic setting.
+In particular, there are RKHS for which $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ implies that the underlying dynamics is affine linear, such as the Gaussian RKHS, see [15].
+For (fractional) Sobolev spaces $H^{s}$ (see Section 3.2 for an introduction), however, which are RKHS if $s>n/2$ and are generated by Wendland or Matérn kernels (see below for a precise definition), Koopman invariance was proven in [24, Theorem 4.2] in the deterministic setting.
 The central assumption is that the flow is smooth enough, such that the Koopman composition operator preserves Sobolev regularity.
-On the other hand, invariance of some RKHS under the Koopman operator is sufficient for boundedness, as shown in [KohnPhil24, Lemma 3.1] for the deterministic case, and analogously in Proposition 2.4 below for the stochastic case.
+On the other hand, invariance of a RKHS under the Koopman operator is sufficient for boundedness, as shown in [24, Lemma 3.1] for the deterministic case, and analogously in Proposition 2.4 below for the stochastic case.
 Proposition 2.4.
@@
 Let $f_{n},f,g\in\mathbb{H}$ such that $f_{n}\to f$ and $\mathcal{K}f_{n}\to g$ in $\mathbb{H}$.
-Recall the canonical features $\Phi_{x}$ for $x\in\Omega$.
+Recall the canonical features $\Phi_{x}$ at $x\in\Omega$.
 Then, the reproducing property (9) yields
-$g(x)$ $=\langle g,\Phi_{x}\rangle_{\mathbb{H}}=\lim_{n\to\infty}\langle\mathcal{K}f_{n},\Phi_{x}\rangle_{\mathbb{H}}=\lim_{n\to\infty}\mathcal{K}f_{n}(x)=\lim_{n\to\infty}\int f_{n}(y)\,d\rho_{x}(y).$
+$g(x)$ $=\langle g,\Phi_{x}\rangle_{\mathbb{H}}=\lim_{n\to\infty}\langle\mathcal{K}f_{n},\Phi_{x}\rangle_{\mathbb{H}}=\lim_{n\to\infty}\mathcal{K}f_{n}(x)=\lim_{n\to\infty}\int_{\Omega}f_{n}(y)\,d\rho_{x}(y).$
 Now, $f_{n}(y)=\langle f_{n},\Phi_{y}\rangle_{\mathbb{H}}\to\langle f,\Phi_{y}\rangle_{\mathbb{H}}=f(y)$ for all $y\in\Omega$ as $n\to\infty$ by continuity of the scalar product, and $|\langle f_{n},\Phi_{y}\rangle_{\mathbb{H}}|\leq\|f_{n}\|_{\mathbb{H}}\|\Phi_{y}\|_{\mathbb{H}}\leq C$ for all $n\in\mathbb{N}$.
@@
 ∎
-To sum up, the assumption of Theorem 2.3 is met for Sobolev RKHS, and error bounds of the projection error (14) are readily available from the literature for the generating kernels, leading to a pointwise and deterministic error bound for EDMD for deterministic systems.
+To sum up, the assumption of Theorem 2.3 is met for Sobolev RKHS, and bounds on the projection error in the form (15) are well-known for many generating kernels, leading to a pointwise and deterministic error bound for deterministic systems.
 2.3. Kernel EDMD for stochastic systems
-In contrast to the deterministic case, the stochastic Koopman operator is defined by means of conditional expectations.
-These are usually not readily available in applications with unknown dynamics such that the analogue to the regression problem (6) is not directly solvable if only observations are available.
-More precisely, to implement the counterpart of the approximate identity (5), i.e.,
-$K\Psi_{\mathcal{X}}\approx\Psi_{\mathcal{Y}}=(\mathbb{E}[\Psi_{i}(Y)\mid X=x_{j}])_{(i,j)\in[M]\times[d]},$ (15)
-or the corresponding regression problem (6), the data matrix $\Psi_{\mathcal{Y}}$ has to be further approximated.
-To this end, we build upon the rich theory of Monte Carlo methods to perform an entry-wise approximation, where we draw $m_{j}\in\mathbb{N}$ i.i.d.
+In constrast, the stochastic Koopman operator propagates conditional expectations of observables. Here, the counterpart of the approximate identity (6) takes the form
+$K\Psi_{\mathcal{X}}\approx\Psi_{\mathcal{Y}}=\big(\mathbb{E}[\Psi_{i}(Y)\mid X=x_{j}]\big)_{i,j=1}^{M,d}.$ (16)
+Consequently, for the finite-dimensional approximation $K\in\mathbb{R}^{M\times M}$ in (16), setting
+$Z^{+}:=\Psi(Y)\approx KZ,\qquad Z:=\Psi(X),$
+defines a linear surrogate system in feature space.
+The surrogate system describes the evolution of the conditional mean of the lifted observables, i.e., $\mathbb{E}[Z^{+}\,|\,X]\approx KZ=K\Psi(X)$.
+Here, probability enters through the conditional expectation: the surrogate system approximates the expected evolution of the lifted observables, while stochastic fluctuations around this mean remain.
+However, the conditional expectations in (16) are usually not readily available in applications with unknown dynamics, such that the analogue to the regression problem (7) cannot be directly solved.
+Instead, the data matrix $\Psi_{\mathcal{Y}}$ has to be approximated from observations of the stochastic system.
+For this, we build upon the rich theory of Monte Carlo methods to perform an entry-wise approximation.
+We draw $m_{j}\in\mathbb{N}$ i.i.d.
 samples $y_{j}^{(l)}\sim\rho_{x_{j}}$, $l=1,\dots,m_{j}$, and invoke the law of large numbers such that
 $\mathbb{E}[\Psi_{i}(Y)\mid X=x_{j}]\approx\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}\Psi_{i}(y_{j}^{(l)}).$ (17)
-For the kernel case, we then have, in view of $\Psi_{i}=k(x_{i},\cdot)$,
-$K\mathbf{K}_{\mathcal{X}}\approx\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\qquad\mathrm{with}\qquad(\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}})_{ij}=\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}k(x_{i},y_{j}^{(l)})$ (17)
+For the kernel case, we then have, in view of $\Psi_{i}=\text{\sf{k}}(x_{i},\cdot)$,
+$K\mathbf{K}_{\mathcal{X}}\approx\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\qquad\mathrm{with}\qquad(\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}})_{ij}=\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}\text{\sf{k}}(x_{i},y_{j}^{(l)})$ (18)
 This yields the matrix approximant $K=\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\mathbf{K}_{\mathcal{X}}^{-1}$, which corresponds to the operator approximant
-$\widehat{\mathcal{K}}^{\mathrm{MC}}f=f_{\mathcal{X}}^{\top}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{k}_{\mathcal{X}}$ (18)
-which mirrors the deterministic approximant from (13).
+$\widehat{\mathcal{K}}^{\mathrm{MC}}f=f_{\mathcal{X}}^{\top}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{k}_{\mathcal{X}},$ (19)
+mirroring the deterministic approximant from (14).
 However, the entry-wise Monte Carlo approximation $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ introduces a further estimation error since (17) only holds approximately.
@@
 Remark 2.5.
-Another possible design choice more in line with the covariance operator approach circumvents the estimation of estimating the expected value under certain assumptions on the underlying stochastic system and sampling scheme (see [NuskPeit23, PhilSchaWort2025] for details).
+Another possible design choice more in line with the covariance operator approach circumvents the estimation of the expected value in (17) under certain assumptions on the underlying stochastic system and sampling scheme (see [36, 41] for details).
 In the setting of a governing stochastic differential equation, the authors either sample the design points in $\mathcal{X}$ from the invariant measure $\mu$ or assume ergodicity and sample from a single long trajectory $x_{0},x_{1},\dots,x_{d}$.
-In this paper, we pursue the first approach.
-Future work utilizing the second approach could be the subject of further work and will be discussed later.
+In this paper, we pursue the Monte Carlo approach illustrated above, which requires weaker assumptions on the system but relies on i.i.d.
+samples.
 2.4. Main contribution: error analysis of kEDMD for stochastic systems
@@
 Let $\mathbb{H}$ be an RKHS of functions over $\Omega$, and assume $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b}(\Omega),C_{b}(\Omega))$.
-Furthermore, for $\mathcal{X}=\{x_{1},\ldots,x_{d}\}$, let $m_{j}=m\in\mathbb{N}$ i.i.d.
-samples from every $\rho_{x_{j}}$, $j\in\{1,\dots,d\}$ be given.
-Then, for every $\varepsilon>0$ and for all $f\in\mathbb{H}$, $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
-$\|(\mathcal{K}-\widehat{\mathcal{K}}^{\mathrm{MC}})f\|_{\infty}\leq\big[1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\big]\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}+\varepsilon\cdot\|k\|_{\infty}^{1/2}\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$ (19)
-with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right)$.
-The proof of the above theorem consists of two parts.
+Furthermore, for $\mathcal{X}=\{x_{1},\ldots,x_{d}\}\subset\Omega$, let $m_{j}=m\in\mathbb{N}$ i.i.d.
+samples from every $\rho_{x_{j}}$, $j\in\{1,\dots,d\}$, be given.
+Then, for every $\varepsilon>0$ satisfying $m\geq\frac{2\|\text{\sf{k}}\|_{\infty}}{\varepsilon^{2}}\log(2d)$ and for all $f\in\mathbb{H}$, $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
+$\big\|\big(\mathcal{K}-\widehat{\mathcal{K}}^{\mathrm{MC}}\big)f\big\|_{\infty}\leq\big[1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\big]\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}+\varepsilon\cdot\|\text{\sf{k}}\|_{\infty}^{1/2}\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$ (20)
+with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right)$, where $\mathbbm{1}=\{v\in\mathbb{R}^{d}:|v_{i}|=1\}$.
+The proof of the above theorem (cf.
+Section 4.3) consists of two parts.
 First, the finite number of data points and observables given by $\mathcal{X}\subset\Omega$ introduces a deterministic error.
@@
 $\|\mathcal{K}-\widehat{\mathcal{K}}^{\mathrm{MC}}\|_{\mathbb{H}\to C_{b}}\leq\underbrace{\|\mathcal{K}-\widehat{\mathcal{K}}\|_{\mathbb{H}\to C_{b}}}_{\text{deterministic error}}+\underbrace{\|\widehat{\mathcal{K}}-\widehat{\mathcal{K}}^{\mathrm{MC}}\|_{\mathbb{H}\to C_{b}}}_{\text{probabilistic error}}.$
-The probabilistic error is controlled by the accuracy of the entry-wise Monte Carlo estimator (5), i.e., the number $m$ of i.i.d.
-samples from $\rho_{x_{j}}$.
+In Theorem 2.6, the parameter $m$ is chosen sufficiently large (depending on $\varepsilon$ and $d$) to ensure that the stated lower bound on the probability is nonnegative.
+More precisely, for fixed $d$ and $\varepsilon>0$, choosing $m\geq\frac{2\|\text{\sf{k}}\|_{\infty}}{\varepsilon^{2}}\log(2d)$ guarantees that the probability bound is meaningful.
+In practice, the parameters $d$ and $m$ play different roles: the number of design points $d$ controls the deterministic approximation error, while $m$ determines the accuracy of the Monte Carlo estimation.
+For a prescribed accuracy level $\varepsilon$, one therefore chooses $d$ large enough to control the deterministic error and then selects $m=m(d,\varepsilon)$ sufficiently large to achieve the desired error bound with high probability.
+We refer the reader to Section 5.1 and Appendix D for further discussion on the interplay between $d$, $\varepsilon$, and $m$.
+The assumption $\mathcal{K}\in L(C_{b}(\Omega),C_{b}(\Omega))$ is a regularity condition on the transition kernel $\rho_{x}$, which is characterized and discussed in Section 4.1, and is independent of the continuity of the RKHS kernel k defining $\mathbb{H}$.
 Besides, the two main components in Theorem 2.6 are the operator norms (which are assumed to be finite) and the projection error governed by $\mathcal{I}-S_{\mathcal{X}}$.
@@
 Boundedness of the Koopman operator.
-Invariance of reproducing kernel Hilbert spaces for Koopman operators is a non-trivial issue.
+Invariance of reproducing kernel Hilbert spaces under Koopman operators is a non-trivial issue.
 Hence, in Section 3 we provide characterizations and sufficient conditions for the invariance of Sobolev spaces.
-Here, we present the most important and practically relevant result.
-Its proof follows from the more general result of Theorem 3.2.
+Here, we present the most important and practically relevant result, which follows from the more general result of Theorem 3.2.
 Theorem 2.7.
-If the transition kernel $\rho_{x}$ is given by a density $p:\Omega\times\Omega\to[0,\infty)$, i.e.,
-$\rho_{x}(\Delta)=\int_{\Delta}p(x,y)\,dy\qquad\forall x\in\Omega,\,\Delta\in\mathcal{B}_{\Omega},$
-that satisfies
+If the transition kernel $\rho_{x}$ is given by a density $p:\Omega\times\Omega\to[0,\infty)$, i.e., $\rho_{x}(\Delta)=\int_{\Delta}p(x,y)\,dy$ for all $x\in\Omega$ and $\Delta\in\mathcal{B}_{\Omega}$, that satisfies
 $\sum_{|\alpha|\leq\ell}\left\|D_{1}^{\alpha}p(\cdot,\cdot)\right\|_{L^{2}(\Omega\times\Omega)}\leq C$ (21)
-for some constant $C>0$ and $\ell\in\mathbb{N}$, then $\mathcal{K}\in L(H^{s},H^{s})$ for all $s\in(0,\ell]$.
-We point out that Koopman invariance, as shown in [KohnPhil24] for suitably chosen reproducing kernel Hilbert spaces, was extended to stochastic systems in the preprint [OlguOsseRami2025], under the stronger regularity assumption $p\in C^{m}(\Omega\times\Omega)$ for the transition density.
+for some constant $C>0$ and $\ell\in\mathbb{N}$, then $\mathcal{K}\in L(H^{s}(\Omega),H^{s}(\Omega))$ for all $s\in(0,\ell]$.
+We point out that Koopman invariance, as shown in [24] for suitably chosen reproducing kernel Hilbert spaces, was extended to stochastic systems in the preprint [38], under the stronger regularity assumption $p\in C^{m}(\Omega\times\Omega)$ for the transition density.
 On the one hand, we generalize this result by significantly weakening the assumptions.
@@
 Projection error bounds in terms of the fill distance.
-In this work, we apply our results to two popular choices of kernels, Wendland kernels [Wend04] and Matérn kernels [BerlThom2004] which are in particular consistent with the above mentioned invariance as they generate Sobolev spaces (see [Adams2003] and Section 3.2 for an introduction).
+In this work, we apply our results to two popular choices of kernels, Wendland kernels [50] and Matérn kernels [6] which are in particular consistent with the above mentioned invariance as they generate Sobolev spaces (see [1] and Section 3.2 for an introduction).
 We briefly present the error bounds for Wendland kernels, as the bounds for Matérn kernels (see Appendix C for an introduction) are structurally very similar.
-The Wendland kernel $k_{n,l}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ is bounded, continuous and strictly positive-definite (see [Wend04] and Appendix B for details).
+The Wendland kernel $\text{\sf{k}}_{n,l}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ is bounded, continuous and strictly positive-definite (see [50] and Appendix B for details).
 For $\sigma_{n,l}=\frac{n+1}{2}+l$, it generates the (fractional) Sobolev space $H^{\sigma_{n,l}}(\mathbb{R}^{n})$ as RKHS with equivalence of the standard Sobolev norm and the induced (native) RKHS norm.
-Restricted to a bounded domain $\Omega\subset\mathbb{R}^{n}$ with Lipschitz boundary, as is of interest in the setting of this paper, the Wendland kernel generates the corresponding Sobolev space $H^{\sigma_{n,l}}(\Omega)$ [KohnPhil24, Theorem 4.1].
+Restricted to a bounded domain $\Omega\subset\mathbb{R}^{n}$ with Lipschitz boundary, as is of interest in the setting of this paper, the Wendland kernel generates the corresponding Sobolev space $H^{\sigma_{n,l}}(\Omega)$ [24, Theorem 4.1].
 Results from kernel approximation theory allow for a more explicit bound on the projection error in terms of the fill distance.
 Theorem 2.8 ($L^{\infty}$-error bound for Wendland native spaces).
-Let $l\in\mathbb{N}$, define $\sigma_{n,l}=\frac{n+1}{2}+l$ and denote by $\mathbb{H}$ the Sobolev space $H^{\sigma_{n,l}}(\Omega)$ induced by the Wendland kernel $k_{n,l}$ as RKHS.
+Let $l\in\mathbb{N}$, define $\sigma_{n,l}=\frac{n+1}{2}+l$ and denote by $\mathbb{H}$ the Sobolev space $H^{\sigma_{n,l}}(\Omega)$ induced by the Wendland kernel $\text{\sf{k}}_{n,l}$ as RKHS.
 Assume that $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b}(\Omega),C_{b}(\Omega))$.
-Then, for every $\varepsilon>0$ and all $f\in\mathbb{H}$ with $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
-$\|(\mathcal{K}-\mathcal{K}^{\mathrm{MC}})f\|_{\infty}\leq$ $C_{1}\,(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot h_{\mathcal{X}}^{l+1/2}+\varepsilon\cdot C_{2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$ (21)
-with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2C_{2}^{2}}\right)$, with constants $C_{1},C_{2}>0$ only depending on the kernel $k_{n,l}$.
+Furthermore, for $\mathcal{X}=\{x_{1},\ldots,x_{d}\}\subset\Omega$, let $m_{j}=m\in\mathbb{N}$ i.i.d.
+samples from every $\rho_{x_{j}}$, $j\in\{1,\dots,d\}$, be given.
+Then, for every $\varepsilon>0$ satisfying $m\geq\frac{2C_{2}^{2}}{\varepsilon^{2}}\log(2d)$ and all $f\in\mathbb{H}$ with $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
+$\left\|(\mathcal{K}-\mathcal{K}^{\mathrm{MC}})f\right\|_{\infty}\leq$ $C_{1}\,(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot h_{\mathcal{X}}^{l+1/2}+\varepsilon\cdot C_{2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$ (22)
+with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2C_{2}^{2}}\right)$, where $C_{1},C_{2}>0$ are constants only depending on the kernel $\text{\sf{k}}_{n,l}$, and $\mathbbm{1}=\{v\in\mathbb{R}^{d}:|v_{i}|=1\}$.
 Note that the bounds of Theorems 2.6 and 2.8 depend on the smallest eigenvalue of the kernel matrix $\mathbf{K}_{\mathcal{X}}$.
-Tikhonov regularization is a means to control the latter, for the price of introducing a deterministic bias Section 4 addresses the impact of regularization on the deterministic (Proposition 4.5), probabilistic (Corollary 4.9) and total error bounds (Corollary 4.12).
+Tikhonov regularization is a means to control the latter, for the price of introducing a deterministic bias.
+Section 4 addresses the impact of regularization on the deterministic (Proposition 4.5), probabilistic (Corollary 4.9), and total error bounds (Corollary 4.12).
 3. $H^{s}$-invariance under the Koopman operator
-Since Wendland (and Matérn) kernels generate the Sobolev spaces as their respective native spaces, we study the invariance of Sobolev spaces $H^{s}$ under the stochastic Koopman operator.
+Since Wendland (and Matérn) kernels generate the Sobolev spaces $H^{s}(\Omega)$ as their respective native spaces, we study their invariance under the stochastic Koopman operator.
 As those spaces are defined via partial integration, a condition for preserving weak derivatives under Koopman operator action is paramount.
@@
 For $g\in L^{\infty}(\Omega)$, define $\nu_{g}:\mathcal{B}_{\Omega}\to\mathbb{R}$ by
-$\nu_{g}(\Delta):=\int_{\Omega}\rho_{x}(\Delta)g(x)\,dx.$ (22)
+$\nu_{g}(\Delta):=\int_{\Omega}\rho_{x}(\Delta)g(x)\,dx,\qquad\Delta\in\mathcal{B}_{\Omega}.$ (23)
 The following lemma links the signed measure $\nu_{g}$ to the action of the dual of the Koopman operator on $L^{1}(\Omega)$ and $L^{2}(\Omega)$. Its proof is given in Section 3.1.
 Lemma 3.1.
-Assume that ($L^{1}$) holds and consider $\mathcal{K}_{1}:L^{1}(\Omega)\to L^{1}(\Omega)$ as well as $\mathcal{K}_{2}:L^{2}(\Omega)\to L^{2}(\Omega)$.
+Assume that (4) holds and consider $\mathcal{K}_{1}:L^{1}(\Omega)\to L^{1}(\Omega)$ as well as $\mathcal{K}_{2}:L^{2}(\Omega)\to L^{2}(\Omega)$.
 Then, the measure $\nu_{g}$ defined in (23) is absolutely continuous w.r.t.
 the Lebesgue measure, and its density $\frac{d\nu_{g}}{dx}\in L^{\infty}(\Omega)$ satisfies
-$\mathcal{K}_{1}^{*}g=\mathcal{K}_{2}^{*}g=\frac{d\nu_{g}}{dx},\quad\mathrm{for\ all}\quad g\in L^{\infty}(\Omega).$
+$\mathcal{K}_{1}^{*}g=\mathcal{K}_{2}^{*}g=\frac{d\nu_{g}}{dx}\quad\mathrm{for\ all}\quad g\in L^{\infty}(\Omega),$
 where $\mathcal{K}_{1}^{\ast}$ and $\mathcal{K}_{2}^{\ast}$ denote the adjoints w.r.t. the respective dual pairing.
@@
 Theorem 3.2 (Koopman invariance of Sobolev spaces).
-Assume that the condition ($L^{1}$) holds.
+Assume that (4) holds.
   • (i)
-    Let $\ell\in\mathbb{N}_{\geq 1}$. Then $H^{\ell}$ is invariant under $\mathcal{K}$ if and only if there is $C\geq 0$ such that
+    Let $\ell\in\mathbb{N}_{\geq 1}$. Then $H^{\ell}(\Omega)$ is invariant under $\mathcal{K}$ if and only if there is $C\geq 0$ such that
     $\sum_{|\alpha|\leq\ell}\left\|\frac{d\nu_{D^{\alpha}\varphi}}{dx}\right\|_{H^{-|\alpha|}}\leq C\|\varphi\|_{L^{2}}\qquad\forall\varphi\in C_{c}^{\infty}.$ (24)
   • (ii)
-    If (23) holds, then $H^{s}$ is invariant under $\mathcal{K}$ for all $s\in(0,\ell]$.
+    If (24) holds, then $H^{s}(\Omega)$ is invariant under $\mathcal{K}$ for all $s\in(0,\ell]$.
   • (iii)
@@
     $\rho_{x}(\Delta)=\int_{\Delta}p(x,y)\,dy\quad\mathrm{for\ all}\quad x\in\Omega,\,\Delta\in\mathcal{B}_{\Omega}.$
-    If for some constant $C>0$ and $\ell\in\mathbb{N}_{\geq 1}$
-    $\sum_{|\alpha|\leq\ell}\left\|D_{1}^{\alpha}\,p(\cdot,\cdot)\right\|_{L^{2}(\Omega\times\Omega)}\leq C$ (24)
-    then $H^{s}$ is invariant under $\mathcal{K}$ for all $s\in(0,\ell]$.
+    If for some constant $C>0$ and $\ell\in\mathbb{N}_{\geq 1}$, we have
+    $\sum_{|\alpha|\leq\ell}\left\|D_{1}^{\alpha}\,p(\cdot,\cdot)\right\|_{L^{2}(\Omega\times\Omega)}\leq C,$ (25)
+    then $H^{s}(\Omega)$ is invariant under $\mathcal{K}$ for all $s\in(0,\ell]$.
 Note that part (iii) of Theorem 3.2 is precisely the statement of Theorem 2.7.
@@
 Proposition 3.3 (Truncated Normal Distribution).
-For $n=1$ and $\Omega=(0,1)\subset\mathbb{R}$, assume $Y\mid\{X=x\}\sim\rho_{x}:=\mathcal{N}_{(0,1)}(x,1)$ where $\mathcal{N}_{(0,1)}(x,1)$ denotes the (doubly) truncated normal distribution on $(0,1)$ with location parameter $x\in(0,1)$ and scale parameter set to $1$.
+For $n=1$ and $\Omega=(0,1)\subset\mathbb{R}$, assume $Y\mid\{X=x\}\sim\rho_{x}:=\mathcal{N}_{(0,1)}(x,1)$, where $\mathcal{N}_{(0,1)}(x,1)$ denotes the (doubly) truncated normal distribution on $(0,1)$ with location parameter $x\in(0,1)$ and scale parameter set to $1$.
 Then, $H^{1}(\Omega)$ is invariant under the associated Koopman operator that acts on $f\in H^{1}(\Omega)$ via $f\in H^{1}\bigl((0,1)\bigr)$ via $\mathcal{K}f(x)=\mathbb{E}[f(Y)\mid X=x]$.
 Proof.
-The doubly truncated normal distribution $\rho_{x}$ is the distribution of a random variable $Z\sim\mathcal{N}(x,1)$ conditioned on the event $0<Z<1$.
-To apply part (iii) of Theorem 3.2, we study.
-the density of $\rho_{x}$.
+The doubly truncated normal distribution $\rho_{x}=\mathcal{N}_{(0,1)}(x,1)$ is defined as the distribution of a random variable $Z\sim\mathcal{N}(x,1)$ conditioned on the event $0<Z<1$.
+To apply part (iii) of Theorem 3.2, we study the density of the transition kernel $\rho_{x}$.
 Denote the cumulative distribution function of the standard normal distribution by $F(\xi)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\xi}e^{-t^{2}/2}\,dt$.
@@
 $p(x,y)=\frac{\frac{1}{\sqrt{2\pi}}e^{-(y-x)^{2}/2}}{F(1-x)-F(-x)}=\frac{G(x,y)}{H(x)}$
-Numerator and denominator both depend smoothly on $x$, so we can easily compute derivatives with respect to $x$ and bound them. Using $F^{\prime}(\xi)=\frac{1}{\sqrt{2\pi}}e^{-\xi^{2}/2}$, we get
-$\partial_{1}p(x,y)=\frac{\partial_{1}G(x,y)H(x)-G(x,y)H^{\prime}(x)}{H(x)^{2}}=G(x,y)\left((y-x)-\frac{H^{\prime}(x)}{H(x)^{2}}\right)$
+The numerator and denominator both depend smoothly on $x$. With $F^{\prime}(\xi)=\frac{1}{\sqrt{2\pi}}e^{-\xi^{2}/2}$, we compute the derivative with respect to $x$:
+$\partial_{1}p(x,y)=\frac{\partial_{1}G(x,y)H(x)-G(x,y)H^{\prime}(x)}{H(x)^{2}}=G(x,y)\left((y-x)-\frac{H^{\prime}(x)}{H(x)^{2}}\right).$
 Note that $0<C_{1}\leq H(x)\leq C_{2}$ since $H(x)=\mathbb{P}(U\in(-x,1-x))$ for some $U\sim\mathcal{N}(0,1)$ and $x\in(0,1)$.
-Using the fact that $G(x,y)\leq\frac{1}{\sqrt{2\pi}}$ for all $x,y\in(0,1)$ and $|H^{\prime}(x)|\leq\frac{1}{\sqrt{2\pi}}$, we obtain the basic bound
+With $G(x,y)\leq\frac{1}{\sqrt{2\pi}}$ for all $x,y\in(0,1)$ and $|H^{\prime}(x)|\leq\frac{1}{\sqrt{2\pi}}$, we obtain the bound
 $|\partial_{1}p(x,y)|\leq\frac{1}{\sqrt{2\pi}}\left(1+\frac{1}{\sqrt{2\pi}C_{1}}\right)$
@@
 3.2. Sobolev Spaces
-For self-containedness, we briefly recapitulate the notion of Sobolev spaces following [Adams2003].
-Readers who are familiar with Sobolev spaces may only skim this section for notation or skip it entirely.
-Recall the bounded domain $\Omega\subset\mathbb{R}^{n}$ with Lipschitz boundary and write $C_{c}^{\infty}(\Omega)$ for the space of infinitely differentiable and compactly supported functions on $\Omega$.
-If the domain is clear from context, we might write $C_{c}^{\infty}=C_{c}^{\infty}(\Omega)$.
-Given some multi-index $\alpha\in\mathbb{N}_{0}^{n}$, we write $D^{\alpha}\varphi$ for the $\alpha$-th partial derivative of a sufficiently smooth function $\varphi$.
-Note that $\varphi\in C_{c}^{\infty}(\Omega)$ implies $D^{\alpha}\varphi(x)=0$ for all $x\in\partial\Omega$ and $\alpha\in\mathbb{N}^{n}$ since $\varphi\equiv 0$ in an open neighborhood of the boundary.
-Next, recall the notion of the weak derivative of a function: Consider a locally integrable function $f\in L_{\mathrm{loc}}^{1}(\Omega)$; that is, $f$ is integrable on every compact set $K\subset\Omega$, for which $f\in L^{2}(\Omega)$ is clearly sufficient.
+For self-containedness, we briefly recapitulate the notion of Sobolev spaces following [1]. Readers who are familiar with Sobolev spaces may only skim this section for notation or skip it entirely.
+Consider the bounded domain $\Omega\subset\mathbb{R}^{n}$ with Lipschitz boundary and the space $C_{c}^{\infty}(\Omega)$ of infinitely differentiable and compactly supported functions on $\Omega$.
+We recall the notion of the weak derivative of a function: Consider a locally integrable function $f\in L_{\mathrm{loc}}^{1}(\Omega)$; that is, $f$ is integrable on every compact set $K\subset\Omega$, for which $f\in L^{2}(\Omega)$ is clearly sufficient.
 If there exists a function $g\in L_{\mathrm{loc}}^{1}(\Omega)$ satisfying
 $\int_{\Omega}f(x)\,(D^{\alpha}\varphi)(x)\,dx=(-1)^{|\alpha|}\int_{\Omega}g(x)\,\varphi(x)\,dx$ (26)
-for every test function $\varphi\in C_{c}^{\infty}(\Omega)$, we call $g$ a $\alpha$-th weak derivative of $f$.
-The weak derivative is unique up to a set of measure zero, i.e., two weak derivatives of $f$ agree almost everywhere, and needs not be defined pointwise (the elements of $L_{\mathrm{loc}}^{1}(\Omega)$ are equivalence classes, after all).
-In the case of a “conventionally” partially differentiable function, the weak derivative coincides (almost everywhere) with the respective partial derivative.
-Hence, in a slight abuse of notation, we write $g=D^{\alpha}f$ for the weak derivative.
+for every test function $\varphi\in C_{c}^{\infty}(\Omega)$, we call $g$ the $\alpha$-th weak derivative of $f$ and write $g=D^{\alpha}f$.
 Using the notion of a weak derivative, given some $\ell\in\mathbb{N}$ we introduce
@@
 There are two predominant definitions of fractional Sobolev spaces in the literature, either via the Fourier transform or via weak derivatives and interpolation between integer-order Sobolev spaces.
-For a domain with Lipschitz boundary, as is $\Omega\subset\mathbb{R}^{n}$, those definitions coincide [McLe2000, Theorem 3.30], so we proceed with the more suitable second definition for the remainder of this paper.
+For a domain with Lipschitz boundary, as is $\Omega\subset\mathbb{R}^{n}$, those definitions coincide by [31, Theorem 3.30], so we proceed with the second definition which is more suitable for the remainder of this paper.
 Given $s\geq 0$, write $s=\lfloor s\rfloor+r$ with $r\in[0,1)$ and define the norm
 $\|f\|_{H^{s}(\Omega)}^{2}:=\sum_{|\alpha|\leq\,s}\|D^{\alpha}f\|_{L^{2}(\Omega)}^{2}+\chi_{(0,1)}(r)\int_{\Omega}\int_{\Omega}\frac{\left|D^{\alpha}f(x)-D^{\alpha}f(y)\right|^{2}}{\|x-y\|_{2}^{n+2r}}\,dx\,dy.$ (27)
-Under this definition, the Sobolev space $H^{s}(\Omega)=\{f\in L^{2}(\Omega):\|f\|_{H^{s}(\Omega)}<\infty\}$ presents itself as the interpolation space between $H^{\lfloor s\rfloor}(\Omega)$ and $H^{\lceil s\rceil}(\Omega)$.
+Under this definition, the Sobolev space $H^{s}(\Omega)=\{f\in L^{2}(\Omega):\|f\|_{H^{s}(\Omega)}<\infty\}$ is an interpolation space between $H^{\lfloor s\rfloor}(\Omega)$ and $H^{\lceil s\rceil}(\Omega)$.
+More precisely, $H^{s}(\Omega)=[H^{\lfloor s\rfloor}(\Omega),H^{\lceil s\rceil}(\Omega)]_{r}$.
 3.3. Proof of Lemma 3.1 and Theorem 3.2
 We start by proving the important auxiliary result from the beginning of this section.
-While Lemma 3.1 essentially follows from [PhilScha24, Appendix B], we provide a concise proof below for self-containedness.
+While Lemma 3.1 essentially follows from [39, Appendix B], we provide a concise proof below for self-containedness.
 Proof of Lemma 3.1.
-Let $g\in L^{\infty}(\Omega)$ and $\Delta\in\mathcal{B}_{\Omega}$. If ($L^{1}$) holds, the definition of (22) immediately gives
-$|\nu_{g}(\Delta)|\,\leq\,\|g\|_{L^{\infty}}\int\rho_{x}(\Delta)\,dx\,\leq\,C\|g\|_{L^{\infty}}|\Delta|.$
+Let $g\in L^{\infty}(\Omega)$ and $\Delta\in\mathcal{B}_{\Omega}$. If (4) holds, the definition of (23) immediately gives
+$|\nu_{g}(\Delta)|\,\leq\,\|g\|_{L^{\infty}}\int_{\Omega}\rho_{x}(\Delta)\,dx\,\leq\,C\|g\|_{L^{\infty}}|\Delta|.$
 Hence, $\nu_{g}$ is absolutely continuous w.r.t.
@@
 Moreover, for $f\in L^{1}(\Omega)$ and the corresponding Koopman operator $\mathcal{K}_{1}:L^{1}(\Omega)\to L^{1}(\Omega)$, we have
-$\langle\mathcal{K}_{1}f,g\rangle_{L^{1},L^{\infty}}=\int\mathcal{K}_{1}f\cdot g\,dx=\int\int f\,d\rho_{x}g(x)\,dx=\int f\,d\nu_{g}=\left\langle f,\frac{d\nu_{g}}{dx}\right\rangle_{L^{1},L^{\infty}}.$
-This shows $\mathcal{K}_{1}^{*}g=\frac{d\nu_{g}}{dx}$. Finally, if $f,g\in L^{\infty}(\Omega)$, we have
+$\langle\mathcal{K}_{1}f,g\rangle_{L^{1},L^{\infty}}=\int_{\Omega}\mathcal{K}_{1}f\cdot g\,dx=\int_{\Omega}\int_{\Omega}f\,d\rho_{x}g(x)\,dx=\int_{\Omega}f\,d\nu_{g}=\left\langle f,\frac{d\nu_{g}}{dx}\right\rangle_{L^{1},L^{\infty}}.$
+This shows $\mathcal{K}_{1}^{*}g=\frac{d\nu_{g}}{dx}$. Finally, since $|\Omega|<\infty$ implies the inclusion $L^{\infty}(\Omega)\subset L^{2}(\Omega)$, we have for $f,g\in L^{\infty}(\Omega)$:
 $\langle\mathcal{K}_{2}^{*}f,g\rangle_{L^{1},L^{\infty}}=\langle\mathcal{K}_{2}^{*}f,g\rangle_{L^{2}}=\langle f,\mathcal{K}_{2}g\rangle_{L^{2}}=\langle f,\mathcal{K}_{1}g\rangle_{L^{1},L^{\infty}}=\langle\mathcal{K}_{1}^{*}f,g\rangle_{L^{1},L^{\infty}}.$
@@
 The inequality above implies
-$\left\|\frac{d\nu_{D^{\alpha}\varphi}}{dx}\right\|_{H^{-1}}=\|\mathcal{K}_{2}^{*}(D^{\alpha}\varphi)\|_{H^{-1}}$ $=\sup\left\{|\langle f,\mathcal{K}_{2}^{*}(D^{\alpha}\varphi)\rangle_{H^{1},H^{-1}}|:f\in H^{1},\,\|f\|_{H^{1}}=1\right\}\,$
-$\leq\,C_{1}\|\varphi\|_{L^{2}}.$
+$\left\|\frac{d\nu_{D^{\alpha}\varphi}}{dx}\right\|_{H^{-1}}=\|\mathcal{K}_{2}^{*}(D^{\alpha}\varphi)\|_{H^{-1}}$ $=\sup\left\{\left|\langle f,\mathcal{K}_{2}^{*}(D^{\alpha}\varphi)\rangle_{H^{1},H^{-1}}\right|:f\in H^{1},\,\|f\|_{H^{1}}=1\right\}\,$
+$\leq\,C_{1}\|\varphi\|_{L^{2}}$
 for every $|\alpha|=1$.
@@
 Hence, $|\langle\mathcal{K}f,D^{\alpha}\varphi\rangle|=|\langle f,K_{2}^{\ast}(D^{\alpha}\varphi)\rangle|\leq C_{\alpha}\|f\|_{H^{1}}\|\varphi\|_{L^{2}}$ holds for all $f\in H^{1}$ and all $\varphi\in C_{c}^{\infty}$.
-Therefore, for fixed $f\in H^{1}$, the functional $\varphi\mapsto\langle\mathcal{K}f,D^{\alpha}\varphi\rangle$ is bounded on $C_{c}^{\infty}$ and can thus be extended to a bounded functional $F_{f}$ on $L^{2}$.
+Therefore, for fixed $f\in H^{1}$, the functional $\varphi\mapsto\langle\mathcal{K}f,D^{\alpha}\varphi\rangle$ is bounded on $C_{c}^{\infty}$ and can be extended to a bounded functional $F_{f}$ on $L^{2}$.
 By Riesz’ representation theorem, $F_{f}\varphi=\langle Lf,\varphi\rangle$ for all $\varphi\in L^{2}$ with some $Lf\in L^{2}$.
@@
 Again, this gives the boundedness of $\mathcal{K}:H^{\ell}\to H^{\ell}$ by Proposition 2.4, and for all $\alpha\in\mathbb{N}^{n}$ with $|\alpha|\leq\ell$ we have $D^{\alpha}(\mathcal{K}f)\in L^{2}$.
-By the definition of the weak derivative (25)
+By the definition of the weak derivative (26), we have
 $\left|\left\langle\mathcal{K}f,D^{\alpha}\varphi\right\rangle\right|=\left|\left\langle D^{\alpha}(\mathcal{K}f),\varphi\right\rangle\right|$
@@
 In the next step, we use an interpolation argument to deduce a sufficient condition for fractional Sobolev space invariance.
-Step 3: $H^{s}$-invariance, $s\in(0,\ell]$. By [BrenScot2008, Theorem 14.2.3], for real-valued $s\in(0,\ell]$, the space $H^{s}$ coincides with the real interpolation space
-$H^{s}(\Omega)=[L^{2}(\Omega),H^{\ell}(\Omega)]_{\theta},\quad\theta=s/\ell\in(0,1)$
+Step 3: $H^{s}$-invariance, $s\in(0,\ell]$. By [8, Theorem 14.2.3], for real-valued $s\in(0,\ell]$, the space $H^{s}$ coincides with the real interpolation space
+$H^{s}(\Omega)=[L^{2}(\Omega),H^{\ell}(\Omega)]_{\theta},\quad\theta=s/\ell\in(0,1),$
 with equivalent norms.
-Under condition ($L^{1}$), we have the invariance of $L^{2}(\Omega)$ under $\mathcal{K}$.
+Under condition (4), we have the invariance of $L^{2}(\Omega)$ under $\mathcal{K}$.
 Let (24) hold for $\ell\in\mathbb{N}$.
 Then $H^{\ell}(\Omega)$ is also invariant under $\mathcal{K}$ by step 2.
-[Tart2007, Theorem 22.3] now implies that the Koopman operator
+[49, Theorem 22.3] now implies that the Koopman operator
 $\mathcal{K}:[L^{2}(\Omega),H^{\ell}(\Omega)]_{\theta}\to[L^{2}(\Omega),H^{\ell}(\Omega)]_{\theta}$
@@
 Let $k:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ be a continuous strictly positive-definite symmetric kernel and $\mathbb{H}$ the induced reproducing kernel Hilbert space on $\Omega$ with $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b}(\Omega),C_{b}(\Omega))$}.
-To verify especially the condition $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ in the assumption above, we refer to the sufficient conditions of the previous Section 3, where Theorem 3.2 guarantees the invariance of $\mathbb{H}$ since Sobolev spaces $\mathbb{H}=H^{s}$ are induced by Wendland kernels as RKHS [KohnPhil24, Theorem 4.1].
+To verify the condition $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ in Assumption 4.1, we refer to the sufficient conditions of the previous Section 3, where Theorem 3.2 guarantees the invariance of $\mathbb{H}$ since Sobolev spaces $\mathbb{H}=H^{s}$ are induced by Wendland kernels as RKHS [24, Theorem 4.1].
 We start by constructing the kernel-based approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$ of the stochastic Koopman operator from (19) more thoroughly.
-In an intermediate step, we derive a kernel EDMD approximant $\widehat{\mathcal{K}}$, significantly building upon the analogue from [KohnPhil24] in the deterministic setting.
+In an intermediate step, we derive a kernel EDMD approximant $\widehat{\mathcal{K}}$, significantly building upon the analogue from [24] in the deterministic setting.
 This allows to split the approximation error as follows:
@@
 The intermediary approximation $\widehat{\mathcal{K}}$ is employed in Section 4.1 to prove a bound on the deterministic projection error, see Theorem 4.2.
-Therein, we appeal to similar arguments as [KohnPhil24], where the invariance of a reproducing kernel Hilbert space under the Koopman operator is assumed as in our Assumption 4.1.
-Since $\widehat{\mathcal{K}}$ requires the exact knowledge of propagated observables under the stochastic Koopman operator, i.e., conditional expectation, Section 4.2 proposes the random, but computable approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$.
+Therein, we appeal to similar arguments as [24], where the invariance of a reproducing kernel Hilbert space under the Koopman operator is assumed as in our Assumption 4.1.
+Since $\widehat{\mathcal{K}}$ requires the exact knowledge of propagated observables under the stochastic Koopman operator, i.e., conditional expectations, Section 4.2 proposes the random, but computable approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$.
 The approximation of the “exact”, but not computable $\widehat{\mathcal{K}}$ introduces a probabilistic error that is dealt with in Theorem 4.6.
 4.1. Deterministic approximation step: interpolation
-The first (deterministic) approximation mimicks the approach of [KohnPhil24].
+The first (deterministic) approximation mimicks the approach of [24].
 However, this approximation will not be useful in practice as it involves expectation values and thus is not computable, in general.
@@
 $\widehat{\mathcal{K}}=S_{\mathcal{X}}\mathcal{K}S_{\mathcal{X}},$ (32)
-where $S_{\mathcal{X}}$ is the $\mathbb{H}$-orthogonal projection onto
-$V_{\mathcal{X}}=\operatorname{span}\{\Phi_{x_{i}}:i\in\{1,\dots,d\}\}\subset\mathbb{H}$
-with the canonical features $\Phi_{x_{i}}=k(x_{i},\cdot)\in\mathbb{H}$.
+where $S_{\mathcal{X}}$ is the $\mathbb{H}$-orthogonal projection onto $V_{\mathcal{X}}=\operatorname{span}\{\Phi_{x_{i}}:i\in\{1,\dots,d\}\}\subset\mathbb{H}$ with the canonical features $\Phi_{x_{i}}=\text{\sf{k}}(x_{i},\cdot)\in\mathbb{H}$.
 With the notation from Section 2.3, recall the explicit projection formula (11) given by $S_{\mathcal{X}}g=g_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}$ for some $g\in\mathbb{H}$.
-Hence, application of the abstract operator defined in (31) may explicitly be computed via
-$\widehat{\mathcal{K}}f=S_{\mathcal{X}}\mathcal{K}S_{\mathcal{X}}f=(\mathcal{K}S_{\mathcal{X}}f)_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}=\sum_{i=1}^{d}\big[\mathbf{K}_{\mathcal{X}}^{-1}\big(\underbrace{\mathbb{E}[(S_{\mathcal{X}}f)(Y)|X=\cdot]}_{\mathcal{K}S_{\mathcal{X}}f}\big)_{\mathcal{X}}\big]_{i}\Phi_{x_{i}}.$
-Using linearity of the expected value and the projection formula (10), we get, for each $j\in\{1,\dots,d\}$,
-$\mathbb{E}[S_{\mathcal{X}}f(Y)|X=x_{j}]=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbb{E}[\mathbf{k}_{\mathcal{X}}(Y)|X=x_{j}]=\mathbb{E}[\mathbf{k}_{\mathcal{X}}^{\top}(Y)|X=x_{j}]\mathbf{K}_{\mathcal{X}}^{-1}f_{\mathcal{X}}.$
-Hence, defining the coefficients
-$\alpha_{\widehat{\mathcal{K}}f}=\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}f_{\mathcal{X}},\qquad\mathrm{where}\qquad(\mathbf{K}_{\mathcal{X},\mathcal{Y}})_{i,j}=\mathbb{E}(k(Y,x_{i})|X=x_{j}),$
+The action of the abstract operator defined in (32) now may explicitly be computed via
+$\widehat{\mathcal{K}}f=S_{\mathcal{X}}\mathcal{K}S_{\mathcal{X}}f=(\mathcal{K}S_{\mathcal{X}}f)_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}=\sum_{i=1}^{d}\big[\mathbf{K}_{\mathcal{X}}^{-1}\big(\underbrace{\mathbb{E}[(S_{\mathcal{X}}f)(Y)\mid X=\cdot]}_{\mathcal{K}S_{\mathcal{X}}f}\big)_{\mathcal{X}}\big]_{i}\Phi_{x_{i}}.$
+Using linearity of the expected value and the projection formula (11), we get
+$\mathbb{E}[S_{\mathcal{X}}f(Y)\mid X=x_{j}]=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbb{E}[\mathbf{k}_{\mathcal{X}}(Y)\mid X=x_{j}]=\mathbb{E}[\mathbf{k}_{\mathcal{X}}^{\top}(Y)\mid X=x_{j}]\mathbf{K}_{\mathcal{X}}^{-1}f_{\mathcal{X}}$
+for each $j\in\{1,\dots,d\}$. Hence, defining the coefficients
+$\alpha_{\widehat{\mathcal{K}}f}=\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}f_{\mathcal{X}},\qquad\mathrm{where}\qquad(\mathbf{K}_{\mathcal{X},\mathcal{Y}})_{i,j}=\mathbb{E}[\text{\sf{k}}(Y,x_{i})\mid X=x_{j}],$
 yields the approximant
@@
 $\forall\,x_{0}\in\Omega\;\forall\,\Delta\in\mathcal{B}_{\Omega}(x_{0}):\quad\rho_{x}(\Delta)\to\rho_{x_{0}}(\Delta)\quad\text{ as $x\to x_{0}$},$ (34)
-where $\mathcal{B}_{\Omega}(x)=\{\Delta\in\mathcal{B}_{\Omega}:\rho_{x}(\partial\Delta)=0\}$ for $x\in\Omega$, where $\partial\Delta$ denotes the boundary of the Borel-set $\Delta$.
+where $\mathcal{B}_{\Omega}(x)=\{\Delta\in\mathcal{B}_{\Omega}:\rho_{x}(\partial\Delta)=0\}$ for $x\in\Omega$, with $\partial\Delta$ denoting the boundary of the Borel-set $\Delta$.
 Note that this holds in the deterministic case where $\rho_{x}=\delta_{F(x)}$ with continuous $F:\Omega\to\Omega$: Then, $\rho_{x_{0}}(\partial\Delta)=\delta_{F(x_{0})}(\partial\Delta)=0$ is equivalent to $F(x_{0})\notin\partial\Delta$, so $F(x_{0})$ lies either in the interior or the exterior of $\Delta$, and continuity of $F$ gives the convergence.
-By the Portmanteau theorem [Klen2008, Theorem 13.16], the condition (33) is equivalent to the fact that $\mathcal{K}f$ is continuous for each $f\in C_{b}(\Omega)$.
-Hence, if (33) is satisfied, the Koopman operator is a contraction from $C_{b}(\Omega)$ into itself as for $f\in C_{b}(\Omega)$,
+By the Portmanteau theorem [20, Theorem 13.16], the condition (34) is equivalent to the fact that $\mathcal{K}f$ is continuous for each $f\in C_{b}(\Omega)$.
+Hence, if (34) is satisfied, the Koopman operator is a contraction from $C_{b}(\Omega)$ into itself: for $f\in C_{b}(\Omega)$, we have
 $\|\mathcal{K}f\|_{\infty}=\sup_{x\in\Omega}\left|\int_{\Omega}f\,d\rho_{x}\right|\leq\sup_{x\in\Omega}\int_{\Omega}\|f\|_{\infty}\,d\rho_{x}=\|f\|_{\infty}.$
@@
 Theorem 4.2 (Deterministic error bound).
-Under Assumption 4.1, i.e. $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b},C_{b})$, the approximant $\widehat{\mathcal{K}}$ from (32) admits the error bound
+Under Assumption 4.1, i.e. $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b}(\Omega),C_{b}(\Omega))$, the approximant $\widehat{\mathcal{K}}$ from (33) admits the error bound
 $\big\|\mathcal{K}-\widehat{\mathcal{K}}\big\|_{\mathbb{H}\to C_{b}}\,\leq\,(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}.$ (35)
 Proof.
-The proof is a particular case of the proof of [KohnPhil24, Theorem 3.4] and we repeat it here for the sake of clarity.
+The proof is a particular case of the proof of [24, Theorem 3.4] and we repeat it here for the sake of clarity.
 As $\mathcal{K}-\widehat{\mathcal{K}}=\mathcal{K}-S_{\mathcal{X}}\mathcal{K}S_{\mathcal{X}}=\mathcal{K}(\mathcal{I}-S_{\mathcal{X}})+(\mathcal{I}-S_{\mathcal{X}})\mathcal{K}S_{\mathcal{X}}$, we get
 $\big\|\mathcal{K}-\widehat{\mathcal{K}}\big\|_{\mathbb{H}\to C_{b}}$ $\leq\|\mathcal{K}(\mathcal{I}-S_{\mathcal{X}})\|_{\mathbb{H}\to C_{b}}+\|(\mathcal{I}-S_{\mathcal{X}})\mathcal{K}S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}$
-$\leq\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}(\|\mathcal{K}\|_{C_{b}\to C_{b}}+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\|S_{\mathcal{X}}\|_{\mathbb{H}\to\mathbb{H}}).$
-Under the assumption $\mathcal{K}\in L(C_{b},C_{b})$, $\mathcal{K}$ is a contraction, and for constant $f\equiv c\in\Omega$
+$\leq\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}\,\left(\|\mathcal{K}\|_{C_{b}\to C_{b}}+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\|S_{\mathcal{X}}\|_{\mathbb{H}\to\mathbb{H}}\right).$
+Under the assumption $\mathcal{K}\in L(C_{b},C_{b})$, $\mathcal{K}$ is a contraction. Any constant $f\equiv c\in\Omega$ satisfies
 $\mathcal{K}f(x)=\mathbb{E}[f(Y)\mid X=x]=\mathbb{E}[c\mid X=x]=c=f(x)\quad\forall x\in\Omega,$
@@
 The projection error in (35) attains a particularly intuitive form when considering Sobolev spaces as RKHS.
-As stated before, Wendland and Matérn kernels both generate such Sobolev spaces, allowing for the application of results of approximation theory ([Wend04, Theorem 11.13]) to further specify Theorem 4.2 in these cases.
+As stated before, Wendland and Matérn kernels both generate such Sobolev spaces, allowing for the application of results of approximation theory ([50, Theorem 11.13]) to further specify Theorem 4.2 in these cases.
 Under the appropriate sufficient condition from Section 3, the order of convergence can be quantified in terms of the smoothness of the underlying kernels.
@@
 Let $l\in\mathbb{N}$ and define $\sigma_{n,l}=\frac{n+1}{2}+l$.
-Consider the Wendland kernel $k_{n,l}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ that generates the Sobolev space $H^{\sigma_{n,l}}(\Omega)$ as reproducing kernel Hilbert space with equivalence between the native norm $\|\cdot\|_{\mathbb{H}}$ and the classical Sobolev norm $\|\cdot\|_{H^{\sigma_{n,l}}}$.
+Consider the Wendland kernel $\text{\sf{k}}_{n,l}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ that generates the Sobolev space $H^{\sigma_{n,l}}(\Omega)$ as reproducing kernel Hilbert space $\mathbb{H}$ with equivalence between the native norm $\|\cdot\|_{\mathbb{H}}$ and the Sobolev norm $\|\cdot\|_{H^{\sigma_{n,l}}}$.
 Assume that $\mathcal{K}\in L(C_{b},C_{b})$ and that (24) is satisfied for some $\ell\in\mathbb{N}$ with $\ell>\sigma_{n,l}$.
@@
 Under (24), the Sobolev space $H^{\sigma_{n,l}}$ is invariant under the Koopman operator $\mathcal{K}$.
-By [Wend04, Theorem 11.17] or the restatement [KohnPhil24, Theorem 5.1], we may bound the projection error by
+By [50, Theorem 11.17] or the restatement [24, Theorem 5.1], we may bound the projection error by
 $\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}\leq Ch_{\mathcal{X}}^{l+1/2}.$
@@
 For a large data set $\mathcal{X}$, the kernel matrix $K_{\mathcal{X}}$ is typically badly conditioned, leading to numerical instabilities.
-To alleviate this, one might regularize the kernel-based interpolation problem as in [BoldPhil24] to get $R_{\mathcal{X}}^{\lambda}f:={\arg\min}_{{g\in V_{\mathcal{X}}}}\sum_{i=1}^{d}\left|f(x_{i})-g(x_{i})\right|^{2}+\lambda\|g\|_{\mathbb{H}}^{2}$ with regularization parameter $\lambda\geq 0$ and solution operator $R_{\mathcal{X}}^{\lambda}$.
-The linear solution operator satisfies $R_{\mathcal{X}}^{\lambda}f=f_{\mathcal{X}}^{\top}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}$} for $f\in\mathbb{H}$, so for $\lambda=0$ we recover $S_{\mathcal{X}}$, but for $\lambda>0$ $R_{\mathcal{X}}^{\lambda}:\mathbb{H}\to\mathbb{H}$ is no longer a projection.
-However, as is shown in [BoldPhil24, Appendix A], in analogy to (31) we can still define the approximant
+To alleviate this, one might regularize the kernel-based interpolation problem as in [7] to get $R_{\mathcal{X}}^{\lambda}f:={\arg\min}_{{g\in V_{\mathcal{X}}}}\sum_{i=1}^{d}\left|f(x_{i})-g(x_{i})\right|^{2}+\lambda\|g\|_{\mathbb{H}}^{2}$ with regularization parameter $\lambda\geq 0$ and solution operator $R_{\mathcal{X}}^{\lambda}$.
+The linear solution operator satisfies $R_{\mathcal{X}}^{\lambda}f=f_{\mathcal{X}}^{\top}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}$} for $f\in\mathbb{H}$, so for $\lambda=0$ we recover $S_{\mathcal{X}}$, while for $\lambda>0$, $R_{\mathcal{X}}^{\lambda}:\mathbb{H}\to\mathbb{H}$ is no longer a projection.
+However, as is shown in [7, Appendix A], in analogy to (32) we can still define the approximant
 $\widehat{\mathcal{K}}_{\lambda}:=R_{\mathcal{X}}^{\lambda}\,\mathcal{K}\,R_{\mathcal{X}}^{\lambda}$ (37)
@@
 $\bigl\|\mathcal{K}-\widehat{\mathcal{K}}_{\lambda}\bigr\|_{\mathbb{H}\to C_{b}}\leq\left(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\right)\,\|\mathcal{I}-R_{\mathcal{X}}^{\lambda}\|_{\mathbb{H}\to C_{b}}.$ (39)
-Specifically, for the RKHS $\mathbb{H}=H^{\sigma_{n,l}}$ generated by the Wendland kernel $k_{n,l}$, where $\sigma_{n,l}=\frac{n+1}{2}+l$, the deterministic error admits the bound
+Specifically, for the RKHS $\mathbb{H}=H^{\sigma_{n,l}}(\Omega)$ generated by the Wendland kernel $\text{\sf{k}}_{n,l}$, where $\sigma_{n,l}=\frac{n+1}{2}+l$, the deterministic error admits the bound
 $\|(\mathcal{K}-\widehat{\mathcal{K}}_{\lambda})f\|_{\infty}\leq C\left(h_{\mathcal{X}}^{l+1/2}+\lambda^{1/2}\right)\|f\|_{\mathbb{H}}.$ (40)
 Proof.
-The proof of (38) follows the lines of the proof of Theorem 4.2 and uses the fact that $\bigl\|R_{\mathcal{X}}^{\lambda}\bigr\|_{\mathbb{H}\to\mathbb{H}}\leq 1$ as shown in [BoldPhil24, Proposition B.1 (iii)].
-The special case covered in (39) follows from [BoldPhil24, Theorem 2.4].
+The proof of (39) follows the lines of the proof of Theorem 4.2 and uses the fact that $\bigl\|R_{\mathcal{X}}^{\lambda}\bigr\|_{\mathbb{H}\to\mathbb{H}}\leq 1$ as shown in [7, Proposition B.1 (iii)].
+The special case covered in (40) follows from [7, Theorem 2.4].
 ∎
@@
 In practice, we cannot form $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$, as we do not have access to the expected values.
-For this reason, we introduce a random matrix ${\mathbf{K}}^{\mathrm{MC}}_{\mathrm{prop}}$, whose entries are Monte Carlo approximations of the expected values in $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$.
+For this reason, we introduce a random matrix ${\mathbf{K}}^{\mathrm{MC}}_{\mathcal{X},\mathcal{Y}}$, whose entries are Monte Carlo approximations of the expected values in $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$.
 We denote the $j$-th column of $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$ by $(\mathbf{K}_{\mathcal{X},\mathcal{Y}})_{j}$.
 The $j$-th column of $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ is then defined by
-$\big(\mathbf{K}^{\mathrm{MC}}_{\mathcal{X},\mathcal{Y}}\big)_{j}:=\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}\mathbf{k}_{\mathcal{X}}(y_{l,j})\approx\mathbb{E}[\mathbf{k}_{\mathcal{X}}(Y)|X=x_{j}],$ (40)
+$\big(\mathbf{K}^{\mathrm{MC}}_{\mathcal{X},\mathcal{Y}}\big)_{j}:=\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}\mathbf{k}_{\mathcal{X}}(y_{l,j})\approx\mathbb{E}[\mathbf{k}_{\mathcal{X}}(Y)\mid X=x_{j}],$ (41)
 where $y_{l,j}$, $l\in\{1,\dots,m_{j}\}$, are drawn i.i.d. from $\rho(x_{j},\cdot)$, $j\in\{1,\dots,d\}$.
-This leads to a computable approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$ which is analogously defined as $\widehat{\mathcal{K}}$ in (32) with the only difference that we replace $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$ with $\mathbf{K}^{\mathrm{MC}}_{\mathcal{X},\mathcal{Y}}$ as defined in (40), that is
+This leads to a computable approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$, which is defined analogously to $\widehat{\mathcal{K}}$ in (33), with the only difference that we replace $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$ with $\mathbf{K}^{\mathrm{MC}}_{\mathcal{X},\mathcal{Y}}$ as defined in (41).
+Thus, we set
 $\widehat{\mathcal{K}}^{\mathrm{MC}}f=f_{\mathcal{X}}^{\top}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{k}_{\mathcal{X}}.$ (42)
-Decomposing the full error by inserting the auxiliary approximation $\widehat{\mathcal{K}}$, i.e,
+Decomposing the full error by inserting the auxiliary approximation $\widehat{\mathcal{K}}$, i.e.,
 […]
@@
 Due to the random sampling employed in $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$, this error is probabilistic.
-In the remainder of this part, we derive a probabilistic bound on the error
-$(\widehat{\mathcal{K}}-\widehat{\mathcal{K}}^{\mathrm{MC}})f=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}(\mathbf{K}_{\mathcal{X},\mathcal{Y}}-\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}})\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}.$
+In the remainder of this part, we derive a probabilistic uniform bound on the error
+$(\widehat{\mathcal{K}}-\widehat{\mathcal{K}}^{\mathrm{MC}})f=f_{\mathcal{X}}^{\top}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\left(\mathbf{K}_{\mathcal{X},\mathcal{Y}}-\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\right)\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{k}_{\mathcal{X}}.$
 For this, let us define the vectors
@@
 Theorem 4.6 (Monte Carlo error bound).
-Let $\varepsilon>0$ and choose $m_{j}=m\in\mathbb{N}$ i.i.d.
-samples from every $\rho_{x_{j}}$, $j=1,\dots,d$.
+Choose $m_{j}=m\in\mathbb{N}$ i.i.d.
+samples from every $\rho_{x_{j}}$, $j=1,\dots,d$, and let $\varepsilon>0$ such that $m\geq\frac{2\|\text{\sf{k}}\|_{\infty}}{\varepsilon^{2}}\log(2d)$.
 Then, for all $f\in\mathbb{H}$, the Koopman approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$ from (42) satisfies
-$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|k\|_{\infty}^{1/2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}\|f\|_{\mathbb{H}}$ (42)
-with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right)$.
+$\left\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\right\|_{\infty}\,\leq\,\varepsilon\|\text{\sf{k}}\|_{\infty}^{1/2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}\cdot\|f\|_{\mathbb{H}}$ (43)
+with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right)$, where $\mathbbm{1}=\{v\in\mathbb{R}^{d}:|v_{i}|=1\}$.
 Proof.
 For all $x\in\Omega$, we may interpret this as a perturbed interpolation problem, where the perturbation of the function values stems from approximation of the expected value:
-$|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f(x)|=|(g-g^{\mathrm{MC}})^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)|\leq\|g-g^{\mathrm{MC}}\|_{\infty}\,\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}.$
-As shown in the proof of [BoldPhil24, Theorem 4.2], the deterministic second factor is bounded by $\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}\leq\|k\|_{\infty}^{1/2}\cdot\big(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big)^{1/2}$, where $\mathbbm{1}=\{v\in\mathbb{R}^{d}:|v_{i}|=1\}$.
+$\left|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f(x)\right|=\left|(g-g^{\mathrm{MC}})^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\right|\leq\left\|g-g^{\mathrm{MC}}\right\|_{\infty}\,\left\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\right\|_{1}.$
+As shown in the proof of [7, Theorem 4.2], the deterministic second factor is bounded by $\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}\leq\|\text{\sf{k}}\|_{\infty}^{1/2}\cdot\big(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big)^{1/2}$.
 Moreover,
-$\|g-g^{\mathrm{MC}}\|_{\infty}=\max_{x_{j}\in\mathcal{X}}\left|\mathbb{E}[(S_{\mathcal{X}}f)(Y)|X=x_{j}]-\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}(S_{\mathcal{X}}f)(y_{l,j})\right|.$
-Setting $\mathbb{E}_{j}:=g_{j}=\mathbb{E}[(S_{\mathcal{X}}f)(Y)|X=x_{j}]$, we obtain $\mathbb{E}[(S_{\mathcal{X}}f)(y_{l,j})]=\int(S_{\mathcal{X}}f)(y)\,\rho(x_{j},dy)=(\mathcal{K}S_{\mathcal{X}}f)(x_{j})=\mathbb{E}_{j}$.
-Moreover,
-$|(S_{\mathcal{X}}f)(y)|=|\langle S_{\mathcal{X}}f,\Phi_{y}\rangle_{\mathbb{H}}|\,\leq\,\|S_{\mathcal{X}}f\|_{H}\|\Phi_{y}\|_{\mathbb{H}}\,\leq\,\|k\|_{\infty}^{1/2}\|f\|_{\mathbb{H}}$
-for all $y\in\Omega$, where the last inequality follows from $\|S_{\mathcal{X}}\|_{H\to H}=1$ since $S_{\mathcal{X}}$ is a projection.
+$\|g-g^{\mathrm{MC}}\|_{\infty}=\max_{x_{j}\in\mathcal{X}}\left|\mathbb{E}[(S_{\mathcal{X}}f)(Y)\mid X=x_{j}]-\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}(S_{\mathcal{X}}f)(y_{l,j})\right|.$
+Setting $\mathbb{E}_{j}:=g_{j}=\mathbb{E}[(S_{\mathcal{X}}f)(Y)\mid X=x_{j}]$, we obtain $\mathbb{E}[(S_{\mathcal{X}}f)(y_{l,j})]=\int(S_{\mathcal{X}}f)(y)\,\rho(x_{j},dy)=(\mathcal{K}S_{\mathcal{X}}f)(x_{j})=\mathbb{E}_{j}$.
+We also have
+$|(S_{\mathcal{X}}f)(y)|=|\langle S_{\mathcal{X}}f,\Phi_{y}\rangle_{\mathbb{H}}|\,\leq\,\|S_{\mathcal{X}}f\|_{H}\|\Phi_{y}\|_{\mathbb{H}}\,\leq\,\|\text{\sf{k}}\|_{\infty}^{1/2}\|f\|_{\mathbb{H}}$
+for all $y\in\Omega$, where the last inequality follows from $\|S_{\mathcal{X}}\|_{H\to H}=1$ since $S_{\mathcal{X}}$ is an orthogonal projection.
 Hence, the random variables $(S_{\mathcal{X}}f)(y_{l,j})$ are bounded and i.i.d., so that we may apply Hoeffding’s inequality:
 $\mathbb{P}\left(\left|\mathbb{E}_{j}-\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}(S_{\mathcal{X}}f)(y_{l,j})\right|\,\geq\,\varepsilon\|f\|_{H}\right)$ $=\mathbb{P}\left(\left|\sum_{l=1}^{m_{j}}\big[(S_{\mathcal{X}}f)(y_{l,j})-\mathbb{E}_{j}\big]\right|\,\geq\,\varepsilon m_{j}\|f\|_{H}\right)$
-$\leq 2\exp\left(-\frac{2m_{j}^{2}\varepsilon^{2}\|f\|_{H}^{2}}{4m_{j}\|k\|_{\infty}\|f\|_{H}^{2}}\right)=2\exp\left(-\frac{m_{j}\varepsilon^{2}}{2\|k\|_{\infty}}\right).$
+$\leq 2\exp\left(-\frac{2m_{j}^{2}\varepsilon^{2}\|f\|_{H}^{2}}{4m_{j}\|\text{\sf{k}}\|_{\infty}\|f\|_{H}^{2}}\right)=2\exp\left(-\frac{m_{j}\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right).$
 Plugging this entry-wise result into a union bound, we readily obtain:
@@
 $\leq\sum_{j=1}^{d}\mathbb{P}\left(\left|\mathbb{E}_{j}-\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}(S_{\mathcal{X}}f)(y_{l,j})\right|\,\geq\,\varepsilon\|f\|_{H}\right)$
-$\leq 2\sum_{j=1}^{d}\exp\left(-\frac{m_{j}\varepsilon^{2}}{2\|k\|_{\infty}}\right).$
+$\leq 2\sum_{j=1}^{d}\exp\left(-\frac{m_{j}\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right).$
 Hence, if $m_{j}=m$ for all $j\in\{1,\dots,d\}$, then
-$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{\infty}\geq\varepsilon\|f\|_{H})$ $\leq 2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right).$
-In total, for every $f\in H$, with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right)$, we obtain
-$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|k\|_{\infty}^{1/2}\cdot\big(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big)^{1/2}\|f\|_{H},$
-i.e., (42). ∎
-Note that in the bound (42) we have $\|k\|_{\infty}<\infty$ by Assumption 4.1, since the kernel $k$ is continuous on $\mathbb{R}^{n}\times\mathbb{R}^{n}$ and hence bounded on the bounded open set $\Omega\times\Omega$.
+$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{\infty}\geq\varepsilon\|f\|_{H}\big)$ $\leq 2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right).$
+In total, for every $f\in H$, we obtain the upper bound
+$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|\text{\sf{k}}\|_{\infty}^{1/2}\cdot\Big(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\Big)^{1/2}\|f\|_{H},$
+i.e., (43), with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right)$, which is nonnegative due to the imposed inequality $m\geq\frac{2\|\text{\sf{k}}\|_{\infty}}{\varepsilon^{2}}\log(2d)$.
+∎
+Note that in the bound (43) we have $\|\text{\sf{k}}\|_{\infty}<\infty$ by Assumption 4.1, since the kernel k is continuous on $\mathbb{R}^{n}\times\mathbb{R}^{n}$ and hence bounded on the bounded open set $\Omega\times\Omega$.
 Remark 4.7 (Alternative probabilistic bound).
@@
 Applying the Hilbert-space Hoeffding inequality circumvents the union bound used in the proof of Theorem 4.6 to directly get
-$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{2}\geq\varepsilon\|f\|_{H})\leq 2\exp\left(-\frac{m\varepsilon^{2}}{8d\|k\|_{\infty}}\right).$ (43)
+$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{2}\geq\varepsilon\|f\|_{H})\leq 2\exp\left(-\frac{m\varepsilon^{2}}{8d\|\text{\sf{k}}\|_{\infty}}\right).$ (44)
 By the $\ell^{1}$-$\ell^{2}$-inequality, the deterministic second term is smaller than the one in Theorem 4.6, i.e., $\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{2}\leq\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}$.
 We refer the reader to Appendix A for detailed calculations.
-In total, with probability at least $1-2\exp\left(-\frac{m\varepsilon^{2}}{8d\|k\|_{\infty}}\right)$, the bound
+In total, with probability at least $1-2\exp\left(-\frac{m\varepsilon^{2}}{8d\|\text{\sf{k}}\|_{\infty}}\right)$, the bound
 $\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{2}\,\|f\|_{H}.$
@@
 The bound
-$\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}\leq\|k\|_{\infty}^{1/2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$
+$\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}\leq\|\text{\sf{k}}\|_{\infty}^{1/2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$
 from the proof of Theorem 4.6 warrants further discussion.
@@
 $\leq d\cdot\lambda_{\mathrm{max}}(\mathbf{K}_{\mathcal{X}}^{-1})\,\mathbf{k}_{\mathcal{X}}(x)^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{k}_{\mathcal{X}}(x).$
-From the positive definiteness of the kernel $k$, we get the positive semidefiniteness of the augmented kernel matrix $\mathbf{K}_{\mathcal{X}\cup\{x\}}=\begin{pmatrix}\mathbf{K}_{\mathcal{X}}&\mathbf{k}_{\mathcal{X}}\\ \mathbf{k}_{\mathcal{X}}^{\top}&k(x,x)\end{pmatrix}$ with positive semidefinite (thus, non-negative) Schur complement $k(x,x)-\mathbf{k}_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}\geq 0$.
+From the positive definiteness of the kernel k, we get the positive semidefiniteness of the augmented kernel matrix $\mathbf{K}_{\mathcal{X}\cup\{x\}}=\left(\begin{smallmatrix}\mathbf{K}_{\mathcal{X}}&\mathbf{k}_{\mathcal{X}}\\ \mathbf{k}_{\mathcal{X}}^{\top}&\text{\sf{k}}(x,x)\end{smallmatrix}\right)$ with positive semidefinite (thus, non-negative) Schur complement $\text{\sf{k}}(x,x)-\mathbf{k}_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}\geq 0$.
 With $\lambda_{\mathrm{max}}(\mathbf{K}_{\mathcal{X}}^{-1})=\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})^{-1}$, where $\lambda_{\max}$ and $\lambda_{\min}$ are the largest and smallest eigenvalue of the respective matrices, we get
-$\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}^{2}\leq k(x,x)\cdot\frac{d}{\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})}.$
+$\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}^{2}\leq\text{\sf{k}}(x,x)\cdot\frac{d}{\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})}.$
 The term in the denominator usually decays rapidly as the condition of the kernel matrix gets worse.
-If our kernel is given by a radial basis function (RBF) $B:\mathbb{R}^{n}\to\mathbb{R}$, $k(x,x)=B(0)$ is constant, so
+If our kernel is given by a radial basis function (RBF) $B:\mathbb{R}^{n}\to\mathbb{R}$, $\text{\sf{k}}(x,x)=B(0)$ is constant, so
 $\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}^{2}\leq B(0)\frac{d}{\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})}.$
-For example, for the Wendland kernel $k_{n,l}$ with RBF $B_{n,l}$, we have $B_{n,l}(0)=\frac{1}{2^{l}}\binom{\lfloor n/2\rfloor+2l+1}{l}$, while for Matérn kernels $k_{\nu}$ with RBF $\mathcal{M}_{\nu}$, we have $\mathcal{M}_{\nu}(0)=1$.
+For example, for the Wendland kernel $\text{\sf{k}}_{n,l}$ with RBF $B_{n,l}$, we have $B_{n,l}(0)=\frac{1}{2^{l}}\binom{\lfloor n/2\rfloor+2l+1}{l}$, while for Matérn kernels $\text{\sf{k}}_{\nu}$ with RBF $\mathcal{M}_{\nu}$, we have $\mathcal{M}_{\nu}(0)=1$.
 To balance the decay of the eigenvalue $\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})$ to zero, the computable Monte Carlo approximation (42) also lends itself to a Tikhonov regularization in style of (37).
@@
 Corollary 4.9 (Regularized Monte Carlo error bound).
-Let the assumptions of of Theorem 4.6 hold. For $f\in\mathbb{H}$, the regularized approximant $\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}$ from (44) satisfies
-$\|(\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}-\widehat{\mathcal{K}}_{\lambda})f\|_{\infty}\,\leq\,\varepsilon\cdot\|k\|_{\infty}^{1/2}\cdot\sqrt{\frac{d}{\lambda}}\|f\|_{H}$ (45)
-with probability at least $1-2de^{-m\varepsilon^{2}/(2\|k\|_{\infty})}$.
+Let the assumptions of Theorem 4.6 hold. For $f\in\mathbb{H}$, the regularized approximant $\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}$ from (45) satisfies
+$\left\|(\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}-\widehat{\mathcal{K}}_{\lambda})f\right\|_{\infty}\,\leq\,\varepsilon\cdot\|\text{\sf{k}}\|_{\infty}^{1/2}\cdot\sqrt{\frac{d}{\lambda}}\,\|f\|_{\mathbb{H}}$ (46)
+with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right)$.
 Proof.
@@
 Following Remark 4.8, this yields
-$\|(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}^{2}\leq k(x,x)\frac{d}{\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}}+\lambda I)}\leq k(x,x)\frac{d}{\lambda}.$
+$\|(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}^{2}\leq\text{\sf{k}}(x,x)\frac{d}{\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}}+\lambda I)}\leq\text{\sf{k}}(x,x)\frac{d}{\lambda}.$
 For the other term, we have
-$\|g_{\lambda}-g_{\lambda}^{\mathrm{MC}}\|_{\infty}=\max_{x_{j}\in\mathcal{X}}\left|\mathbb{E}[(R_{\mathcal{X}}^{\lambda}f)(Y)|X=x_{j}]-\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}(R_{\mathcal{X}}^{\lambda}f)(y_{l,j})\right|$
+$\|g_{\lambda}-g_{\lambda}^{\mathrm{MC}}\|_{\infty}=\max_{x_{j}\in\mathcal{X}}\left|\mathbb{E}[(R_{\mathcal{X}}^{\lambda}f)(Y)\mid X=x_{j}]-\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}(R_{\mathcal{X}}^{\lambda}f)(y_{l,j})\right|$
 with the operator $R_{\mathcal{X}}^{\lambda}$ from (38).
-By [BoldPhil24, Proposition B.1], that “projection-like” operator satisfies $\|R_{\mathcal{X}}^{\lambda}\|_{\mathbb{H}\to\mathbb{H}}\leq\|S_{\mathcal{X}}\|_{\mathbb{H}\to\mathbb{H}}=1$.
+By [7, Proposition B.1], that “projection-like” operator satisfies $\|R_{\mathcal{X}}^{\lambda}\|_{\mathbb{H}\to\mathbb{H}}\leq\|S_{\mathcal{X}}\|_{\mathbb{H}\to\mathbb{H}}=1$.
 Thus, the arguments from the proof of Theorem 4.6 carry over, and for every $f\in\mathbb{H}$, we get
-$\mathbb{P}\left(\|g_{\lambda}-g_{\lambda}^{\mathrm{MC}}\|_{\infty}\geq\varepsilon\|f\|_{\mathbb{H}}\right)\leq 2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right).$
+$\mathbb{P}\left(\|g_{\lambda}-g_{\lambda}^{\mathrm{MC}}\|_{\infty}\geq\varepsilon\|f\|_{\mathbb{H}}\right)\leq 2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right).$
 In total, this yields (46). ∎
 The appeal of Corollary 4.9 lies in the more explicit bound in terms of the number $d$ of data points in $\mathcal{X}$.
-The regularization parameter $\lambda$ puts a “floor” on the conditioning of the kernel matrix $\mathbf{K}_{\mathcal{X}}$, making the tradeoff between deterministic regularization bias and kernel matrix conditioning evident.
-In the final subsection, we combine the deterministic and Monte Carlo error bounds to get a bound on the full $L^{\infty}$-error of the approximations $\widehat{\mathcal{K}}^{\mathrm{MC}}$ and $\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}$.
+The regularization parameter $\lambda$ puts a “floor” on the smallest eigenvalue of the kernel matrix $\mathbf{K}_{\mathcal{X}}$, making the tradeoff between deterministic regularization bias and kernel matrix conditioning evident.
+In the final subsection, we combine the deterministic and Monte Carlo error bounds to get a bound on the full $L^{\infty}$-error of the approximations $\widehat{\mathcal{K}}^{\mathrm{MC}}$ and $\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}$, proving the remaining main results from Section 2.
 4.3. Proof of Theorems 2.6 and 2.8
@@
 Theorem 0 ($L^{\infty}$-error bound, restated).
-Let $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b},C_{b})$.
+Let $\mathbb{H}$ be an RKHS of functions over $\Omega$, and assume $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b}(\Omega),C_{b}(\Omega))$.
 Furthermore, for $\mathcal{X}=\{x_{1},\ldots,x_{d}\}$, let $m_{j}=m\in\mathbb{N}$ i.i.d.
-samples from every $\rho_{x_{j}}$, $j\in\{1,\dots,d\}$.
-Then, for every $\varepsilon>0$ and for all $f\in\mathbb{H}$, $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
-$\|(\mathcal{K}-\widehat{\mathcal{K}}^{\mathrm{MC}})f\|_{\infty}\leq\big[1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\big]\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}+\varepsilon\cdot\|k\|_{\infty}^{1/2}\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$
-with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right)$.
+samples from every $\rho_{x_{j}}$, $j\in\{1,\dots,d\}$, be given.
+Then, for every $\varepsilon>0$ satisfying $m\geq\frac{2\|\text{\sf{k}}\|_{\infty}}{\varepsilon^{2}}\log(2d)$ and for all $f\in\mathbb{H}$, $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
+$\big\|\big(\mathcal{K}-\widehat{\mathcal{K}}^{\mathrm{MC}}\big)f\big\|_{\infty}\leq\big[1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\big]\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}+\varepsilon\cdot\|\text{\sf{k}}\|_{\infty}^{1/2}\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$
+with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right)$, where $\mathbbm{1}=\{v\in\mathbb{R}^{d}:|v_{i}|=1\}$.
 Proof.
@@
 By Theorem 4.6, the second summand admits the bound
-$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|k\|_{\infty}^{1/2}\cdot\big(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big)^{1/2}\|f\|_{H}$
-with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right)$. With $\|f\|_{\mathbb{H}}\leq 1$, those two inequalities yield the claim. ∎
+$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|\text{\sf{k}}\|_{\infty}^{1/2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}\,\|f\|_{H}$
+with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right)$. With $\|f\|_{\mathbb{H}}\leq 1$, those two inequalities yield the claim. ∎
 The projection error $\mathcal{I}-S_{\mathcal{X}}$ and the expression involving the kernel matrix $\mathbf{K}_{\mathcal{X}}^{-1}$ warrant a closer inspection in certain special cases.
-For Sobolev spaces and Wendland RKHS, the projection error in Theorem 2.8 is bounded in terms of the fill distance $h_{\mathcal{X}}$.
+For Sobolev spaces as Wendland RKHS, the projection error in Theorem 2.8 is bounded in terms of the fill distance $h_{\mathcal{X}}$.
 Theorem 0 ($L^{\infty}$-error bound for Wendland native spaces, restated).
-Let $l\in\mathbb{N}$, define $\sigma_{n,l}=\frac{n+1}{2}+l$ and denote by $\mathbb{H}$ the Sobolev space $H^{\sigma_{n,l}}$ induced by the Wendland kernel $k_{n,l}$ as RKHS.
-Assume that $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b},C_{b})$.
-Then, for every $\varepsilon>0$ and all $f\in\mathbb{H}$ with $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
+Let $l\in\mathbb{N}$, define $\sigma_{n,l}=\frac{n+1}{2}+l$ and denote by $\mathbb{H}$ the Sobolev space $H^{\sigma_{n,l}}(\Omega)$ induced by the Wendland kernel $\text{\sf{k}}_{n,l}$ as RKHS.
+Assume that $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b}(\Omega),C_{b}(\Omega))$.
+Furthermore, for $\mathcal{X}=\{x_{1},\ldots,x_{d}\}\subset\Omega$, let $m_{j}=m\in\mathbb{N}$ i.i.d.
+samples from every $\rho_{x_{j}}$, $j\in\{1,\dots,d\}$, be given.
+Then, for every $\varepsilon>0$ satisfying $m\geq\frac{2C_{2}^{2}}{\varepsilon^{2}}\log(2d)$ and all $f\in\mathbb{H}$ with $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
 $\|(\mathcal{K}-\mathcal{K}^{\mathrm{MC}})f\|_{\infty}\leq$ $C_{1}\,(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot h_{\mathcal{X}}^{l+1/2}+\varepsilon\cdot C_{2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$
-with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2C_{2}^{2}}\right)$, with constants $C_{1},C_{2}>0$ only depending on the kernel $k_{n,l}$.
+with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2C_{2}^{2}}\right)$, where $C_{1},C_{2}>0$ are constants only depending on the kernel $\text{\sf{k}}_{n,l}$, and $\mathbbm{1}=\{v\in\mathbb{R}^{d}:|v_{i}|=1\}$.
 Proof.
@@
 for some $C_{1}>0$. The second summand is bounded by Theorem 4.6, which gives
-$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|k_{n,l}\|_{\infty}^{1/2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}\|f\|_{H}$
-with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k_{n,l}\|_{\infty}}\right)$.
-Setting $C_{2}=\|k_{n,l}\|_{\infty}^{1/2}$ and taking both inequalities for $\|f\|_{\mathbb{H}}\leq 1$ yields the claim.
+$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|\text{\sf{k}}_{n,l}\|_{\infty}^{1/2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}\|f\|_{H}$
+with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}_{n,l}\|_{\infty}}\right)$.
+Setting $C_{2}=\|\text{\sf{k}}_{n,l}\|_{\infty}^{1/2}$ and taking both inequalities for $\|f\|_{\mathbb{H}}\leq 1$ yields the claim.
 ∎
-Note that Wendland kernels are defined via $k_{n,l}(x,y)=B_{n,l}(\|x-y\|_{2})$ with the radial basis function $B_{n,l}:[0,\infty)\to\mathbb{R}$, see Appendix B.
-It can be shown that $\|k_{n,l}\|_{\infty}=\sup_{x\in\mathbb{R}^{n}}|k_{n,l}(x,x)|=B_{n,l}(0)=p_{n,l}(0)$ (see, e.g., [PhilSchaWort2025]), where $p_{n,l}$ is the recursively defined polynomial underlying the Wendland kernel [Wend04].
+Note that Wendland kernels are defined via $\text{\sf{k}}_{n,l}(x,y)=B_{n,l}(\|x-y\|_{2})$ with the radial basis function $B_{n,l}:[0,\infty)\to\mathbb{R}$, see Appendix B.
+It can be shown that $\|\text{\sf{k}}_{n,l}\|_{\infty}=\sup_{x\in\mathbb{R}^{n}}|\text{\sf{k}}_{n,l}(x,x)|=B_{n,l}(0)=p_{n,l}(0)$ (see, e.g., [41]), where $p_{n,l}$ is the recursively defined polynomial underlying the Wendland kernel [50].
 Thus, we get the explicitly computable expression $C_{2}=\sqrt{p_{n,l}(0)}$ for the constant from (22) that depends on dimension $n$ and smoothness $l$.
@@
 Thus, to control the probabilistic error, one must choose a small enough $\varepsilon>0$ that compensates for the conditioning.
-Lastly, we want our error bound to hold with probability close to one, so the amount of Monte Carlo samples $m$ for the entries of $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ has to guarantee that $2d\exp(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}})$ is close to zero.
+Lastly, we want our error bound to hold with probability close to one, so the amount of Monte Carlo samples $m$ for the entries of $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ has to guarantee that $2d\exp(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}})$ is close to zero.
 Since the factor involving the inverse kernel matrix $\mathbf{K}_{\mathcal{X}}^{-1}$ depends on the size $d$ of data set $\mathcal{X}$, regularization as indicated in Corollary 4.9 prevents this expression from blowing up with $d$.
@@
 Let the assumptions of Theorem 2.6 hold and let $\lambda\geq 0$. Then, for every $\varepsilon>0$ and every $f\in\mathbb{H}$ with $\|f\|_{\mathbb{H}}\leq 1$, we have
-$\|(\mathcal{K}-\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}})f\|_{\infty}\leq\left(\|1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\right)\,\|\mathcal{I}-R_{\mathcal{X}}^{\lambda}\|_{\mathbb{H}\to C_{b}}\,+\varepsilon\cdot\|k\|_{\infty}^{1/2}\cdot\sqrt{\frac{d}{\lambda}}$ (46)
-with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right)$.
-Specifically, for the RKHS $\mathbb{H}=H^{\sigma_{n,l}}$ induced by the Wendland kernel $k_{n,l}$, where $\sigma_{n,l}=\frac{n+1}{2}+l$, we have the pointwise error bound
-$\|(\mathcal{K}-\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}})f\|_{\infty}\leq C_{1}\left(h_{\mathcal{X}}^{l+1/2}+\lambda^{1/2}\right)+\varepsilon\cdot C_{2}\cdot\sqrt{\frac{d}{\lambda}}$ (47)
-for every $\varepsilon>0$ and all $f\in H^{\sigma_{n,l}}$ with $\|f\|_{H^{\sigma_{n,l}}}\leq 1$ with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2C_{2}^{2}}\right)$ and constants $C_{1},C_{2}>0$ only depending on $k_{n,l}$.
+$\big\|\big(\mathcal{K}-\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}\big)f\big\|_{\infty}\leq\left(\|1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\right)\,\|\mathcal{I}-R_{\mathcal{X}}^{\lambda}\|_{\mathbb{H}\to C_{b}}\,+\varepsilon\cdot\|\text{\sf{k}}\|_{\infty}^{1/2}\cdot\sqrt{\frac{d}{\lambda}}$ (47)
+with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right)$.
+Specifically, for the RKHS $\mathbb{H}=H^{\sigma_{n,l}}(\Omega)$ induced by the Wendland kernel $\text{\sf{k}}_{n,l}$, where $\sigma_{n,l}=\frac{n+1}{2}+l$, we have the pointwise error bound
+$\big\|\big(\mathcal{K}-\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}\big)f\big\|_{\infty}\leq C_{1}\left(h_{\mathcal{X}}^{l+1/2}+\lambda^{1/2}\right)+\varepsilon\cdot C_{2}\cdot\sqrt{\frac{d}{\lambda}}$ (48)
+for every $\varepsilon>0$ and all $f\in H^{\sigma_{n,l}}(\Omega)$ with $\|f\|_{H^{\sigma_{n,l}}}\leq 1$ with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2C_{2}^{2}}\right)$ and constants $C_{1},C_{2}>0$ only depending on $\text{\sf{k}}_{n,l}$.
 Proof.
@@
 and bounding the deterministic first part by Proposition 4.5. The estimation error in the second part is bounded by Corollary 4.9 with the respective probability and yields (47).
-The second claim combines the second part of Proposition 4.5 for the deterministic error bound with setting $C_{2}=\|k_{n,l}\|_{\infty}^{1/2}$, obtaining (47). ∎
+The second claim combines the second part of Proposition 4.5 for the deterministic error bound with setting $C_{2}=\|\text{\sf{k}}_{n,l}\|_{\infty}^{1/2}$, obtaining (48). ∎
 The Matérn kernel analogue to the Wendland special case of Corollary 4.12 is stated in Corollary C.3 in Appendix C.
@@
 $dX_{t}$ $=b(X_{t})\,dt\,+\sigma(X_{t})\,dW_{t},$ (49)
-$X_{0}$ $=x_{0}\in\Omega$
-with Lipschitz continuous drift $b:\mathbb{R}\to\mathbb{R}$ and diffusion $\sigma:\mathbb{R}\to\mathbb{R}$, and Brownian motion $(W_{t})_{t\geq 0}$, via the Euler-Maruyama method (see, e.g., [KloePlat1992]).
-[Okse2000, Theorem 5.2.1] guarantees the existence of a unique solution $(X_{t})_{t\geq 0}$.
+$X_{0}$ $=x_{0}\in\Omega,$
+with Lipschitz continuous drift $b:\mathbb{R}\to\mathbb{R}$ and diffusion $\sigma:\mathbb{R}\to\mathbb{R}$, and Brownian motion $(W_{t})_{t\geq 0}$, via the Euler-Maruyama method (see, e.g., [21]).
+[37, Theorem 5.2.1] guarantees the existence of a unique solution $(X_{t})_{t\geq 0}$.
 A basic example for this case is the Ornstein-Uhlenbeck process (OU process) given by the linear SDE
 $dX_{t}$ $=-\theta X_{t}\,dt\,+\sigma\,dW_{t},$ (50)
-$X_{0}$ $=x_{0}\in\Omega$
+$X_{0}$ $=x_{0}\in\Omega,$
 with drift coefficient $\theta>0$ and diffusion coefficient $\sigma>0$.
@@
 To apply the setting of Sections 3 and 4, the solution process should remain in some bounded domain $\Omega=(a,b)$ until time $T$.
-[PhilSchaWort2025] elaborates on a sufficient condition on the drift and diffusion for the invariance of a compact interval under the SDE flow.
+[41] elaborates on a sufficient condition on the drift and diffusion for the invariance of a compact interval under the SDE flow.
 Another way to guarantee the boundedness is by reflection at the boundaries.
-We consider this doubly reflected OU process on $\Omega=(0,1)$.
+For this, we consider the doubly reflected OU process on $\Omega=(0,1)$.
 It is formally given by the SDE
@@
 $X_{0}$ $=x_{0}\in(0,1),$
-where the $L_{t}\geq 0$ and $U_{t}\geq 0$ are the minimal nondecreasing processes that ensure $X_{t}\geq 0$ and $X_{t}\leq 1$ with reflection at the boundaries for $t\geq 0$, respectively.
+where $L_{t}\geq 0$ and $U_{t}\geq 0$ are the minimal nondecreasing processes that ensure $X_{t}\geq 0$ and $X_{t}\leq 1$ with reflection at the boundaries for $t\geq 0$, respectively.
 The processes $L$ and $U$ can be interpreted as “idleness process” and “loss process”; they satisfy $\int_{0}^{T}\chi_{\{X_{t}>0\}}\,dL_{t}=0$ and $\int_{0}^{T}\chi_{\{X_{t}<1\}}\,dU_{t}=0$ for any $T>0$, respectively.
-We refer to [HuanMandSpre2014] for details.
+We refer to [18] for details.
 As an example of a nonlinear stochastic differential equation, we consider the double-well potential $V(x)=\frac{1}{2}x^{2}-\frac{1}{4}x^{4}$ and set the drift to $b(x)=V^{\prime}(x)$, yielding the SDE
 $dX_{t}$ $=(X_{t}-X_{t}^{3})\,dt\,+\sigma\,dW_{t},$ (52)
-$X_{0}$ $=x_{0}\in\Omega$
+$X_{0}$ $=x_{0}\in\Omega,$
 with $\sigma>0$.
@@
 3:Form the kernel matrix
-$\mathbf{K}_{\mathcal{X}}=\bigl(k(x_{i},x_{j})\bigr)_{i,j=1}^{d}.$
+$\mathbf{K}_{\mathcal{X}}=\bigl(\text{\sf{k}}(x_{i},x_{j})\bigr)_{i,j=1}^{d}.$
 4:Define $f_{\mathcal{X}}=(f(x_{1}),\dots,f(x_{d}))^{\top}$.
@@
 7:Estimate the propagated kernel matrix
-$(\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}})_{ij}=\frac{1}{m_{\mathrm{train}}}\sum_{m=1}^{m_{\mathrm{train}}}k(x_{i},y_{j}^{(m)}).$
+$(\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}})_{ij}=\frac{1}{m_{\mathrm{train}}}\sum_{m=1}^{m_{\mathrm{train}}}\text{\sf{k}}(x_{i},y_{j}^{(m)}).$
 8:Koopman estimate
@@
 10:  Form kernel vector
-$k_{\mathcal{X}}(x_{\ell}^{\mathrm{grid}})=(k(x_{1},x_{\ell}^{\mathrm{grid}}),\dots,k(x_{d},x_{\ell}^{\mathrm{grid}}))^{\top}.$
+$\mathbf{k}_{\mathcal{X}}(x_{\ell}^{\mathrm{grid}})=(\text{\sf{k}}(x_{1},x_{\ell}^{\mathrm{grid}}),\dots,\text{\sf{k}}(x_{d},x_{\ell}^{\mathrm{grid}}))^{\top}.$
 11:  Compute the estimated Koopman action
-$\widehat{\mathcal{K}}^{\mathrm{MC}}f(x_{\ell}^{\mathrm{grid}})=f_{\mathcal{X}}^{\top}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\,\mathbf{K}_{\mathcal{X}}^{-1}\,k_{\mathcal{X}}(x_{\ell}^{\mathrm{grid}}).$
+$\widehat{\mathcal{K}}^{\mathrm{MC}}f(x_{\ell}^{\mathrm{grid}})=f_{\mathcal{X}}^{\top}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{k}_{\mathcal{X}}(x_{\ell}^{\mathrm{grid}}).$
 12:end for
@@
 In the setting of the stochastic Koopman operator (3), we set $X=X_{0}$ and $Y=X_{T}$.
-This means that we are interested in the evolution of observables $f:\Omega\to\mathbb{R}$ under the stochastic process that solves (48), i.e.
+This means that we are interested in the evolution of observables $f:\Omega\to\mathbb{R}$ under the stochastic process that solves (49), i.e.,
 $\mathcal{K}f(x)=\mathbb{E}[f(X_{T})\mid X_{0}=x].$ (53)
 For numerical testing, we can generally proceed as usual in kernel EDMD methods using the approximation (42), that is,
-$\widehat{\mathcal{K}}^{\mathrm{MC}}f=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}.$
+$\widehat{\mathcal{K}}^{\mathrm{MC}}f=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}},$
 or the regularized approximant (45), i.e.,
-$\widehat{\mathcal{K}}^{\mathrm{MC}}f=f_{\mathcal{X}}^{\top}(\mathbf{K}_{\mathcal{X}}^{-1}+\lambda I)^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}.$
+$\widehat{\mathcal{K}}^{\mathrm{MC}}f=f_{\mathcal{X}}^{\top}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}.$
 Aside from the observable $f:\Omega\to\mathbb{R}$ and design points $\mathcal{X}=\{x_{1},\dots,x_{d}\}\subset\Omega$ that need to be chosen beforehand, we introduce the parameter $m_{\mathrm{train}}\in\mathbb{N}$ that specifies the number of samples we draw from the conditional distributions $\rho_{x_{j}}$, $j=1,\dots,d$.
@@
   1. (1)
-    Conditional expectations of the form $\mathcal{K}f(x_{0})=\mathbb{E}[f(X_{T})\mid X_{0}=x_{0}]$ for some $x_{0}\in\Omega$ usually do not necessarily have an analytical expression.
-Therefore, we do not have access to the exact action of the Koopman operator on the observable $f$ (”ground truth”).
+    Conditional expectations of the form $\mathcal{K}f(x_{0})=\mathbb{E}[f(X_{T})\mid X_{0}=x_{0}]$ for some $x_{0}\in\Omega$ do not necessarily admit an analytical expression.
+Therefore, we usually do not have access to the exact action of the Koopman operator on the observable $f$ (”ground truth”).
 Instead, we use a further Monte Carlo approximation
@@
     The randomness of the propagated matrix $K_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ necessitates a comparison over multiple realizations of the Koopman approximation.
-In our SDE setting, this means that we build matrices $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC,n}}$, $n=1,\dots,n_{\mathrm{pred}}$ and use the mean values and standard deviations of the empirical $L^{\infty}$-errors $\widehat{\mathrm{err}}^{(n)}$ per realization to judge the performance of the approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$.
+In our SDE setting, this means that we build matrices $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC,n}}$, $n=1,\dots,n_{\mathrm{pred}}$, and use the mean values and standard deviations of the empirical $L^{\infty}$-errors $\widehat{\mathrm{err}}^{(n)}$ per realization to judge the performance of the approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$.
 Theorem 2.6 decomposes the $L^{\infty}$-error of the stochastic Koopman approximation into a deterministic part controlled by the fill distance $h_{\mathcal{X}}$ and a probabilistic part influenced by the Monte Carlo sample size, which we denote by $m_{\mathrm{train}}$ to differentiate it from the number of Monte Carlo samples $m_{\mathrm{truth}}$ used for ground truth approximation.
-For numerical testing, we specify $h_{\mathcal{X}}$ (or corresponding design points $\mathcal{X}$) and $m_{\mathrm{train}}$.
 Throughout the testing procedure, we frequently sample the SDE solution with initial values $x_{0}\in\Omega$ at time $T>0$ from the conditional distribution $\rho_{x_{0},T}$.
 First, we start trajectories at chosen evaluation grid points $\mathcal{X}^{\mathrm{grid}}\subset\Omega$ for ground truth computation.
-Later, we repeatedly simulate trajectories starting at the design points $\mathcal{X}$ to construct realizations $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ in the Koopman approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$ Algorithm 2 illustrates the procedure.
+Later, we repeatedly simulate trajectories starting at the design points $\mathcal{X}$ to construct realizations $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ in the Koopman approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$.
+For numerical demonstration on the domain $\Omega=(0,1)$, we first chose fill distances $h_{\mathcal{X}}$ such that $d=2/h_{\mathcal{X}}\in\mathbb{N}$.
+The fill distance was then achieved by spacing out the $d$ design points $\mathcal{X}=\{x_{1},\dots,x_{d}\}$ evenly on $\Omega=(0,1)$: We set $x_{1}=h_{\mathcal{X}}$, $x_{d}=1-h_{\mathcal{X}}$, and $x_{i}=(2i-1)h_{\mathcal{X}}$ for $i=2,\dots,d-1$.
+Algorithm 2 illustrates the procedure.
 Algorithm 2 Error Analysis for the Stochastic Kernel EDMD Koopman Approximation
@@
 13:return mean and standard deviation of $\widehat{\mathrm{err}}_{n}$, $n=1,\dots,n_{\mathrm{pred}}$.
-For numerical stability, we apply a Tikhonov regularization with parameter $\lambda$ as in (44) to alleviate the bad conditioning of kernel matrices $\mathbf{K}_{\mathcal{X}}$, .
+For numerical stability, we apply a Tikhonov regularization with parameter $\lambda$ as in (45) to alleviate the bad conditioning of kernel matrices $\mathbf{K}_{\mathcal{X}}$.
 The parameters $m_{\mathrm{train}}$ and $\lambda$ are occasionally adapted to the chosen fill distance $h_{\mathcal{X}}$ as indicated in the table of Figure 1.
@@
 We chose to compare the
-$\text{Wendland kernel}\quad k_{1,1}^{\text{wendland}}(x,y)$ $=(1-r)_{+}^{3}\cdot(3r+1),~\text{and}$
+$\text{Wendland kernel}\quad\text{\sf{k}}_{1,1}^{\text{wendland}}(x,y)$ $=(1-r)_{+}^{3}\cdot(3r+1),~\text{and}$
 […]
@@
 […]
-$X_{0}$ $=x_{0}\in\Omega$
+$X_{0}$ $=x_{0}\in\Omega,$
 with initial value $x_{0}$ and auxiliary processes $L$ and $U$ to achieve reflection at the boundaries of $\Omega=(0,1)$.
@@
 Figure 2.
-Ornstein-Uhlenbeck process: mean empirical $L^{\infty}$-error with standard deviation of $n_{\mathrm{pred}}=30$ realizations of the stochastic Koopman approximation in dependence of the number of Monte Carlo simulations $m_{\mathrm{train}}$.
+Ornstein-Uhlenbeck process: mean empirical $L^{\infty}$-error of $n_{\mathrm{pred}}=30$ realizations of the stochastic Koopman approximation in dependence of the number of Monte Carlo simulations $m_{\mathrm{train}}$.
 $\star$ and $\diamond$ indicate Wendland and Matérn kernels, respectively.
@@
 For better comparability, especially in terms of the fill distance, we want to restrict the corresponding SDE to the same domain $\Omega=(0,1)$.
-Since the wells in the regular double-well SDE (51) are at $\pm 1$, we apply the affine transformation $Z_{t}=(X_{t}+2)/4$ to mimic the behavior of the original double-well potential on (2,2)(-2,2) on our domain $\Omega=(0,1)$.
+Since the wells in the regular double-well SDE (52) are at $\pm 1$, we apply the affine transformation $Z_{t}=(X_{t}+2)/4$ to mimic the behavior of the original double-well potential on (2,2)(-2,2) on $\Omega=(0,1)$.
 Solving for $X_{t}$ gives $X_{t}=4Z_{t}-2$, so that the chain rule yields the SDE
-$dZ_{t}$ $=\frac{1}{4}\left[(4Z_{t}-2)-(4Z_{t}-2)^{3}\right]\,dt\,+\frac{\sigma}{4}\,dW_{t}+dL_{t}-dU_{t}$ (55)
-$Z_{0}$ $=z_{0}\in\Omega$
+$dZ_{t}$ $=\frac{1}{4}\left[(4Z_{t}-2)-(4Z_{t}-2)^{3}\right]\,dt\,+\frac{\sigma}{4}\,dW_{t}+dL_{t}-dU_{t},$ (56)
+$Z_{0}$ $=z_{0}\in\Omega,$
 with the auxiliary processes $L$ and $U$ achieving reflection at the boundaries $a=0$ and $b=1$.
@@
 We proposed a Monte Carlo-based estimator in a kernel EDMD approach for the Koopman operator of stochastic systems and analyzed its pointwise error.
-Extending the results of [KohnPhil24], we decompose the error into the well-known deterministic projection error and the probabilistic error introduced by the entry-wise estimation of expected values in the exact “propagation matrix” $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$.
-To apply known projection error bounds, we derive sufficient conditions for the invariance of (fractional) Sobolev spaces $H^{s}$ under the stochastic Koopman operator, which are induced as RKHS by Wendland and Matérn kernels.
+Extending the results of [24], we decompose the error into the well-known deterministic projection error and the probabilistic error introduced by the entry-wise estimation of expected values in the exact “propagation matrix” $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$.
+To apply known projection error bounds, we derive sufficient conditions for the invariance of (fractional) Sobolev spaces $H^{s}(\Omega)$ under the stochastic Koopman operator, which are induced as RKHS by Wendland and Matérn kernels.
 Novel probabilistic error bounds for our estimator are proven to obtain a pointwise bound on the full approximation error with high probability.
@@
 $\leq\sum_{j=1}^{d}\mathbb{P}\left(\left|\sum_{l=1}^{m_{j}}\big[(S_{\mathcal{X}}f)(y_{l,j})-\mathbb{E}_{j}\big]\right|\,\geq\,\frac{\varepsilon m_{j}\|f\|_{H}}{\sqrt{d}}\right)$
-$\leq 2\sum_{j=1}^{d}\exp\left(-\frac{m_{j}\varepsilon^{2}}{2d\|k\|_{\infty}}\right).$
+$\leq 2\sum_{j=1}^{d}\exp\left(-\frac{m_{j}\varepsilon^{2}}{2d\|\text{\sf{k}}\|_{\infty}}\right).$
 Hence, if $m_{j}=m$ for all $j\in\{1,\dots,d\}$, then
-$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{2}\geq\varepsilon\|f\|_{H})\,\leq\,2d\exp\left(-\frac{m\varepsilon^{2}}{2d\|k\|_{\infty}}\right).$
+$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{2}\geq\varepsilon\|f\|_{H})\,\leq\,2d\exp\left(-\frac{m\varepsilon^{2}}{2d\|\text{\sf{k}}\|_{\infty}}\right).$
 We may improve this by using the Hilbert-space Hoeffding inequality.
 For this, let $m_{j}=m$ for all $j$, and set $\xi_{l}=[(S_{\mathcal{X}}f)(y_{l,j})]_{j=1}^{d}-g\in\mathbb{R}^{d}$, $l\in\{1,\dots,m\}$.
-Then $\mathbb{E}[\xi_{l}]=0$ and
+Then, we have $\mathbb{E}[\xi_{l}]=0$ and
 $\|\xi_{l}\|_{2}^{2}=\sum_{j=1}^{d}\left|(S_{\mathcal{X}}f)(y_{l,j})-\int(S_{\mathcal{X}}f)(y)\,\rho(x,dy)\right|^{2}.$
-Since $|(S_{\mathcal{X}}f)(y)|\leq\|k\|_{\infty}^{1/2}\|f\|_{H}$, we obtain $\|\xi_{l}\|_{2}^{2}\leq 4d\|k\|_{\infty}\|f\|_{H}^{2}$ for all $l\in\{1,\dots,m\}$. Furthermore,
-$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{2}\geq\varepsilon\|f\|_{H})$ $=\mathbb{P}\left(\sum_{j=1}^{d}\left|\frac{1}{m}\sum_{l=1}^{m}\xi_{l,j}\right|^{2}\,\geq\,\varepsilon^{2}\|f\|_{H}^{2}\right)$
-$=\mathbb{P}\left(\left\|\frac{1}{m}\sum_{l=1}^{m}\xi_{l}\right\|\,\geq\,\varepsilon\|f\|_{H}\right)$
-$\leq 2\exp\left(-\frac{m\varepsilon^{2}}{8d\|k\|_{\infty}}\right).$
+Since $|(S_{\mathcal{X}}f)(y)|\leq\|\text{\sf{k}}\|_{\infty}^{1/2}\|f\|_{H}$, we obtain $\|\xi_{l}\|_{2}^{2}\leq 4d\|\text{\sf{k}}\|_{\infty}\|f\|_{H}^{2}$ for all $l\in\{1,\dots,m\}$. Also,
+$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{2}\geq\varepsilon\|f\|_{H})$ $=\mathbb{P}\left(\sum_{j=1}^{d}\left|\frac{1}{m}\sum_{l=1}^{m}\xi_{l,j}\right|^{2}\,\geq\,\varepsilon^{2}\|f\|_{H}^{2}\right)=\mathbb{P}\left(\left\|\frac{1}{m}\sum_{l=1}^{m}\xi_{l}\right\|\,\geq\,\varepsilon\|f\|_{H}\right)$
+$\leq 2\exp\left(-\frac{m\varepsilon^{2}}{8d\|\text{\sf{k}}\|_{\infty}}\right).$
 We are left with estimating $\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{2}$.
@@
 $\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{2}$ $=\sup_{\|v\|_{2}=1}|v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)|=\sup_{f\in V_{\mathcal{X}},\,f\neq 0}\frac{|f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)|}{\|f_{\mathcal{X}}\|_{2}}=\sup_{f\in V_{\mathcal{X}},\,f\neq 0}\frac{|f(x)|}{\|f_{\mathcal{X}}\|_{2}}$
-$=\sup_{f\in V_{\mathcal{X}},\,f\neq 0}\frac{|\langle f,\Phi_{x}\rangle_{H}|}{\|f_{\mathcal{X}}\|_{2}}\,\leq\,\|k\|_{\infty}^{1/2}\sup_{f\in V_{\mathcal{X}},\,f\neq 0}\frac{\|f\|_{H}}{\|f_{\mathcal{X}}\|_{2}}=\|k\|_{\infty}^{1/2}\|\mathbf{K}_{\mathcal{X}}^{-1}\|_{2\to 2}^{1/2}.$
+$=\sup_{f\in V_{\mathcal{X}},\,f\neq 0}\frac{|\langle f,\Phi_{x}\rangle_{H}|}{\|f_{\mathcal{X}}\|_{2}}\,\leq\,\|\text{\sf{k}}\|_{\infty}^{1/2}\sup_{f\in V_{\mathcal{X}},\,f\neq 0}\frac{\|f\|_{H}}{\|f_{\mathcal{X}}\|_{2}}=\|\text{\sf{k}}\|_{\infty}^{1/2}\|\mathbf{K}_{\mathcal{X}}^{-1}\|_{2\to 2}^{1/2}.$
 Alternatively, by the $\ell^{1}$-$\ell^{2}$-inequality, one may bound the deterministic term as in Theorem 4.6 by $\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{2}\leq\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}\leq\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$.
 Appendix B Recap: Wendland kernels
-Wendland kernels are constructed using particular compactly supported radial basis functions, which are constructed in the following way: For some integer $m\in\mathbb{N}$, define the a function $\beta_{m}:\mathbb{R}_{+}\to\mathbb{R},\beta_{m}(r)=(1-r)_{+}^{m}=\max(1-r,0)^{m}$.
+Wendland kernels are constructed using particular compactly supported radial basis functions, which are constructed in the following way: For $m\in\mathbb{N}$, define the function $\beta_{m}:\mathbb{R}_{+}\to\mathbb{R},\beta_{m}(r)=(1-r)_{+}^{m}{=\max(1-r,0)^{m}}$.
 Given the operator $\mathcal{I}$ that acts on measurable non-negative functions $\beta:\mathbb{R}_{+}\to\mathbb{R}_{+}$ by
-$\mathcal{I}\beta(r)=\int_{r}^{\infty}t\beta(t)\,dt$
-and dimension $n\geq 1$, we choose a smoothness $l\in\mathbb{N}$ and set $\beta_{n,l}=\mathcal{I}^{l}\beta_{\lfloor\frac{n}{2}\rfloor+l+1}$. This function is of the form
+$\mathcal{I}\beta(r)=\int_{r}^{\infty}t\beta(t)\,dt,$
+and given dimension $n\geq 1$, we choose a smoothness $l\in\mathbb{N}$ and set $\beta_{n,l}=\mathcal{I}^{l}\beta_{\lfloor\frac{n}{2}\rfloor+l+1}$. This function is of the form
 $\beta_{n,l}(r)=\begin{cases}p_{n,l}(r),&0\leq r\leq 1,\\ 0,&r>1\end{cases}$
-with a univariate polynomial $p_{n,l}$ of degree $\lfloor\frac{n}{2}\rfloor+3l+1$ whose coefficients can be computed recursively by [Wend04, Theorem 9.12].
+with a univariate polynomial $p_{n,l}$ of degree $\lfloor\frac{n}{2}\rfloor+3l+1$ whose coefficients can be computed recursively by [50, Theorem 9.12].
 We can now define the Wendland radial basis function $B_{n,m}:\mathbb{R}^{n}\to\mathbb{R}$ of dimension $n$ and smoothness degree $m$ by
 $B_{n,l}(x):=\beta_{n,l}(\|x\|_{2})$
-and the corresponding Wendland kernel by $k_{n,l}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R},k_{n,l}(x,y)=B_{n,l}(x-y)=\beta_{n,l}(\|x-y\|_{2})$.
+and the corresponding Wendland kernel by $\text{\sf{k}}_{n,l}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R},\text{\sf{k}}_{n,l}(x,y)=B_{n,l}(x-y)=\beta_{n,l}(\|x-y\|_{2})$.
 Appendix C Error bounds for Matérn kernels
@@
 Specifically, the Matérn covariance function $\mathcal{M}_{\nu,\alpha}:\mathbb{R}_{+}\to\mathbb{R}$ with smoothness $\nu>0$ and length-scale parameter $\alpha>0$ is given by
-$\mathcal{M}_{\nu,\alpha}(r)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}\,r}{\alpha}\right)^{\nu}K_{\nu}\left(\frac{\sqrt{2\nu}\,r}{\alpha}\right)$
-where $\Gamma$ denotes the gamma function and $K_{\nu}$ denotes the modified Bessel function of the second kind.
-Denoting the Euclidean distance of $x,y\in\mathbb{R}^{n}$ by $\|\cdot\|_{2}$, the Matérn kernel $k_{\mu,\alpha}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ is given by
-$k_{\nu,\alpha}(x,y)=\mathcal{M}_{\nu,\alpha}(\|x-y\|_{2}).$
-The parameter $\nu$ controls the differentiability of the Matérn kernel at $x=y$ through $k_{\nu,\alpha}\in C^{2\lfloor\nu\rfloor}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ – with increasing $\nu$, the Matérn kernels become smoother.
+$\mathcal{M}_{\nu,\alpha}(r)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}\,r}{\alpha}\right)^{\nu}\text{\sf{k}}_{\nu}\left(\frac{\sqrt{2\nu}\,r}{\alpha}\right),$
+where $\Gamma$ denotes the gamma function and $\text{\sf{k}}_{\nu}$ denotes the modified Bessel function of the second kind.
+Denoting the Euclidean distance of $x,y\in\mathbb{R}^{n}$ by $\|\cdot\|_{2}$, the Matérn kernel $\text{\sf{k}}_{\nu,\alpha}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ is given by
+$\text{\sf{k}}_{\nu,\alpha}(x,y)=\mathcal{M}_{\nu,\alpha}(\|x-y\|_{2}).$
+The parameter $\nu$ controls the differentiability of the Matérn kernel at $x=y$ through $\text{\sf{k}}_{\nu,\alpha}\in C^{2\lfloor\nu\rfloor}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ – with increasing $\nu$, the Matérn kernels become smoother.
 Equivalently, the sample paths of a Gaussian process with the Matérn kernel as covariance function are $\lfloor\nu\rfloor$-times mean-square differentiable, which provides a flexible way to control smoothness and makes Matérn kernels a popular choice in this setting.
@@
 to the squared exponential (Gaussian) kernel, uniformly on compact sets.
-Of particular interest are the cases $\nu=s+1/2$ for $s\in\mathbb{N}$: In that case, the Matérn covariance function is given as the product of a negative exponential and a polynomial of order $s$ [PorcBeviSchabOate2024].
-Up to scaling constants, the Matérn kernel $k_{\nu,\alpha}$ on $\mathbb{R}^{n}$ induces the Sobolev space $H^{\nu+n/2}$ with equivalent norms; for the dimensionless Matérn kernel $k_{\nu}:=k_{\nu,\sqrt{2\nu}}$ the RKHS norm even coincides with the standard Sobolev norm of order $\nu+n/2$ up to a constant factor.
-We refer the reader to [FassYe11, PorcBeviSchabOate2024] for more insights into Matérn kernels.
-Matérn kernels $k_{\nu}(x,y)=\mathcal{M}_{\nu,1}(\|x-y\|_{2})$ of smoothness $\nu>0$ induce the Sobolev spaces $H^{\nu+n/2}$.
-Denote the native norm of the RKHS by $\|f\|_{\mathbb{H}}$.
+Of particular interest are the cases $\nu=s+1/2$ for $s\in\mathbb{N}$: In that case, the Matérn covariance function is given as the product of a negative exponential and a polynomial of order $s$ [42].
+Up to scaling constants, the Matérn kernel $\text{\sf{k}}_{\nu,\alpha}$ on $\mathbb{R}^{n}$ induces the Sobolev space $H^{\nu+n/2}$ with equivalent norms; for the dimensionless Matérn kernel $\text{\sf{k}}_{\nu}:=\text{\sf{k}}_{\nu,\sqrt{2\nu}}$, the RKHS norm even coincides with the standard Sobolev norm of order $\nu+n/2$ up to a constant factor.
+We refer the reader to [14, 42] for more insights into Matérn kernels.
+Matérn kernels $\text{\sf{k}}_{\nu}(x,y)=\mathcal{M}_{\nu,1}(\|x-y\|_{2})$ of smoothness $\nu>0$ induce the Sobolev spaces $H^{\nu+n/2}$.
+Denote the native norm of the RKHS by $\|\cdot\|_{\mathbb{H}}$.
 The Matérn kernel is $2\lfloor\nu\rfloor$-times differentiable.
 Furthermore, (normalized) Matérn kernels are uniformly bounded by $1$.
-Thus, according to the discussion after [Wend04, Theorem 11.13], the bound
+Thus, according to the discussion after [50, Theorem 11.13], the bound
 $\|(\mathcal{I}-S_{\mathcal{X}})f\|_{\infty}\leq Ch_{\mathcal{X}}^{\lfloor\nu\rfloor}\|f\|_{\mathbb{H}}$
 with a constant $C>0$ holds for all $f\in\mathbb{H}$.
-Using the above properties, we obtain quasi-analogous results as in Section 4 for kernel EDMD approximations with underlying Matérn kernels.
+Using the above properties, we obtain analogous results as in Section 4 for kernel EDMD approximations with underlying Matérn kernels.
 Following the derivations for Wendland kernels, we start with the deterministic error bound characterized by the projection error.
@@
 Let $\nu>0$.
-Denote by $k_{\nu}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ the (dimensionless) Matérn kernel generating the Sobolev space $H^{\nu+n/2}(\Omega)$ with equivalence between the native norm $\|\cdot\|_{\mathbb{H}}$ and the classical Sobolev norm $\|\cdot\|_{H^{\nu+n/2}}$.
+Denote by $\text{\sf{k}}_{\nu}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ the (dimensionless) Matérn kernel generating the Sobolev space $H^{\nu+n/2}(\Omega)$ with equivalence between the native norm $\|\cdot\|_{\mathbb{H}}$ and the classical Sobolev norm $\|\cdot\|_{H^{\nu+n/2}}$.
 Assume that $\mathcal{K}\in L(C_{b},C_{b})$ and that (24) is satisfied for some $\ell\in\mathbb{N}$ with $\ell>\nu+n/2$.
 Then, the Koopman approximation $\widehat{\mathcal{K}}$ from (32) satisfies
-$\|\mathcal{K}-\widehat{\mathcal{K}}\|_{\mathbb{H}\to C_{b}}\leq C(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot h_{\mathcal{X}}^{\lfloor\nu\rfloor}$ (56)
+$\|\mathcal{K}-\widehat{\mathcal{K}}\|_{\mathbb{H}\to C_{b}}\leq C(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot h_{\mathcal{X}}^{\lfloor\nu\rfloor}.$ (57)
 Proof.
-The Matérn kernel satisfies $k_{\nu}\in C^{\lfloor 2\nu\rfloor}(\mathbb{R}^{n}\times\mathbb{R}^{n})$. According to [Wend04, Theorem 11.13] and the subsequent discussion, we have
+The Matérn kernel satisfies $\text{\sf{k}}_{\nu}\in C^{\lfloor 2\nu\rfloor}(\mathbb{R}^{n}\times\mathbb{R}^{n})$. According to [50, Theorem 11.13] and the subsequent discussion, we have
 $\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}\leq h_{\mathcal{X}}^{\lfloor\nu\rfloor}$
@@
 ∎
-Since the probabilistic error bound for the Matérn case is virtually identical to the Wendland case, merely inserting $\|k_{\nu}\|_{\infty}=1$, we state the analogue to Theorem 2.8 for the Matérn case below.
+Since the probabilistic error bound for the Matérn case is virtually identical to the Wendland case, merely inserting $\|\text{\sf{k}}_{\nu}\|_{\infty}=1$, we state the analogue to Theorem 2.8 for the Matérn case below.
 Corollary C.2 ($L^{\infty}$-error bound for Matérn native spaces).
 Let the assumptions of Theorem 2.6 hold.
-Let $\nu>0$ and denote by $\mathbb{H}=H^{\nu+n/2}$ the reproducing kernel Hilbert space induced by the Matérn kernel $k_{\nu}$.
-Then, for every $\varepsilon>0$ and all $f\in\mathbb{H}$ with $\|f\|_{\mathbb{H}}$, the full error admits the bound
+Let $\nu>0$ and denote by $\mathbb{H}=H^{\nu+n/2}(\Omega)$ the reproducing kernel Hilbert space induced by the Matérn kernel $\text{\sf{k}}_{\nu}$.
+Then, for every $\varepsilon>0$ and all $f\in\mathbb{H}$ with $\|f\|_{\mathbb{H}}\leq 1$, the full error admits the bound
 $\|(\mathcal{K}-\mathcal{K}^{\mathrm{MC}})f\|_{\infty}\leq$ $C\,(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot h_{\mathcal{X}}^{\lfloor\nu\rfloor}+\varepsilon\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$ (58)
-with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2}\right)$.
+with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2}\right)$ and $\mathbbm{1}=\{v\in\mathbb{R}^{d}:|v_{i}|=1\}$.
 Proof.
-the Matérn covariance function is monotonically decreasing for $r>0$, admitting its maximum at $\mathcal{M}_{\nu,\alpha}(0)=1$, so $\|k_{\nu}\|_{\infty}=1$.
+The Matérn covariance function is monotonically decreasing for $r>0$, admitting its maximum at $\mathcal{M}_{\nu,\alpha}(0)=1$, so $\|\text{\sf{k}}_{\nu}\|_{\infty}=1$.
 Similarly to Theorem 2.8, the result now follows immediately from Theorem 2.6.
@@
 $\|\mathcal{K}f-\widehat{\mathcal{K}}_{\lambda}f\|_{\infty}\leq C\left(h_{\mathcal{X}}^{\lfloor\nu\rfloor}+\lambda^{1/2}\right)\|f\|_{\mathbb{H}}$ (59)
-for $\mathbb{H}=H^{\nu+n/2}$, following the line of argument in Proposition 4.5. To conclude, we state the analogue to the second part Corollary 4.12 for Matérn native spaces.
+for $\mathbb{H}=H^{\nu+n/2}(\Omega)$, following the line of argument in Proposition 4.5. To conclude, we state the analogue to the second part Corollary 4.12 for Matérn native spaces.
 Corollary C.3 (Regularized $L^{\infty}$-error bounds for Matérn native spaces).
-Let the assumptions of Theorem 2.6 hold and let $\lambda\geq 0$. Then, for the RKHS $\mathbb{H}=H^{\nu+n/2}$ induced by the Matérn kernel $k_{\nu}$, we have the pointwise error bound
+Let the assumptions of Theorem 2.6 hold and let $\lambda\geq 0$.
+Then, for the RKHS $\mathbb{H}=H^{\nu+n/2}(\Omega)$ induced by the Matérn kernel $\text{\sf{k}}_{\nu}$, we have the pointwise error bound
 $\|(\mathcal{K}-\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}})f\|_{\infty}\leq C\left(h_{\mathcal{X}}^{l+1/2}+\lambda^{1/2}\right)+\varepsilon\cdot\sqrt{\frac{d}{\lambda}}$ (60)
-for every $\varepsilon>0$ and all $f\in H^{\sigma_{n,l}}$ with $\|f\|_{\mathbb{H}}\leq 1$ with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2}\right)$ and constant $C>0$ only depending on $k_{\nu}$.
+for every $\varepsilon>0$ and all $f\in H^{\sigma_{n,l}}$ with $\|f\|_{\mathbb{H}}\leq 1$ with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2}\right)$ and constant $C>0$ only depending on $\text{\sf{k}}_{\nu}$.
 Proof.
-The proof is analogous to the second part of Corollary 4.12 and leverages the uniform bound $\|k_{\nu}\|_{\infty}=1$. ∎
-Appendix D Balancing between deterministic and probabilistic error
-We discuss here the practical choice of the parameters $h_{\mathcal{X}}$, $m_{\mathrm{train}}$ and $\lambda$ for the stochastic kernel EDMD method when using the Wendland $C^{2}$ kernel ($l=1$) with spatial dimension $n=1$.
+The proof is analogous to the second part of Corollary 4.12 and leverages the uniform bound $\|\text{\sf{k}}_{\nu}\|_{\infty}=1$. ∎
+Appendix D Balancing deterministic and probabilistic error
+Here, we discuss the practical choice of the parameters $h_{\mathcal{X}}$, $m_{\mathrm{train}}$ and $\lambda$ for the stochastic kernel EDMD method when using the Wendland $C^{2}$-kernel ($l=1$) in spatial dimension $n=1$.
 The derivation for the Matérn kernel can be done in a similar fashion.
@@
 As $h_{\mathcal{X}}$ decreases, the number $d$ of design points increases, and the smallest eigenvalue of the kernel matrix $\mathbf{K}_{\mathcal{X}}$ decreases.
-For the unregularized inverse, the bound $\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\leq\frac{d}{\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})}$ and $\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})=O(h_{\mathcal{X}}^{3})=O(d^{-3})$ give
-$\left(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X},\lambda}^{-1}v\right)^{1/2}=O\left(\sqrt{\frac{d}{d^{-3}}}\right)=O(d^{2})=O(h_{\mathcal{X}}^{-2}).$
+For the unregularized inverse, we start from the basic inequality $\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\leq\frac{d}{\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})}$.
+A classical result from kernel interpolation theory provides a lower bound on the smallest eigenvalue via the separation distance $q_{\mathcal{X}}=\min_{i\neq j}\|x_{i}-x_{j}\|_{2}$ of the set $\mathcal{X}=\{x_{1},\dots,x_{d}\}$: Specifically, we have $\lambda_{\min}(\mathbf{K}_{\mathcal{X}})\geq Cq_{\mathcal{X}}^{2s-n}$ with smoothness $s>n/2$ and dimension $n$ of the underlying kernel, see [51, 45, 50].
+For Wendland kernels $\text{\sf{k}}_{n,l}$, where $s=\frac{n+1}{2}+l$, this gives rate $q_{\mathcal{X}}^{2l+1}$.
+We assume equispaced points in dimension $n=1$, for which $O(q_{\mathcal{X}})=O(h_{\mathcal{X}})$ holds, such that $\lambda_{\min}(\mathbf{K}_{\mathcal{X}})=O(q_{\mathcal{X}}^{3})=O(h_{\mathcal{X}}^{3})=O(d^{-3})$, and
+$\left(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\right)^{1/2}=O\left(\sqrt{\frac{d}{d^{-3}}}\right)=O(d^{2})=O(h_{\mathcal{X}}^{-2}).$
 Under Tikhonov regularization, Corollary 4.12 shows that $\left(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X},\lambda}^{-1}v\right)^{1/2}=O\!\left(\sqrt{\frac{d}{\lambda}}\right)$.
@@
 $\|\mathcal{K}f-\widehat{\mathcal{K}}_{\lambda}f\|_{\infty}\leq C\left(h_{\mathcal{X}}^{3/2}+\lambda^{1/2}\right)\|f\|_{\mathbb{H}}.$
-To prevent the regularization bias from dominating, one should require $\lambda=O(h_{\mathcal{X}}^{3})$.
-Monte Carlo estimation of $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$ introduces a stochastic error. From Theorem 4.6, with probability at least $1-2de^{-m\varepsilon^{2}/(2\|k\|_{\infty})}$,
-$\|\widehat{\mathcal{K}}f-\widehat{\mathcal{K}}^{\mathrm{MC}}f\|_{\infty}\leq\varepsilon\,C_{\mathcal{X}}\|f\|_{\mathbb{H}}.$
-Solving the probability expression for $\varepsilon$ gives
-$\varepsilon=O(m^{-1/2}),$
-again ignoring logarithmic terms in $d$. Since equispaced grids satisfy $d=O(h_{\mathcal{X}}^{-1})$, the Tikhonov control factor becomes
+To prevent the regularization bias from dominating, one requires $\lambda=O(h_{\mathcal{X}}^{3})$.
+Moreover, Monte Carlo estimation of $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$ introduces a stochastic error.
+From Theorem 4.6, with probability at least $1-2de^{-m\varepsilon^{2}/(2\|\text{\sf{k}}\|_{\infty})}$,
+$\|\widehat{\mathcal{K}}f-\widehat{\mathcal{K}}^{\mathrm{MC}}f\|_{\infty}\leq\varepsilon\,C_{\mathcal{X}}\|f\|_{\mathbb{H}},$
+where $C_{\mathcal{X}}$ depends on the data points $\mathcal{X}$.
+Solving the probability expression for $\varepsilon$ gives $\varepsilon=O(m^{-1/2})$, again ignoring logarithmic terms in $d$.
+Since equispaced grids satisfy $d=O(h_{\mathcal{X}}^{-1})$, the Tikhonov control factor becomes
 $\sqrt{\max_{v}v^{\top}\mathbf{K}_{\mathcal{X},\lambda}^{-1}v}=O\!\left(\sqrt{\frac{d}{\lambda}}\right)=O\!\left(\frac{h_{\mathcal{X}}^{-1/2}}{h_{\mathcal{X}}^{3/2}}\right)=O(h_{\mathcal{X}}^{-2}).$
@@
 Since this is infeasible in practice, we conducted numerical experiments to test more moderate orders of growth of $m_{\mathrm{train}}$. We found the orders
-$m_{\mathrm{train}}(h_{\mathcal{X}})=O\!\left(h_{\mathcal{X}}^{-3/2}\right),\quad\lambda(h_{\mathcal{X}})=O\!\left(h_{\mathcal{X}}^{-3}\right)$
+$m_{\mathrm{train}}(h_{\mathcal{X}})=O\!\left(h_{\mathcal{X}}^{-3/2}\right),\quad\lambda(h_{\mathcal{X}})=O\!\left(h_{\mathcal{X}}^{-3}\right),$
 to produce satisfying balancing of the two error sources, as displayed in Figure 1.